GNU PROLOG

A Native Prolog Compiler with Constraint Solving over Finite Domains

Edition 1.50, for GNU Prolog version 1.5.0 May 12, 2021

by Daniel Diaz

Copyright (C) 1999-2021 Daniel Diaz

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Free Software Foundation^{1}, 51 Franklin St, Fifth Floor, Boston, MA 02110-1301, USA.

1http://www.fsf.org/

Contents

1 Acknowledgements  9
2 GNU Prolog License Conditions  11
3 Introduction  11
4 Using GNU Prolog  13
4.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13
4.2 The GNU Prolog interactive interpreter . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13
4.2.1 Starting/exiting the interactive interpreter . . . . . . . . . . . . . . . . . . . . . . 13
4.2.2 The interactive interpreter read-execute-write loop . . . . . . . . . . . . . . . . . 15
4.2.3 Consulting a Prolog program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17
4.2.4 Scripting Prolog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18
4.2.5 Interrupting a query . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18
4.2.6 The line editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18
4.3 Adjusting the size of Prolog data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20
4.4 The GNU Prolog compiler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21
4.4.1 Different kinds of codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21
4.4.2 Compilation scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21
4.4.3 Using the compiler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23
4.4.4 Running an executable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27
4.4.5 Generating a new interactive interpreter . . . . . . . . . . . . . . . . . . . . . . . 27
4.4.6 The name mangling scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28
5 Debugging  31
5.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31
5.2 The procedure box model . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31
5.3 Debugging predicates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31
5.3.1 Running and stopping the debugger . . . . . . . . . . . . . . . . . . . . . . . . . 31
5.3.2 Leashing ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32
5.3.3 Spy-points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32
5.4 Debugging messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33
5.5 Debugger commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33
5.6 The WAM debugger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35
6 Format of definitions  37
6.1 General format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37
6.2 Types and modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37
6.3 Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39
6.3.1 General format and error context . . . . . . . . . . . . . . . . . . . . . . . . . . . 39
6.3.2 Instantiation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39
6.3.3 Uninstantiation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40
6.3.4 Type error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40
6.3.5 Domain error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40
6.3.6 Existence error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41
6.3.7 Permission error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41
6.3.8 Representation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41
6.3.9 Evaluation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42
6.3.10 Resource error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42
6.3.11 Syntax error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42
6.3.12 System error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42
7 Prolog directives and control constructs  45

7.1 Prolog directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45
7.1.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45
7.1.2 dynamic/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45
7.1.3 public/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45
7.1.4 multifile/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46
7.1.5 discontiguous/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46
7.1.6 ensure linked/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

7.1.7 built in/0, built in/1, built in fd/0, built in fd/1 . . . . . . . . . . . . . . 47

7.1.8 include/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
7.1.9 if/1, else/0, endif/0, elif/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
7.1.10 ensure loaded/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
7.1.11 op/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
7.1.12 char conversion/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
7.1.13 set prolog flag/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
7.1.14 initialization/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
7.1.15 foreign/2, foreign/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
7.2 Prolog control constructs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51
7.2.1 true/0, fail/0, !/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51
7.2.2 (',')/2 - conjunction, (;)/2 - disjunction, (->)/2 - if-then, (*->)/2 - soft-cut (soft if-then) 51
7.2.3 call/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
7.2.4 catch/3, throw/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53
8 Prolog built-in predicates  55
8.1 Type testing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

8.1.1 var/1, nonvar/1, atom/1, integer/1, float/1, number/1, atomic/1,compound/1, callable/1, ground/1, is list/1, list/1,partial list/1, list or partial list/1 55

8.2 Term unification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56
8.2.1 (=)/2 - Prolog unification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56
8.2.2 unify with occurs check/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56
8.2.3 (\=)/2 - not Prolog unifiable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57
8.3 Term comparison . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57
8.3.1 Standard total ordering of terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57
8.3.2 (==)/2 - term identical, (\==)/2 - term not identical,(@<)/2 - term less than, (@=<)/2 - term less than or equal to,(@>)/2 - term greater than, (@>=)/2 - term greater than or equal to 57
8.3.3 compare/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58
8.4 Term processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59
8.4.1 functor/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59
8.4.2 arg/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59
8.4.3 (=..)/2 - univ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60
8.4.4 copy term/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60
8.4.5 term variables/2, term variables/3 . . . . . . . . . . . . . . . . . . . . . . . . 61
8.4.6 subsumes term/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61
8.4.7 acyclic term/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61
8.4.8 term hash/4, term hash/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62
8.4.9 setarg/4, setarg/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63
8.5 Variable naming/numbering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63
8.5.1 name singleton vars/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63
8.5.2 name query vars/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

8.5.3 bind variables/2, numbervars/3, numbervars/1 . . . . . . . . . . . . . . . . . . 64

8.5.4 term ref/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65
8.6 Arithmetic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66
8.6.1 Evaluation of an arithmetic expression . . . . . . . . . . . . . . . . . . . . . . . . 66
8.6.2 (is)/2 - evaluate expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69
8.6.3 (=:=)/2 - arithmetic equal, (=\=)/2 - arithmetic not equal,(<)/2 - arithmetic less than, (=<)/2 - arithmetic less than or equal to,(>)/2 - arithmetic greater than, (>=)/2 - arithmetic greater than or equal to 69
8.6.4 succ/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70
8.7 Dynamic clause management . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

8.7.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70
8.7.2 asserta/1, assertz/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71
8.7.3 retract/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72
8.7.4 retractall/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72
8.7.5 clause/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73
8.7.6 abolish/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73
8.8 Predicate information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74
8.8.1 current predicate/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74
8.8.2 predicate property/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75
8.9 All solutions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76
8.9.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76
8.9.2 findall/4, findall/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76
8.9.3 bagof/3, setof/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77
8.10 Streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78
8.10.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78
8.10.2 current input/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79
8.10.3 current output/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79
8.10.4 set input/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80
8.10.5 set output/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80
8.10.6 open/4, open/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81
8.10.7 close/2, close/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82
8.10.8 flush output/1, flush output/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 83
8.10.9 current stream/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83
8.10.10 stream property/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84
8.10.11 at end of stream/1, at end of stream/0 . . . . . . . . . . . . . . . . . . . . . . 85
8.10.12 stream position/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85
8.10.13 set stream position/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86
8.10.14 seek/4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86
8.10.15 character count/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87
8.10.16 line count/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88
8.10.17 line position/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88
8.10.18 stream line column/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89
8.10.19 set stream line column/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89
8.10.20 add stream alias/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90
8.10.21 current alias/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90
8.10.22 add stream mirror/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91
8.10.23 remove stream mirror/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91
8.10.24 current mirror/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92
8.10.25 set stream type/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92
8.10.26 set stream eof action/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93
8.10.27 set stream buffering/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93
8.11 Constant term streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94
8.11.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

8.11.2 open input atom stream/2, open input chars stream/2,open input codes stream/2 94
8.11.3 close input atom stream/1, close input chars stream/1,close input codes stream/1 95
8.11.4 open output atom stream/1, open output chars stream/1,open output codes stream/1 96
8.11.5 close output atom stream/2, close output chars stream/2,close output codes stream/2 96

8.12 Character input/output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

8.12.1 get char/2, get char/1, get code/1, get code/2 . . . . . . . . . . . . . . . . . 97
8.12.2 get key/2, get key/1 get key no echo/2, get key no echo/1 . . . . . . . . . . 98
8.12.3 peek char/2, peek char/1, peek code/1, peek code/2 . . . . . . . . . . . . . . . 99
8.12.4 unget char/2, unget char/1, unget code/2, unget code/1 . . . . . . . . . . . . 100
8.12.5 put char/2, put char/1, put code/1, put code/2, nl/1, nl/0 . . . . . . . . . . 100

8.13 Byte input/output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

8.13.1 get byte/2, get byte/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101
8.13.2 peek byte/2, peek byte/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102
8.13.3 unget byte/2, unget byte/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102
8.13.4 put byte/2, put byte/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103
8.14 Term input/output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104
8.14.1 read term/3, read term/2, read/2, read/1 . . . . . . . . . . . . . . . . . . . . . 104

8.14.2 read atom/2, read atom/1, read integer/2, read integer/1,read number/2, read number/1105

8.14.3 read token/2, read token/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106
8.14.4 syntax error info/4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107
8.14.5 last read start line column/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 107

8.14.6 write term/3, write term/2, write/2, write/1, writeq/2, writeq/1,write canonical/2, write canonical/1, display/2, display/1, print/2,print/1108

8.14.7 format/3, format/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110
8.14.8 portray clause/2, portray clause/1 . . . . . . . . . . . . . . . . . . . . . . . . 112
8.14.9 get print stream/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113
8.14.10 op/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113
8.14.11 current op/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115
8.14.12 char conversion/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116
8.14.13 current char conversion/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116
8.15 Input/output from/to constant terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117

8.15.1 read term from atom/3, read from atom/2, read token from atom/2 . . . . . . 117
8.15.2 read term from chars/3, read from chars/2, read token from chars/2 . . . . 117
8.15.3 read term from codes/3, read from codes/2, read token from codes/2 . . . . 118
8.15.4 write term to atom/3, write to atom/2, writeq to atom/2,write canonical to atom/2, display to atom/2, print to atom/2,format to atom/3118
8.15.5 write term to chars/3, write to chars/2, writeq to chars/2,write canonical to chars/2, display to chars/2, print to chars/2,format to chars/3119
8.15.6 write term to codes/3, write to codes/2, writeq to codes/2,write canonical to codes/2, display to codes/2, print to codes/2,format to codes/3120

8.16 DEC-10 compatibility input/output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120
8.16.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120
8.16.2 see/1, tell/1, append/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121
8.16.3 seeing/1, telling/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121
8.16.4 seen/0, told/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122
8.16.5 get0/1, get/1, skip/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122
8.16.6 put/1, tab/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123
8.17 Term expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123
8.17.1 Definite clause grammars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123
8.17.2 expand term/2, term expansion/2 . . . . . . . . . . . . . . . . . . . . . . . . . . 124
8.17.3 phrase/3, phrase/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125
8.18 Logic, control and exceptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

8.18.1 abort/0, stop/0, top level/0, break/0, halt/1, halt/0 . . . . . . . . . . . . . 126
8.18.2 false/0, once/1, (\+)/1 - not provable, call/2-11, call with args/1-11, call det/2, forall/2126

8.18.3 repeat/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127
8.18.4 between/3, for/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128
8.19 Atomic term processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128
8.19.1 atom length/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129
8.19.2 atom concat/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129
8.19.3 sub atom/5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129
8.19.4 char code/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130
8.19.5 lower upper/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130
8.19.6 atom chars/2, atom codes/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131

8.19.7 number atom/2, number chars/2, number codes/2 . . . . . . . . . . . . . . . . . 132

8.19.8 name/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133
8.19.9 new atom/2, new atom/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133
8.19.10 current atom/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134
8.19.11 atom property/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134
8.20 List processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

8.20.1 append/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135
8.20.2 member/2, memberchk/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135
8.20.3 reverse/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136
8.20.4 delete/3, select/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136
8.20.5 subtract/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136
8.20.6 permutation/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137
8.20.7 prefix/2, suffix/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137
8.20.8 sublist/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138
8.20.9 last/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138
8.20.10 flatten/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138
8.20.11 length/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139
8.20.12 nth/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139
8.20.13 max list/2, min list/2, sum list/2 . . . . . . . . . . . . . . . . . . . . . . . . . 139
8.20.14 maplist/2-8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140

8.20.15 sort/2, msort/2, keysort/2 sort/1, msort/1, keysort/1 . . . . . . . . . . . . . 140

8.21 Global variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141
8.21.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141
8.21.2 g assign/2, g assignb/2, g link/2 . . . . . . . . . . . . . . . . . . . . . . . . . 143
8.21.3 g read/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143
8.21.4 g array size/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144

8.21.5 g inc/3, g inc/2, g inco/2, g inc/1, g dec/3, g dec/2, g deco/2, g dec/1 . . . 144
8.21.6 g set bit/2, g reset bit/2, g test set bit/2, g test reset bit/2 . . . . . . 145

8.21.7 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146
8.22 Prolog state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148
8.22.1 set prolog flag/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148
8.22.2 current prolog flag/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151
8.22.3 set bip name/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151
8.22.4 current bip name/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151

8.22.5 write pl state file/1, read pl state file/1 . . . . . . . . . . . . . . . . . . . 152

8.23 Program state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153
8.23.1 consult/1, '.'/2 - program consult . . . . . . . . . . . . . . . . . . . . . . . . . 153
8.23.2 load/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153
8.23.3 listing/1, listing/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154
8.24 System statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155
8.24.1 statistics/0, statistics/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155

8.24.2 user time/1, system time/1, cpu time/1, real time/1 . . . . . . . . . . . . . . 155

8.25 Random number generator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156
8.25.1 set seed/1, randomize/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156
8.25.2 get seed/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156
8.25.3 random/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157
8.25.4 random/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157
8.26 File name processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158
8.26.1 absolute file name/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158

8.26.2 is absolute file name/1, is relative file name/1 . . . . . . . . . . . . . . . 158

8.26.3 decompose file name/4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159
8.26.4 prolog file name/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159
8.27 Operating system interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160
8.27.1 argument counter/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160
8.27.2 argument value/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160
8.27.3 argument list/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161
8.27.4 environ/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

8.27.5 make directory/1, delete directory/1, change directory/1 . . . . . . . . . . 162

8.27.6 working directory/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162
8.27.7 directory files/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162

8.27.8 rename file/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163
8.27.9 delete file/1, unlink/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163
8.27.10 file permission/2, file exists/1 . . . . . . . . . . . . . . . . . . . . . . . . . 164
8.27.11 file property/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165
8.27.12 temporary name/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166
8.27.13 temporary file/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167
8.27.14 date time/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167
8.27.15 host name/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168
8.27.16 os version/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168
8.27.17 architecture/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168
8.27.18 shell/2, shell/1, shell/0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169
8.27.19 system/2, system/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169
8.27.20 spawn/3, spawn/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170
8.27.21 popen/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170
8.27.22 exec/5, exec/4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171
8.27.23 fork prolog/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172
8.27.24 create pipe/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172
8.27.25 wait/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173
8.27.26 prolog pid/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173
8.27.27 send signal/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174
8.27.28 sleep/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174
8.27.29 select/5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175
8.28 Sockets input/output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176
8.28.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176
8.28.2 socket/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176
8.28.3 socket close/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176
8.28.4 socket bind/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177
8.28.5 socket connect/4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178
8.28.6 socket listen/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178
8.28.7 socket accept/4, socket accept/3 . . . . . . . . . . . . . . . . . . . . . . . . . 179
8.28.8 hostname address/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179
8.29 Linedit management . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180
8.29.1 get linedit prompt/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180
8.29.2 set linedit prompt/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180
8.29.3 add linedit completion/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181
8.29.4 find linedit completion/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181
9 Finite domain solver and built-in predicates  183
9.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183
9.1.1 Finite Domain variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183
9.2 FD variable parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184
9.2.1 fd max integer/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184
9.2.2 fd vector max/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185
9.2.3 fd set vector max/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185
9.3 Initial value constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185
9.3.1 fd domain/3, fd domain bool/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185
9.3.2 fd domain/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186
9.4 Type testing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187

9.4.1 fd var/1, non fd var/1, generic var/1, non generic var/1 . . . . . . . . . . . 187

9.5 FD variable information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187
9.5.1 fd min/2, fd max/2, fd size/2, fd dom/2 . . . . . . . . . . . . . . . . . . . . . . 187

9.5.2 fd has extra cstr/1, fd has vector/1, fd use vector/1 . . . . . . . . . . . . . 188

9.6 Arithmetic constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188
9.6.1 FD arithmetic expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188

9.6.2 Partial AC: (#=)/2 - constraint equal, (#\=)/2 - constraint not equal,(#<)/2 - constraint less than, (#=<)/2 - constraint less than or equal,(#>)/2 - constraint greater than, (#>=)/2 - constraint greater than or equal189
9.6.3 Full AC: (#=#)/2 - constraint equal, (#\=#)/2 - constraint not equal,(#<#)/2 - constraint less than, (#=<#)/2 - constraint less than or equal,(#>#)/2 - constraint greater than, (#>=#)/2 - constraint greater than or equal190
9.6.4 fd prime/1, fd not prime/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191
9.7 Boolean and reified constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191
9.7.1 Boolean FD expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191
9.7.2 fd reified in/4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192
9.7.3 (#\)/1 - constraint NOT, (#<=>)/2 - constraint equivalent,(#\<=>)/2 - constraint different, (##)/2 - constraint XOR,(#==>)/2 - constraint imply, (#\==>)/2 - constraint not imply,(#/\)/2 - constraint AND, (#\/\)/2 - constraint NAND,(#\/)/2 - constraint OR, (#\\/)/2 - constraint NOR193

9.7.4 fd cardinality/2, fd cardinality/3, fd at least one/1, fd at most one/1,fd only one/1194

9.8 Symbolic constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195
9.8.1 fd all different/1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195
9.8.2 fd element/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195
9.8.3 fd element var/3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196

9.8.4 fd atmost/3, fd atleast/3, fd exactly/3 . . . . . . . . . . . . . . . . . . . . . 196

9.8.5 fd relation/2, fd relationc/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 197
9.9 Labeling constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198

9.9.1 fd labeling/2, fd labeling/1, fd labelingff/1 . . . . . . . . . . . . . . . . . 198

9.10 Optimization constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199
9.10.1 fd minimize/2, fd maximize/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199
10 Interfacing Prolog and C  201
10.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201
10.2 Including and using gprolog.h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201
10.3 Calling C from Prolog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202
10.3.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202
10.3.2 foreign/2 directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202
10.3.3 The C function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204
10.3.4 Input arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204
10.3.5 Output arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204
10.3.6 Input/output arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205
10.3.7 Writing non-deterministic C code . . . . . . . . . . . . . . . . . . . . . . . . . . . 205
10.3.8 Example: input and output arguments . . . . . . . . . . . . . . . . . . . . . . . . 206
10.3.9 Example: non-deterministic code . . . . . . . . . . . . . . . . . . . . . . . . . . . 207
10.3.10 Example: input/output arguments . . . . . . . . . . . . . . . . . . . . . . . . . . 209
10.4 Manipulating Prolog terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210
10.4.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210
10.4.2 Managing Prolog atoms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210
10.4.3 Reading Prolog terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211
10.4.4 Unifying Prolog terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212
10.4.5 Creating Prolog terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214
10.4.6 Testing the type of Prolog terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214
10.4.7 Comparing Prolog terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215
10.4.8 Term processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215
10.4.9 Comparing and evaluating arithmetic expressions . . . . . . . . . . . . . . . . . . 216
10.5 Raising Prolog errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217
10.5.1 Managing the error context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217
10.5.2 Instantiation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217
10.5.3 Uninstantiation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217
10.5.4 Type error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217
10.5.5 Domain error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217
10.5.6 Existence error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218
10.5.7 Permission error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218
10.5.8 Representation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218
10.5.9 Evaluation error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218
10.5.10 Resource error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

10.5.11 Syntax error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219
10.5.12 System error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219
10.6 Calling Prolog from C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220
10.6.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220
10.6.2 Example: my call/1 - a call/1 clone . . . . . . . . . . . . . . . . . . . . . . . . . 221
10.6.3 Example: recovering the list of all operators . . . . . . . . . . . . . . . . . . . . . 223
10.7 Defining a new C main() function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224
10.7.1 Example: asking for ancestors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225
References  229
Index  231

\underline{9}

1 Acknowledgements

I am grateful to the members of the Loco project at INRIA Rocquencourt^{2} for their encouragement. Their involvement in this work led to useful feedback and exchange (1995).

Many thanks to all those people at GNU^{3} who helped me to finalize the GNU Prolog project (1999).

I would like to thank everybody who tested preliminary releases and helped me to put the finishing touches to this system (1999).

I would particularly like to thank Jonathan Hodgson^{4} for the time and effort he put into the proofreading of this manual. His suggestions, both regarding ISO technical aspects as well as the language in which it was expressed, proved invaluable (1999-2017).

Thanks to Richard A. O’Keefe for his advice regarding the implementation of some Prolog built-in predicates and for suggesting me the in-place installation feature (1999).

The on-line HTML version of this document was created using H^{E}V^{E}A^{5} developed by Luc Maranget who kindly devoted so much of his time extending the capabilities of H^{E}V^{E}A in order to handle such a sizeable manual (2000).

Jean-Christophe Aude kindly improved the visual aspect of both the illustrations and the GNU Prolog web pages (2000).

Many thanks to the following contributors:

• Clive Cox^{6} and Edmund Grimley Evans^{7} for their port to ix86/SCO (1999).

• Andreas Stolcke^{8} for his port to ix86/Solaris (1999).

• Alexander Diemand^{9} for his initial port to alpha/linux (2000).

• Nicolas Ollinger^{10} to for his port to ix86/FreeBSD (2000).

• Brook Milligan^{11} for his port to ix86/NetBSD and for general configuration improvements (2000).

• Lindsey Spratt^{12} for his port to powerpc/Darwin (MacOS X) (2002).

• Gwenolé Beauchesne^{13} for his port to x86 64/Linux (2002).

• Jason Beegan^{14} for his port to sparc/NetBSD and to powerpc/NetBSD (2003).

• Cesar Rabak^{15} for his initial port to ix86/MinGW (2004).

• Scott L. Burson^{16} for his port to x86 64/Solaris (2007).

2http://www.inria.fr/Unites/ROCQUENCOURT-eng.html

3http://www.gnu.org

4http://www.sju.edu/~jhodgson

5http://pauillac.inria.fr/~maranget/hevea/

6clive@laluna.demon.co.uk

7http://www.rano.org/

8http://www.speech.sri.com/people/stolcke/

9ax@apax.net

10nollinge@ens-lyon.fr

11brook@nmsu.edu

12spratt@alum.mit.edu

13gbeauchesne@mandrakesoft.com

14jtb@netbsd.org

15csrabak@ig.com.br

16Scott@coral8.com

• David Holland^{17} for his port to x86 64/BSD systems (2008).

• Ozaki Kiichi^{18} for his port to x86 64/Darwin (Mac OS X) (2012).

• Jasper Taylor^{19} for his port to x86 64/MinGW64 (2011).

• Jasper Taylor^{20} for his ports to arm32/Linux and arm64/Linux (2021).

Many thanks to Paulo Moura for his continuous help (in particular about Darwin ports), for his ISO Prolog unit tests and for including GNU Prolog in his logtalk^{21} system (2000-).

Many thanks to John Collins, the latexmk^{22} maintainer, who greatly helped me to simplify the building of the documentation using latexmk (2021).

17dholland@netbsd.org

18gclient.gaap@gmail.com

19jasper@simulistics.com

20jasper@simulistics.com

21http://logtalk.org

22https://ctan.org/pkg/latexmk/

2 GNU Prolog License Conditions

GNU Prolog is free software. Since version 1.4.0, GNU Prolog distributed under a dual license: LGPL or GPL. So, you can redistribute it and/or modify it under the terms of either:

– the GNU Lesser General Public License (LGPL) as published by the Free Software Foundation; either version 3 of the License, or (at your option) any later version.

or

– the GNU General Public License (GPL) as published by the Free Software Foundation; either version 2 of the License, or (at your option) any later version.

or both in parallel (as here).

GNU Prolog is distributed in the hope that it will be useful, but WITHOUT ANY WARRANTY; without even the implied warranty of MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE. See the GNU General Public License for more details.

You should have received copies of the GNU General Public License and the GNU Lesser General Public License along with this program. If not, see http://www.gnu.org/licenses/^{23}.

Remark: versions of GNU Prolog prior to 1.4.0 were entirely released under the GNU General Public License (GPL).

3 Introduction

GNU Prolog [5] is a free Prolog compiler with constraint solving over finite domains. For recent information about GNU Prolog please consult the GNU Prolog page^{24}.

GNU Prolog is a Prolog compiler based on the Warren Abstract Machine (WAM) [9, 1]. It first compiles a Prolog program to a WAM file which is then translated to a low-level machine independent language called mini-assembly specifically designed for GNU Prolog. The resulting file is then translated to the assembly language of the target machine (from which an object is obtained). This allows GNU Prolog to produce a native stand alone executable from a Prolog source (similarly to what does a C compiler from a C program). The main advantage of this compilation scheme is to produce native code and to be fast. Another interesting feature is that executables are small. Indeed, the code of most unused built-in predicates is not included in the executables at link-time.

A lot of work has been devoted to the ISO compatibility. Indeed, GNU Prolog is very close to the ISO standard for Prolog [6].

GNU Prolog also offers various extensions very useful in practice (global variables, OS interface, sockets,...). In particular, GNU Prolog contains an efficient constraint solver over Finite Domains (FD). This opens constraint logic programming to the user combining the power of constraint programming to the declarativity of logic programming. The key feature of the GNU Prolog solver is the use of a single (lowlevel) primitive to define all (high-level) FD constraints. There are many advantages of this approach: constraints can be compiled, the user can define his own constraints (in terms of the primitive), the solver is open and extensible (as opposed to black-box solvers like CHIP),… Moreover, the GNU Prolog solver is rather efficient, often more than commercial solvers.

23http://www.gnu.org/licenses/

24http://www.gprolog.org

GNU Prolog is inspired from two systems:

• wamcc: a Prolog to C compiler [3]. the key point of wamcc was its ability to produce stand alone executables using an original compilation scheme: the translation of Prolog to C via the WAM. Its drawback was the time needed by gcc to compile the produced sources. GNU Prolog can also produce stand alone executables but using a faster compilation scheme.

• clp(FD): a constraint programming language over FD [4]. Its key feature was the use of a single primitive to define FD constraints. GNU Prolog is based on the same idea but offers an extended constraint definition language. In comparison to clp(FD), GNU Prolog offers new predefined constraints, new predefined heuristics, reified constraints,…

Here are some features of GNU Prolog:

• Prolog system:

– conforms to the ISO standard for Prolog (floating point numbers, streams, dynamic code,…).

– a lot of extensions: global variables, definite clause grammars (DCG), sockets interface, operating system interface,…

– more than 300 Prolog built-in predicates.

– Prolog debugger and a low-level WAM debugger.

– line editing facility under the interactive interpreter with completion on atoms.

– powerful bidirectional interface between Prolog and C.

• Compiler:

– native-code compiler producing stand alone executables.

– simple command-line compiler accepting a wide variety of files: Prolog files, C files, WAM files,…

– direct generation of assembly code 15 times faster than wamcc + gcc.

– most of unused built-in predicates are not linked (to reduce the size of the executables).

– compiled predicates (native-code) as fast as wamcc on average.

– consulted predicates (byte-code) 5 times faster than wamcc.

• Constraint solver:

– FD variables well integrated into the Prolog environment (full compatibility with Prolog variables and integers). No need for explicit FD declarations.

– very efficient FD solver (comparable to commercial solvers).

– high-level constraints can be described in terms of simple primitives.

– a lot of predefined constraints: arithmetic constraints, boolean constraints, symbolic constraints, reified constraints,…

– several predefined enumeration heuristics.

– the user can define his own new constraints.

– more than 50 FD built-in constraints/predicates.

4 Using GNU Prolog

4.1 Introduction

GNU Prolog offers two ways to execute a Prolog program:

• interpreting it using the GNU Prolog interactive interpreter.

• compiling it to a (machine-dependent) executable using the GNU Prolog native-code compiler.

Running a program under the interactive interpreter allows the user to list it and to make full use of the debugger on it (section 5, page 31). Compiling a program to native code makes it possible to obtain a stand alone executable, with a reduced size and optimized for speed. Running a Prolog program compiled to native-code is around 3-5 times faster than running it under the interpreter. However, it is not possible to make full use of the debugger on a program compiled to native-code. Nor is it possible to list the program. In general, it is preferable to run a program under the interpreter for debugging and then use the native-code compiler to produce an autonomous executable. It is also possible to combine these two modes by producing an executable that contains some parts of the program (e.g. already debugged predicates whose execution-time speed is crucial) and interpreting the other parts under this executable. In that case, the executable has the same facilities as the GNU Prolog interpreter but also integrates the native-code predicates. This way to define a new enriched interpreter is detailed later (section 4.4.5, page 27).

4.2 The GNU Prolog interactive interpreter

4.2.1 Starting/exiting the interactive interpreter

GNU Prolog offers a classical Prolog interactive interpreter also called top-level. It allows the user to execute queries, to consult Prolog programs, to list them, to execute them and to debug them. The top-level can be invoked using the following command:

% gprolog [OPTION ]…  (the % symbol is the operating system shell prompt)

Options:

--init-goal GOAL  execute GOAL before entering the top-level
--consult-file FILE  consult FILE inside the top-level
--entry-goal GOAL  execute GOAL inside the top-level
--query-goal GOAL  execute GOAL as a query for the top-level
--help  print a help and exit
--version  print version number and exit
--  do not parse the rest of the command-line

The main role of the gprolog command is to execute the top-level itself, i.e. to execute the built-in predicate top level/0 (section 8.18.1, page 126) which will produce something like:

     GNU Prolog 1.5.0 (64 bits)
     Compiled May  3 2021, 16:36:43 with gcc
     Copyright (C) 1999-2021 Daniel Diaz
     | ?-

The top-level is ready to execute your queries as explained in the next section.

To quit the top-level type the end-of-file key sequence (Ctl-D) or its term representation: end of file. It is also possible to use the built-in predicate halt/0 (section 8.18.1, page 126).

However, before entering the top-level itself, the command-line is processed to treat all known options (those listed above). All unrecognized arguments are collected together to form the argument list which will be available using argument value/2 (section 8.27.2, page 160) or argument list/1 (section 8.27.3, page 161). The -- option stops the parsing of the command-line, all remainding options are collected into the argument list.

Several options are provided to execute a goal before entering the interaction with the user:

• The --init-goal option executes the GOAL as soon as it is encountered (while the command-line is processed). GOAL is thus executed before entering top level/0.

• The --consult-file option consults the FILE at the entry of top level/0 just after the banner is displayed. --consult-file options are handled before --consult-file options.

• The --entry-goal option executes the GOAL at the entry of top level/0 just after the banner is displayed.

• The --query-goal option executes the GOAL as if the user has typed in (under the top-level).

The above order is thus the order in which each kind of goal (init, entry, query) is executed. If there are several goals of a same kind they are executed in the order of appearance. Thus, all init goals are executed (in the order of appearance) before all entry goals and all entry goals are executed before all query goals.

Each GOAL is passed as a shell argument (i.e. one shell string) and should not contain a terminal dot. Example: --init-goal 'write(hello), nl' under a sh-like. To be executed, a GOAL is transformed

into a term using read term from atom(Goal, Term, [end of term(eof)]). Respecting both the syn-

tax of shell strings and of Prolog can be heavy. For instance, passing a backslash character \ can be difficult since it introduces an escape sequence both in sh and inside Prolog quoted atoms. The use of back quotes can then be useful since, by default, no escape sequence is processed inside back quotes (this behavior can be controlled using the back quotes Prolog flag (section 8.22.1, page 148)).

Since the Prolog argument list is created when the whole command-line is parsed, if a --init-goal option uses argument value/2 or argument list/1 it will obtained the original command-line arguments (i.e. including all recognized arguments).

Here is an example of using execution goal options:

     % gprolog --init-goal 'write(before), nl' --entry-goal 'write(inside), nl'
     --query-goal 'append([a,b],[c,d],X)'

will produce the following:

     before
     GNU Prolog 1.5.0 (64 bits)
     Compiled May  3 2021, 16:36:43 with gcc
     Copyright (C) 1999-2021 Daniel Diaz
     inside
     | ?- append([a,b],[c,d],X).
     X = [a,b,c,d]
     yes
     | ?-

NB: depending on the used shell it may be necessary to use other string delimiters (e.g. use " under Windows cmd.exe).

4.2.2 The interactive interpreter read-execute-write loop

The GNU Prolog top-level is built on a classical read-execute-write loop that also allows for re-executions (when the query is not deterministic) as follows:

• display the prompt, i.e. ’| ?-’.

• read a query (i.e. a goal).

• execute the query.

• in case of success display the values of the variables of the query.

• if there are remaining alternatives (i.e. the query is not deterministic), display a ? and ask the user who can use one of the following commands: RETURN to stop the execution, ; to compute the next solution or a to compute all remaining solution.

Here is an example of execution of a query (“find the lists X and Y such that the concatenation of X and

Y is [a,b]”):
| ?- append(X,Y,[a,b,c]).
X = []

Y = [a,b,c] ?  ;  (here the user presses ; to compute another solution)

X = [a]

Y = [b,c] ?  a  (here the user presses a to compute all remaining solutions)

X = [a,b]

Y = [c]  (here the user is not asked and the next solution is computed)

X = [a,b,c]

Y = []  (here the user is not asked and the next solution is computed)
no  (no more solution)

In some cases the top-level can detect that the current solution is the last one (no more alternatives remaining). In such a case it does not display the ? symbol (and does not ask the user). Example:

| ?- (X=1 ; X=2).

X = 1 ?  ;  (here the user presses ; to compute another solution)
X = 2  (here the user is not prompted since there are no more alternatives)

yes

The user can stop the execution even if there are more alternatives by typing RETURN.

| ?- (X=1 ; X=2).

X = 1 ?  (here the user presses RETURN to stop the execution)

yes

The top-level tries to display the values of the variables of the query in a readable manner. For instance, when a variable is bound to a query variable, the name of this variable appears. When a variable is a singleton an underscore symbol is displayed ( is a generic name for a singleton variable, it is also called an anonymous variable). Other variables are bound to new brand variable names. When a query variable name X appears as the value of another query variable Y it is because X is itself not instantiated otherwise the value of X is displayed. In such a case, nothing is output for X itself (since it is a variable). Example:

     | ?- X=f(A,B, ,A), A=k.

A = k  (the value of A is displayed also in f/3 for X)
X = f(k,B, ,k)  (since B is a variable which is also a part of X, B is not displayed)

     | ?- functor(T,f,3), arg(1,T,X), arg(3,T,X).

T = f(X, ,X)  (the 1^{st} and 3^{rd} args are equal to X, the 2^{nd} is an anonymous variable)

     | ?- read from atom('k(X,Y,X).',T).

T = k(A, ,A)  (the 1^{st} and 3^{rd} args are unified, a new variable name A is introduced)

The top-level uses variable binding predicates (section 8.5, page 63). To display the value of a variable, the top-level calls write term/3 with the following option list: [quoted(true),numbervars(false), namevars(true)] (section 8.14.6, page 108). A term of the form '$VARNAME'(Name) where Name is an atom is displayed as a variable name while a term of the form '$VAR'(N) where N is an integer is displayed as a normal compound term (such a term could be output as a variable name by write term/3). Example:

     | ?- X='$VARNAME'('Y'), Y='$VAR'(1).

X = Y  (the term '$VARNAME'('Y') is displayed as Y)
Y = '$VAR'(1)  (the term '$VAR'(1) is displayed as is)

     | ?- X=Y, Y='$VAR'(1).
     X = '$VAR'(1)
     Y = '$VAR'(1)

In the first example, X is explicitly bound to '$VARNAME'('Y') by the query so the top-level displays Y as the value of X. Y is unified with '$VAR'(1) so the top-level displays it as a normal compound term. It should be clear that X is not bound to Y (whereas it is in the second query). This behavior should be kept in mind when doing variable binding operations.

Finally, the top-level computes the user-time (section 8.24.2, page 155) taken by a query and displays it when it is significant. Example:

     | ?- retractall(p( )), assertz(p(0)),
          repeat,
             retract(p(X)),
             Y is X + 1,
             assertz(p(Y)),
             X = 1000, !.
     X = 1000
     Y = 1001

(180 ms) yes  (the query took 180ms of user time)

4.2.3 Consulting a Prolog program

The top-level allows the user to consult Prolog source files. Consulted predicates can be listed, executed and debugged (while predicates compiled to native-code cannot). For more information about the difference between a native-code predicate and a consulted predicate refer to the introduction of this section (section 4.1, page 13) and to the part devoted to the compiler (section 4.4.1, page 21).

To consult a program use the built-in predicate consult/1 (section 8.23.1, page 153). The argument of this predicate is a Prolog file name or user to specify the terminal. This allows the user to directly input the predicates from the terminal. In that case the input shall be terminated by the end-of-file key sequence (Ctl-D) or its term representation: end of file. A shorthand for consult(FILE) is [FILE]. Example:

     | ?- [user].
     {compiling user for byte code...}
     even(0).
     even(s(s(X))):-
             even(X).

(here the user presses Ctl-D to end the input)

     {user compiled, 3 lines read - 350 bytes written, 1180 ms}
     | ?- even(X).

X = 0 ?  ;  (here the user presses ; to compute another solution)
X = s(s(0)) ?  ;  (here the user presses ; to compute another solution)
X = s(s(s(s(0)))) ? (here the user presses RETURN to stop the execution)

     yes
     | ?- listing.
     even(0).
     even(s(s(A))) :-
             even(A).

When consult/1 (section 8.23.1, page 153) is invoked on a Prolog file it first runs the GNU Prolog compiler (section 4.4, page 21) as a child process to generate a temporary WAM file for byte-code. If the compilation fails a message is displayed and nothing is loaded. If the compilation succeeds, the produced file is loaded into memory using load/1 (section 8.23.2, page 153). Namely, the byte-code of each predicate is loaded. When a predicate P is loaded if there is a previous definition for P it is removed (i.e. all clauses defining P are erased). We say that P is redefined. Note that only consulted predicates can be redefined. If P is a native-code predicate, trying to redefine it will produce an error at load-time: the predicate redefinition will be ignored and the following message displayed:

     native code procedure P cannot be redefined

Finally, an existing predicate will not be removed if it is not re-loaded. This means that if a predicate P is loaded when consulting the file F , and if later the definition of P is removed from the file F , consulting F again will not remove the previously loaded definition of P from the memory.

Consulted predicates can be debugged using the Prolog debugger. Use the debugger predicate trace/0 or debug/0 (section 5.3.1, page 31) to activate the debugger.

4.2.4 Scripting Prolog

Since version 1.4.0 it is possible to use a Prolog source file as a Unix script-file (shebang support). A PrologScript file should begin as follows:

     #!/usr/bin/gprolog --consult-file

GNU Prolog will be invoked as

     /usr/bin/gprolog --consult-file FILE

Then FILE will be consulted. In order to correctly deal with the #! first line, consult/1 treats as a comment a first line of a file which begins with # (if you want to use a predicate name starting with a #, simply skip a line before its definition).

Remark: it is almost never possible to pass additionnal parameters (e.g. query-goal) this way since in most systems the shebang implementation deliver all arguments (following #!/usr/bin/gprolog) as a single string (which cannot then correctly be recognized by gprolog).

4.2.5 Interrupting a query

Under the top-level it is possible to interrupt the execution of a query by typing the interruption key (Ctl-C). This can be used to abort a query, to stop an infinite loop, to activate the debugger,… When an interruption occurs the top-level displays the following message: Prolog interruption (h for help) ? The user can then type one of the following commands:

Command Name Description
a  abort abort the current execution. Same as abort/0 (section 8.18.1, page 126)
e exit quit the current Prolog process. Same as halt/0 (section 8.18.1, page 126)
b  break invoke a recursive top-level. Same as break/0 (section 8.18.1, page 126)
c  continue resume the execution
t  trace start the debugger using trace/0 (section 5.3.1, page 31)
d  debug start the debugger using debug/0 (section 5.3.1, page 31)
h or ?  help  display a summary of available commands

4.2.6 The line editor

The line editor (linedit) allows the user to build/update the current input line using a variety of commands. This facility is available if the linedit part of GNU Prolog has been installed. linedit is implicitly called by any built-in predicate reading from a terminal (e.g. get char/1, read/1,…). This is the case when the top-level reads a query.

Bindings: each command of linedit is activated using a key. For some commands another key is also available to invoke the command (on some terminals this other key may not work properly while the primary key always works). Here is the list of available commands:

Key  Alternate key Description
Ctl-B  ←  go to the previous character
Ctl-F  →  go to the next character
Esc-B       Ctl-←  go to the previous word
Esc-F       Ctl-→  go to the next word
Ctl-A        Home  go to the beginning of the line
Ctl-E         End  go to the end of the line
Ctl-H      Backspace delete the previous character
Ctl-D       Delete  delete the current character
Ctl-U      Ctl-Home  delete from beginning of the line to the current character
Ctl-K       Ctl-End  delete from the current character to the end of the line
Esc-L  lower case the next word
Esc-U  upper case the next word
Esc-C  capitalize the next word
Ctl-T  exchange last two characters
Ctl-V       Insert  switch on/off the insert/replace mode
Ctl-I         Tab  complete word (twice displays all possible completions)
Esc-Ctl-I     Esc-Tab  insert spaces to emulate a tabulation
Ctl-space  mark beginning of the selection
Esc-W  copy (from the begin selection mark to the current character)
Ctl-W  cut (from the begin selection mark to the current character)

Ctl-Y  paste

Ctl-P  ↑  recall previous history line
Ctl-N  ↓  recall next history line
Esc-P  recall previous history line beginning with the current prefix
Esc-N  recall next history line beginning with the current prefix
Esc-<       Page Up  recall first history line
Esc->      Page Down recall last history line
Ctl-C  generate an interrupt signal (section 4.2.5, page 18)
Ctl-D  generate an end-of-file character (at the begin of the line)
RETURN  validate a line
Esc-?  display a summary of available commands

History: when a line is entered (i.e. terminated by RETURN), linedit records it in an internal list called history. It is later possible to recall history lines using appropriate commands (e.g. Ctl-P recall the last entered line) and to modify them as needed. It is also possible to recall a history line beginning with a given prefix. For instance to recall the previous line beginning with write simply type write followed by Esc-P. Another Esc-P will recall an earlier line beginning with write,…

Completion: another important feature of linedit is its completion facility. Indeed, linedit maintains a list of known words and uses it to complete the prefix of a word. Initially this list contains all predefined atoms and the atoms corresponding to available predicates. This list is dynamically updated when a new atom appears in the system (whether read at the top-level, created with a built-in predicate, associated with a new consulted predicate,…). When the completion key (Tab) is pressed linedit acts as follows:

• use the current word as a prefix.

• collect all words of the list that begin with this prefix.

• complete the current word with the longest common part of all matching words.

• if more than one word matches emit a beep (a second Tab will display all possibilities).

Example:

| ?- argu  (here the user presses Tab to complete the word)

     | ?- argument  (linedit completes argu with argument and emits a beep)

(the user presses again Tab to see all possible completions)

argument counter  (linedit shows 3 possible completions)

     argument list
     argument value

| ?- argument  (linedit redisplays the input line)

     | ?- argument c  (to select argument counter the user presses c and Tab)
     | ?- argument counter (linedit completes with argument counter)

Balancing: linedit allows the user to check that (square/curly) brackets are well balanced. For this, when a close bracket symbol, i.e. ), ] or }, is typed, linedit determines the associated open bracket, i.e. (, [ or {, and temporarily repositions the cursor on it to show the match.

Customization: the behavior of linedit can be controlled via an environment variable called LINEDIT. This variable can contain the following substrings:

no  do not activated linedit (should the only value of the variable)
ansi=no  do not use ANSI escape sequence (unix only)
out=N  use the file descriptor N for the output (unix only)
gui=no  even if compiled with the GUI console run in text mode (windows)
gui=silent if the GUI console is not found, silently run in text mode (windows)
cp=N  use code page N (windows text console)
oem put=no do not use Char→Oem conversion when emitting a char (windows text console)
oem get=no do not use Oem→Char conversion when reading a char (windows text console)

4.3 Adjusting the size of Prolog data

GNU Prolog uses several stacks to execute a Prolog program. Each stack has a static size and cannot be dynamically increased during the execution. For each stack there is a default size but the user can define a new size by setting an environment variable. When a GNU Prolog program is run it first consults these variables and if they are not defined uses the default sizes. The following table presents each stack of GNU Prolog with its default size and the name of its associated environment variable:

Since version 1.4.2, the size of the atom table (the table recording all atoms) is managed similarly to stacks. It is then included in the following table (even if actually it is not a stack but an hash table). In this table, the associated name is atoms which is the key used in statistics (section 8.24.1, page 155). The environment variable name is derived from the corresponding Prolog flag max atom, see (section 8.22.1, page 148).

Stack  Default Environment Description
name size (Kb)  variable
local 16384 LOCALSZ control stack (environments and choice-points)
global 32768  GLOBALSZ heap (compound terms)
trail  16384  TRAILSZ  conditional bindings (bindings to undo at backtracking)
cstr  16384  CSTRSZ  finite domain constraint stack (FD variables and constraints)
atoms  32768  MAX ATOM  atom table

In addition, under Windows (since version 1.4.0), registry keys are consulted (key names are the same as environment names). The keys are stored in HKEY CURRENT USER\Software\GnuProlog\.

If the size of a stack is too small an overflow will occur during the execution. In that case GNU Prolog emits the following error message before stopping:

     S stack overflow (size: N Kb, environment variable used: E)

where S is the name of the stack, N is the current stack size in Kb and E the name of the associated environment variable. When such a message occurs it is possible to (re)define the variable E with the new size. For instance to allocate Kb to the local stack under a Unix shell use:

     LOCALSZ=32768; export LOCALSZ (under sh or bash)
     setenv LOCALSZ 32768  (under csh or tcsh)

This method allows the user to adjust the size of Prolog stacks. However, in some cases it is preferable not to allow the user to modify these sizes. For instance, when providing a stand alone executable whose behavior should be independent of the environment in which it is run. In that case the program should not consult environment variables and the programmer should be able to define new default stack sizes. The GNU Prolog compiler offers this facilities via several command-line options such as --local-size or --fixed-sizes (section 4.4.3, page 23).

Finally note that GNU Prolog stacks are virtually allocated (i.e. use virtual memory). This means that a physical memory page is allocated only when needed (i.e. when an attempt to read/write it occurs). Thus it is possible to define very large stacks. At the execution, only the needed amount of space will be physically allocated.

4.4 The GNU Prolog compiler

4.4.1 Different kinds of codes

One of the main advantages of GNU Prolog is its ability to produce stand alone executables. A Prolog program can be compiled to native code to give rise to a machine-dependent executable using the GNU Prolog compiler. However native-code predicates cannot be listed nor fully debugged. So there is an alternative to native-code compilation: byte-code compilation. By default the GNU Prolog compiler produces native-code but via a command-line option it can produce a file ready for byte-code loading. This is exactly what consult/1 does as was explained above (section 4.2.3, page 17). GNU Prolog also manages interpreted code using a Prolog interpreter written in Prolog. Obviously interpreted code is slower than byte-code but does not require the invocation of the GNU Prolog compiler. This interpreter is used each time a meta-call is needed as by call/1 (section 7.2.3, page 52). This also the case of dynamically asserted clauses. The following table summarizes these three kinds of codes:

Type  Speed  Debug ? For what
interpreted-code slow  yes  meta-call and dynamically asserted clauses
byte-code  medium  yes  consulted predicates
native-code  fast  no  compiled predicates

4.4.2 Compilation scheme

Native-code compilation: a Prolog source is compiled in several stages to produce an object file that is linked to the GNU Prolog libraries to produce an executable. The Prolog source is first compiled to obtain a WAM [9] file. For a detailed study of the WAM the interested reader can refer to “Warren’s Abstract Machine: A Tutorial Reconstruction”^{25} [1]. The WAM file is translated to a machine-independent language specifically designed for GNU Prolog. This language is close to a (universal) assembly language and is based on a very reduced instruction set. For this reason this language is called mini-assembly (MA). The mini-assembly file is then mapped to the assembly language of the target machine. This assembly file is assembled to give rise to an object file which is then linked with the GNU Prolog libraries to provide an executable. The compiler also takes into account Finite Domain constraint definition files. It translates them to C and invoke the C compiler to obtain object files. The following figure presents this compilation scheme:

25http://www.isg.sfu.ca/~hak/documents/wam.html

Obviously all intermediate stages are hidden to the user who simply invokes the compiler on his Prolog file(s) (plus other files: C,…) and obtains an executable. However, it is also possible to stop the compiler at any given stage. This can be useful, for instance, to see the WAM code produced (perhaps when learning the WAM). Finally it is possible to give any kind of file to the compiler which will insert it in the compilation chain at the stage corresponding to its type. The type of a file is determined using the suffix of its file name. The following table presents all recognized types/suffixes:

Suffix of the file  Type of the file  Handled by:

 .pl, .pro, .prolog  Prolog source file  pl2wam

.wam  WAM source file  wam2ma
.ma  Mini-assembly source file  ma2asm
.s  Assembly source file  the assembler
.c, .C, .CC, .cc, .cxx, .c++, .cpp C or C++ source file  the C compiler
.fd  Finite Domain constraint source file fd2c
any other suffix (.o, .a,…)  any other type (object, library,…) the linker (C linker)

Byte-code compilation: the same compiler can be used to compile a source Prolog file for byte-code. In that case the Prolog to WAM compiler is invoked using a specific option and produces a WAM for byte-code source file (suffixed .wbc) that can be later loaded using load/1 (section 8.23.2, page 153). Note that this is exactly what consult/1 (section 8.23.1, page 153) does as explained above (section 4.2.3, page 17).

4.4.3 Using the compiler

The GNU Prolog compiler is a command-line compiler similar in spirit to a Unix C compiler like gcc. To invoke the compiler use the gplc command as follows:

% gplc [OPTION ]… FILE …  (the % symbol is the operating system shell prompt)

The arguments of gplc are file names that are dispatched in the compilation scheme depending on the type determined from their suffix as was explained previously (section 4.4.2, page 21). All object files are then linked to produce an executable. Note however that GNU Prolog has no module facility (since there is not yet an ISO reference for Prolog modules) thus a predicate defined in a Prolog file is visible from any other predicate defined in any other file. GNU Prolog allows the user to split a big Prolog source into several files but does not offer any way to hide a predicate from others.

The simplest way to obtain an executable from a Prolog source file prog.pl is to use:

     % gplc prog.pl

This will produce an native executable called prog which can be executed as follows:

     % prog

However, there are several options that can be used to control the compilation:

General options:

-o FILE , --output FILE  use FILE as the name of the output file
-W, --wam-for-native  stop after producing WAM file(s)
-w, --wam-for-byte-code  stop after producing WAM for byte-code file(s) (force --no-call-c)
-M, --mini-assembly  stop after producing mini-assembly file(s)
-S, --assembly  stop after producing assembly file(s)
-F, --fd-to-c  stop after producing C file(s) from FD constraint definition file(s)
-c, --object  stop after producing object file(s)
--temp-dir PATH  use PATH as directory for temporary files
--no-del-temp  do not delete temporary files
--no-demangling  do not decode predicate names (name demangling)
-v, --verbose  print executed commands
-h, --help  print a help and exit
--version  print version number and exit

Prolog to WAM compiler options:

--pl-state FILE  read FILE to set the initial Prolog state

--wam-comment COMMENT  emit COMMENT as a comment in the WAM file

--no-susp-warn  do not show warnings for suspicious predicates
--no-singl-warn  do not show warnings for named singleton variables
--no-redef-error  do not show errors for built-in predicate redefinitions

--foreign-only  only compile foreign/1-2 directives

--no-call-c  do not allow the use of fd tell, '$call c’,…
--no-inline  do not inline predicates
--no-reorder  do not reorder predicate arguments
--no-reg-opt  do not optimize registers
--min-reg-opt  minimally optimize registers
--no-opt-last-subterm  do not optimize last subterm compilation
--fast-math  use fast mathematical mode (assume integer arithmetics)
--keep-void-inst  keep void WAM instructions in the output file
--compile-msg  print a compile message
--statistics  print statistics information

WAM to mini-assembly translator options:

--comment  include comments in the output file

Mini-assembly to assembly translator options:

--comment  include comments in the output file
--pic  produce position independent code (PIC)

C compiler options:

--c-compiler FILE  use FILE as C compiler/linker
-C OPTION  pass OPTION to the C compiler

Assembler options:

-A OPTION  pass OPTION to the assembler

Linker options:

 --linker FILE  use FILE as linker

--local-size N  set default local stack size to N Kb
--global-size N  set default global stack size to N Kb
--trail-size N  set default trail stack size to N Kb
--cstr-size N  set default constraint stack size to N Kb
--max-atom N  set default atom table size to N atoms
--fixed-sizes  do not consult environment variables at run-time (use default sizes)
--gui-console  link with the GUI console (windows only)
--new-top-level  link the top-level main (to recognize top-level command-line options)
--no-top-level  do not link the top-level (force --no-debugger)
--no-debugger  do not link the Prolog/WAM debugger
--min-pl-bips  link only used Prolog built-in predicates
--min-fd-bips  link only used FD solver built-in predicates

 --min-bips  shorthand for: --no-top-level --min-pl-bips --min-fd-bips
 --min-size  shorthand for: --min-bips --strip

--no-fd-lib  do not look for the FD library (maintenance only)
-s, --strip  strip the executable
-L OPTION  Pass OPTION to the linker

It is possible to only give the prefix of an option if there is no ambiguity.

The name of the output file is controlled via the -o FILE option. If present the output file produced will be named FILE . If not specified, the output file name depends on the last stage reached by the compiler. If the link is not done the output file name(s) is the input file name(s) with the suffix associated with the last stage. If the link is done, the name of the executable is the name (without suffix) of the first file name encountered in the command-line. Note that if the link is not done -o has no sense in the presence of multiple input file names. For this reason, several meta characters are available for substitution in

FILE :

• %f is substitued by the whole input file name.

• %F is similar to %f but the directory part is omitted.

• %p is substitued by the whole prefix file name (omitting the suffix).

• %P is similar to %p but the directory part is omitted.

• %s is substitued by the file suffix (including the dot).

• %d is substitued by the directory part (empty if no directory is specified).

• %c is substitued by the value of an internal counter starting from 1 and auto-incremented.

By default the compiler runs in the native-code compilation scheme. To generate a WAM file for byte-code use the --wam-for-byte-code option. The resulting file can then be loaded using load/1 (section 8.23.2, page 153).

To execute the Prolog to WAM compiler in a given read environment (operator definitions, character conversion table,…) use --pl-stateFILE . The state file should be produced by write pl state file/1 (section 8.22.5, page 152).

By default the Prolog to WAM compiler inlines calls to some deterministic built-in predicates (e.g. arg/3 and functor/3). Namely a call to such a predicate will not yield a classical predicate call but a simple C function call (which is obviously faster). It is possible to avoid this using --no-inline.

Another optimization performed by the Prolog to WAM compiler is unification reordering. The arguments of a predicate are reordered to optimize unification. This can be deactivated using --no-reorder. The compiler also optimizes the unification/loading of nested compound terms. More precisely, the compiler emits optimized instructions when the last subterm of a compound term is itself a compound term (e.g. lists). This can be deactivated using --no-opt-last-subterm.

By default the Prolog to WAM compiler fully optimizes the allocation of registers to decrease both the number of instruction produced and the number of used registers. A good allocation will generate many void instructions that are removed from the produced file except if --keep-void-inst is specified. To prevent any optimization use --no-reg-opt while --min-reg-opt forces the compiler to only perform simple register optimizations.

The Prolog to WAM compiler emits an error when a control construct or a built-in predicate is redefined. This can be avoided using --no-redef-error. The compiler also emits warnings for suspicious predicate definitions like -/2 since this often corresponds to an earlier syntax error (e.g. - instead of . This can be deactivated by specifying --no-susp-warn. Finally, the compiler warns when a singleton variable has a name (i.e. not the generic anonymous name ). This can be deactivated specifying --no-singl-warn.

Internally, predicate names are encoded to fit the syntax of (assembly) identifiers. For this GNU Prolog uses it own name mangling scheme. This is explained in more detail later (section 4.4.6, page 28). By default the error messages from the linker (e.g. multiple definitions for a given predicate, reference to an undefined predicate,…) are filtered to replace an internal name representation by the real predicate name (demangling). Specifying the --no-demangling prevents gplc from filtering linker output messages (internal identifiers are then shown).

When producing an executable it is possible to specify default stack sizes (using --STACK NAME-size) and to prevent it from consulting environment variables (using --fixed-sizes) as was explained above (section 4.3, page 20). By default the produced executable will include the top-level, the Prolog/WAM debugger and all Prolog and FD built-in predicates. It is possible to avoid linking the top-level (section 4.2, page 13) by specifying --no-top-level. In this case, at least one initialization/1 directive (section 7.1.14, page 50) should be defined. The option --no-debugger does not link the debugger. To include only used built-in predicates that are actually used the options --no-pl-bips and/or --no-fd-bips can be specified. For the smallest executable all these options should be specified. This can be abbreviated by using the shorthand option --min-bips. By default, executables are not stripped, i.e. their symbol table is not removed. This table is only useful for the C debugger (e.g. when interfacing Prolog and C). To remove the symbol table (and then to reduce the size of the final executable) use --strip. Finally --min-size is a shortcut for --min-bips and --strip, i.e. the produced executable is as small as possible.

Example: compile and link two Prolog sources prog1.pl and prog2.pl. The resulting executable will be named prog1 (since -o is not specified):

     % gplc prog1.pl prog2.pl

Example: compile the Prolog file prog.pl to study basic WAM code. The resulting file will be named

prog.wam:
     % gplc -W --no-inline --no-reorder --keep-void-inst prog.pl

Example: compile the Prolog file prog.pl and its C interface file utils.c to provide an autonomous executable called mycommand. The executable is not stripped to allow the use of the C debugger:

     % gplc -o mycommand prog.pl utils.c

Example: detail all steps to compile the Prolog file prog.pl (the resulting executable is stripped). All intermediate files are produced (prog.wam, prog.ma, prog.s, prog.o and the executable prog):

     % gplc -W prog.pl
     % gplc -M --comment prog.wam

     % gplc -S --comment prog.ma
     % gplc -c prog.s
     % gplc -o prog -s prog.o

4.4.4 Running an executable

In this section we explain what happens when running an executable produced by the GNU Prolog nativecode compiler. The default main function first starts the Prolog engine. This function collects all linked objects (issued from the compilation of Prolog files) and initializes them. The initialization of a Prolog object file consists in adding to appropriate tables new atoms, new predicates and executing its system directives. A system directive is generated by the Prolog to WAM compiler to reflect a (user) directive executed at compile-time such as op/3 (section 7.1.11, page 49). Indeed, when the compiler encounters such a directive it immediately executes it and also generates a system directive to execute it at the start of the executable. When all system directives have been executed the Prolog engine executes all initialization directives defined with initialization/1 (section 7.1.14, page 50). If several initialization directives appear in the same file they are executed in the order of appearance. If several initialization directives appear in different files the order in which they are executed is machine-dependant. However, on most machines the order will be the reverse order in which the associated files have been linked (this is not true under native win32). When all initialization directives have been executed the default main function looks for the GNU Prolog top-level. If present (i.e. it has been linked) it is called otherwise the program simply ends. Note that if the top-level is not linked and if there is no initialization directive the program is useless since it simply ends without doing any work. The default main function detects such a behavior and emits a warning message.

Example: compile an empty file prog.pl without linking the top-level and execute it:

     % gplc --no-top-level prog.pl
     % prog
     Warning: no initial goal executed
        use a directive :- initialization(Goal)
        or remove the link option --no-top-level (or --min-bips or --min-size)

4.4.5 Generating a new interactive interpreter

In this section we show how to define a new top-level extending the GNU Prolog interactive interpreter with new predicate definitions. The obtained top-level can then be considered as an enriched version of the basic GNU Prolog top-level (section 4.2, page 13). Indeed, each added predicate can be viewed as a predefined predicate just like any other built-in predicate. This can be achieved by compiling these predicates and including the top-level at link-time.

The real question is: why would we include some predicates in a new top-level instead of simply consulting them under the GNU Prolog top-level ? There are two reasons for this:

• the predicate cannot be consulted. This is the case of a predicate calling foreign code, like a predicate interfacing with C (section 10, page 201) or a predicate defining a new FD constraint.

• the performance of the predicate is crucial. Since it is compiled to native-code such a predicate will be executed very quickly. Consulting will load it as byte-code. The gain is much more noticeable if the program is run under the debugger. The included version will not be affected by the debugger while the consulted version will be several times slower. Obviously, a predicate should be included in a new top-level only when it is itself debugged since it is difficult to debug native-code.

To define a new top-level simply compile the set of desired predicates and linking them with the GNU Prolog top-level (this is the default) using gplc (section 4.4.3, page 23).

Example: let us define a new top-level called my top level including all predicates defined in prog.pl:

     % gplc -o my top level prog.pl

By the way, note that if prog.pl is an empty Prolog file the previous command will simply create a new interactive interpreter similar to the GNU Prolog top-level.

Example: as before where some predicates of prog.pl call C functions defined in utils.c:

     % gplc -o my top level prog.pl utils.c

To obtain a fully extended executable, it is desirable to accept the same set of opions as the original top-level, see (section 4.2, page 13), e.g. --init-goal. For this it is necessary to link main() function used by the original top-level. This can be achieved passing the --new-top-level to gplc:

     % gplc --new-top-level -o my top level prog.pl utils.c

In conclusion, defining a particular top-level is nothing else but a particular case of the native-code compilation. It is simple to do and very useful in practice.

4.4.6 The name mangling scheme

When the GNU Prolog compiler compiles a Prolog source to an object file it has to associate a symbol to each predicate name. However, the syntax of symbols is restricted to identifiers: string containing only letters, digits or underscore characters. On the other hand, predicate names (i.e. atoms) can contain any character with quotes if necessary (e.g. 'x+y=z' is a valid predicate name). The compiler may thus have to encode predicate names respecting the syntax of identifiers. In addition, Prolog allows the user to define several predicates with the same name and different arities, for this GNU Prolog encodes predicate indicators (predicate name followed by the arity). Finally, to support modules in the future, the module name is also encoded.

Since version 1.4.0, GNU Prolog adopts the following name mangling scheme. A predicate indicator of the form [MODULE:]PRED/N (where the MODULE can be omitted) will give rise to an identifier of the following form: XK [E(MODULE) ]E(PRED) aN where:

K is a digit in 0..5 storing coding information about MODULE and PRED . Possible values are: • 0: no module present, PRED is not encoded

• 1: no module present, PRED is encoded

• 2: MODULE is not encoded, PRED is not encoded

• 3: MODULE is not encoded, PRED is encoded

• 4: MODULE is encoded, PRED is not encoded

• 5: MODULE is encoded, PRED is encoded

E(STR) is a function to encode a string STR which returns:

• STR itself (not encoded) if STR only contains letters, digits or but does not contain the substring and does not begin nor end with (i.e. regexp: [a-zA-Z0-9]([-]?[a-zA-Z0-9])*).

• an hexadecimal representation of each character of the string otherwise. For example: E( x+y=z) returns 782B793D7A since 78 is the hexadecimal representation of the ASCII code of x, 2B of the code of +, etc.

Examples:

Predicate indicator internal identifier

                     father/2          X0 father a2
                     'x+y=z'/3         X1 782B793D7A a3
                     util:same/2       X2 util same a2
                     util:same 1/3     X3 util 73616D655F5F31 a3

So, from the mini-assembly stage, each predicate indicator is handled via its name mangling identifier. The knowledge of this scheme is normally not of interest for the user, i.e. the Prolog programmer. For this reason the GNU Prolog compiler hides this mangling. When an error occurs on a predicate (undefined predicate, predicate with multiple definitions,…) the compiler has to decode the symbol associated with the predicate indicator (name demangling). For this gplc filters each message emitted by the linker to locate and decode eventual predicate indicators. This filtering can be deactivated specifying --no-demangling when invoking gplc (section 4.4.3, page 23).

This filter is provided as an utility that can be invoked using the hexgplc command as follows:

% hexgplc [OPTION ]… FILE … (the % symbol is the operating system shell prompt)

Options:

--decode or --demangling decoding mode (this is the default mode)

 --encode or --mangling  encoding mode

--relax  decode also predicate names (not only predicate indicators)
--printf FORMAT  pass encoded/decoded string to C printf(3) with FORMAT
--aux-father  decode an auxiliary predicate as its father
--aux-father2  decode an auxiliary predicate as its father + auxiliary number
--cmd-line  encode/decode each argument of the command-line

 -E or -M  same as: --cmd-line --encode --relax
 -P or -D  same as: --cmd-line --decode --relax --quote

--help  print a help and exit
--version  print version number and exit

It is possible to give a prefix of an option if there is no ambiguity.

Without arguments hexgplc runs in decoding mode reading its standard input and decoding (demangling) each symbol corresponding to a predicate indicator. To use hexgplc in the encoding (mangling) mode the --encode option must be specified. By default hexgplc only decodes predicate indicators, this can be relaxed using --relax to also take into account simple predicate names (the arity can be omitted). It is possible to format the output of an encoded/decoded string using --printf FORMAT in that case each string S is passed to the C printf(3) function as printf(FORMAT,S).

Auxiliary predicates are generated by the Prolog to WAM compiler when simplifying some control constructs like ';'/2 present in the body of a clause. They are of the form '$NAME/ARITY $auxN' where NAME/ARITY is the predicate indicator of the simplified (i.e. father) predicate and N is a sequential number (a predicate can give rise to several auxiliary predicates). It is possible to force hexgplc to decode an auxiliary predicate as its father predicate indicator using --aux-father or as its father predicate indicator followed by the sequential number using --aux-father2.

If no file is specified, hexgplc processes its standard input otherwise each file is treated sequentially. Specifying the --cmd-line option informs hexgplc that each argument is not a file name but a string that must be encoded (or decoded). This is useful to encode/decode a particular string. For this reason the option -E (encode) and -D (decode) are provided as shorthand. Then, to obtain the mangling representation of a predicate PRED use:

     % hexgplc -E PRED

NB: if PRED is a complex atom it is necessary to quote it (the quotes must be passed to hexgplc). Here is an example under bash:

     % hexgplc -E \'x+y=z\'/3
     X1_782B793D7A__a3

Or even more safely (using bash quotes to prevent bash from interpreting special characters):

     % hexgplc -E \''x+y=z'\'/3
     X1_782B793D7A__a3

5 Debugging

5.1 Introduction

The GNU Prolog debugger provides information concerning the control flow of the program. The debugger can be fully used on consulted predicates (i.e. byte-code). For native compiled code only the calls/exits are traced, no internal behavior is shown. Under the debugger it is possible to exhaustively trace the execution or to set spy-points to only debug a specific part of the program. Spy-points allow the user to indicate on which predicates the debugger has to stop to allow the user to interact with it. The debugger uses the “procedure box control flow model”, also called the Byrd Box model since it is due to Lawrence Byrd.

5.2 The procedure box model

The procedure box model of Prolog execution provides a simple way to show the control flow. This model is very popular and has been adopted in many Prolog systems (e.g. SICStus Prolog, Quintus Prolog,…). A good introduction is the chapter 8 of “Programming in Prolog” of Clocksin & Mellish [2]. The debugger executes a program step by step tracing an invocation to a predicate (call) and the return from this predicate due to either a success (exit) or a failure (fail). When a failure occurs the execution backtracks to the last predicate with an alternative clause. The predicate is then reinvoked (redo). Another source of change of the control flow is due to exceptions. When an exception is raised from a predicate (exception) by throw/1 (section 7.2.4, page 53) the control is given back to the most recent predicate that has defined a handler to recover this exception using catch/3 (section 7.2.4, page 53). The procedure box model shows these different changes in the control flow, as illustrated here:

Each arrow corresponds to a port. An arrow to the box indicates that the control is given to this predicate while an arrow from the box indicates that the control is given back from the procedure. This model visualizes the control flow through these five ports and the connections between the boxes associated with subgoals. Finally, it should be clear that a box is associated with one invocation of a given predicate. In particular, a recursive predicate will give raise to a box for each invocation of the predicate with different entries/exits in the control flow. Since this might get confusing for the user, the debugger associates with each box a unique identifier (i.e. the invocation number).

5.3 Debugging predicates

5.3.1 Running and stopping the debugger

trace/0 activates the debugger. The next invocation of a predicate will be traced.

debug/0 activates the debugger. The next invocation of a predicate on which a spy-point has been set will be traced.

It is important to understand that the information associated with the control flow is only available when the debugger is on. For efficiency reasons, when the debugger is off the information concerning the control flow (i.e. the boxes) is not retained. So, if the debugger is activated in the middle of a computation (by a call to debug/0 or trace/0 in the program or after the interrupt key sequence (Ctl-C) by choosing trace or debug), information prior to this point is not available.

debugging/0: prints onto the terminal information about the current debugging state (whether the debugger is switched on, what are the leashed ports, spy-points defined,…).

notrace/0 or nodebug/0 switches the debugger off.

wam debug/0 invokes the sub-debugger devoted to the WAM data structures (section 5.6, page 35). It can be also invoked using the W debugger command (section 5.5, page 33).

5.3.2 Leashing ports

leash(Ports) requests the debugger to prompt the user, as he creeps through the program, for every port defined in the Ports list. Each element of Ports is an atom in call, exit, redo, fail, exception. Ports can also be an atom defining a shorthand:

   • full: equivalent to [call, exit, redo, fail, exception]
   • half: equivalent to [call, redo]

• loose: equivalent to [call]

• none: equivalent to []

   • tight: equivalent to [call, redo, fail, exception]

When an unleashed port is encountered the debugger continues to show the associated goal but does not stop the execution to prompt the user.

5.3.3 Spy-points

When dealing with big sources it is not very practical to creep through the entire program. It is preferable to define a set of spy-points on interesting predicates to be prompted when the debugger reaches one of these predicates. Spy-points can be added either using spy/1 (or spypoint condition/3) or dynamically when prompted by the debugger using the + (or *) debugger command (section 5.5, page 33). The current mode of leashing does not affect spy-points in the sense that user interaction is requested on every port.

spy(PredSpec) sets a spy-point on all the predicates given by PredSpec. PredSpec defines one or several predicates and has one of the following forms:

• [PredSpec1, PredSpec2,...]: set a spy-point for each element of the list.

• Name: set a spy-point for any predicate whose name is Name (whatever the arity).

• Name/Arity: set a spy-point for the predicate whose name is Name and arity is Arity.

• Name/A1-A2: set a spy-point for the each predicate whose name is Name and arity is between A1 and A2.

It is not possible to set a spy-point on an undefined predicate.

The following predicate is used to remove one or several spy-points:

nospy(PredSpec) removes the spy-points from the specified predicates.

nospyall/0 removes all spy-points:

It is also possible to define conditional spy-points.

spypoint condition(Goal, Port, Test) sets a conditional spy-point on the predicate for Goal. When the debugger reaches a conditional spy-point it only shows the associated goal if the following conditions are verified:

• the actual goal unifies with Goal.

• the actual port unifies with Port.

• the Prolog goal Test succeeds.

5.4 Debugging messages

We here described which information is displayed by the debugger when it shows a goal. The basic format is as follows:

     S N M Port: Goal ?

S is a spy-point indicator: if there is a spy-point on the current goal the + symbol is displayed else a space is displayed. N is the invocation number. This unique number can be used to correlate the trace messages for the various ports, since it is unique for every invocation. M is an index number which represents the number of direct ancestors of the goal (i.e. the current depth of the goal). Port specifies the particular port (call, exit, fail, redo, exception). Goal is the current goal (it is then possible to inspect its current instantiation) which is displayed using write term/3 with quoted(true) and max depth(D) options (section 8.14.6, page 108). Initially D (the print depth) is set to 10 but can be redefined using the < debugger command (section 5.5, page 33). The ? symbol is displayed when the debugger is waiting a command from the user. (i.e. Port is a leashed port). If the port is unleashed, this symbol is not displayed and the debugger continues the execution displaying the next goal.

5.5 Debugger commands

When the debugger reaches a leashed port it shows the current goal followed by the ? symbol. At this point there are many commands available. Typing RETURN will creep into the program. Continuing to creep will show all the control flow. The debugger shows every port for every predicate encountered during the execution. It is possible to select the ports at which the debugger will prompt the user using the built-in predicate leash/1 (section 5.3.2, page 32). Each command is only one character long:

Command  Name  Description
RET or c  creep  single-step to the next port
l  leap  continue the execution only stopping when a goal with a spy-point

is reached

s skip skip over the entire execution of the current goal. No message will

be shown until control returns

G go to ask for an invocation number and continue the execution until a port

is reached for that invocation number

r retry try to restart the invocation of the current goal by failing until reaching the invocation of the goal. The state of execution is the same as when the goal was initially invoked (except when using side-effect predicates)

f  fail  force the current goal to fail immediately
w  write  show the current goal using write/2 (section 8.14.6, page 108)
d  display  show the current goal using display/2 (section 8.14.6, page 108)
p  print  show the current goal using print/2 (section 8.14.6, page 108)
e  exception  show the pending exception. Only applicable to an exception port
g  ancestors  show the list of ancestors of the current goal
A  alternatives  show the list of ancestors of the current goal combined with choice-

points

u unify ask for a term and unify the current goal with this term. This is convenient for getting a specific solution. Only available at a call port

. father file show the Prolog file name and the line number where the current

predicate is defined

n  no debug  switch the debugger off. Same as nodebug/0 (section 5.3.1, page 31)
=  debugging  show debugger information. Same as debugging/0 (section 5.3.1,

page 31)

+ spy this set a spy-point on the current goal. Uses spy/1 (section 5.3.3,

page 32)

- nospy this remove a spy-point on the current goal. Uses nospy/1 (section 5.3.3, page 32)

* spy conditionally ask for a term Goal, Port, Test (terminated by a dot) and set a conditional spy-point on the current predicate. Goal and the current goal must have the same predicate indicator. Uses spypoint condition/3 (section 5.3.3, page 32)

L listing list all the clauses associated with the current predicate. Uses

listing/1 (section 8.23.3, page 154)

a abort abort the current execution. Same as abort/0 (section 8.18.1,

page 126)

b break invoke a recursive top-level. Same as break/0 (section 8.18.1,

page 126)

@ execute goal ask for a goal and execute it < set print depth ask for an integer and set the print depth to this value (-1 for no

depth limit)

h or ? help display a summary of available commands W WAM debugger invoke the low-level WAM debugger (section 5.6, page 35)

5.6 The WAM debugger

In some cases it is interesting to have access to the WAM data structures. This sub-debugger allows the user to inspect/modify the contents of any stack or register of the WAM. The WAM debugger is invoked using the built-in predicate wam debug/0 (section 5.3.1, page 31) or the W debugger command (section 5.5, page 33). The following table presents the specific commands of the WAM debugger:

Command  Description
write A [N ]  write N terms starting at the address A using write/1 (section 8.14.6, page 108)
data A [N ]  display N words starting at the address A
modify A [N ]  display and modify N words starting at the address A
where A  display the real address corresponding to A
what RA  display what corresponds to the real address RA
deref A  display the dereferenced word starting at the address A
envir [SA ]  display the contents of the environment located at SA (or the current one)
backtrack [SA ] display the contents of the choice-point located at SA (or the current one)
backtrack all display all choice-points
quit  quit the WAM debugger
help  display a summary of available commands

In the above table the following conventions apply:

• elements between [ and ] are optional.

• N is an optional integer (defaults to 1).

• A is a WAM address, its syntax is: BANK NAME [ [ N ] ], i.e. a bank name possibly followed by an

index (defaults to 0). BANK NAME is either: – reg: WAM general register (stack pointers, continuation, ...).

– x: WAM X register (temporary variables, i.e. arguments).

– y: WAM Y register (permanent variables).

– ab: WAM X register saved in the current choice-point.

      – STACK NAME : WAM stack (STACK NAME in local, global, trail, cstr).

• SA is a WAM stack address, i.e. STACK NAME [ [ N ] ] (special case of WAM addresses).

• RA is a real address, its syntax is the syntax of C integers (in particular the notation 0x... is recognized).

It is possible to only use the first letters of a commands and bank names when there is no ambiguity. Also the square brackets [ ] enclosing the index of a bank name can be omitted. For instance the following command (showing the contents of 25 consecutive words of the global stack from the index 3): data global[3] 25 can be abbreviated as: d g 3 25.

36  5 DEBUGGING

\underline{37}

6 Format of definitions

6.1 General format

The definition of control constructs, directives and built-in predicates is presented as follows: Templates Specifies the types of the arguments and which of them shall be instantiated (mode). Types and modes are described later (section 6.2, page 37). Description Describes the behavior (in the absence of any error conditions). It is explicitly mentioned when a builtin predicate is re-executable on backtracking. Predefined operators involved in the definition are also mentioned. Errors Details the error conditions. Possible errors are detailed later (section 6.3, page 39). For directives, this part is omitted. Portability Specifies whether the definition conforms to the ISO standard or is a GNU Prolog extension.

6.2 Types and modes

The templates part defines, for each argument of the concerned built-in predicate, its mode and type. The mode specifies whether or not the argument must be instantiated when the built-in predicate is called. The mode is encoded with a symbol just before the type. Possible modes are:

• +: the argument must be instantiated.

• -: the argument must be a variable (will be instantiated if the built-in predicate succeeds).

• ?: the argument can be instantiated or a variable. The type of an argument is defined by the following table:

Type  Description

TYPE list a list whose the type of each element is TYPE TYPE1 or TYPE2 a term whose type is either TYPE1 or TYPE2

atom                 an atom

atom property an atom property (section 8.19.11, page 134)

boolean  the atom true or false

byte an integer ≥ 0 and ≤ 255 callable term an atom or a compound term character a single character atom character code an integer ≥ 1 and ≤ 255 clause a clause (fact or rule) close option a close option (section 8.10.7, page 82) compound term a compound term evaluable an arithmetic expression (section 8.6.1, page 66) fd bool evaluable a boolean FD expression (section 9.7.1, page 191) fd labeling option an FD labeling option (section 9.9.1, page 198) fd evaluable an arithmetic FD expression (section 9.6.1, page 188) fd variable an FD variable flag a Prolog flag (section 8.22.1, page 148) float a floating point number head a head of a clause (atom or compound term)

integer              an integer

in byte an integer ≥ 0 and ≤ 255 or -1 (for the end-of-file) in character a single character atom or the atom end of file (for the end-of-file) in character code an integer ≥ 1 and ≤ 255 or -1 (for the end-of-file)

io mode  an atom in: read, write or append

list the empty list [] or a non-empty list [ | ] nonvar any term that is not a variable number an integer or a floating point number operator specifier an operator specifier (section 8.14.10, page 113) os file property an operating system file property (section 8.27.11, page 165)

predicate indicator a term Name/Arity where Name is an atom and Arity an integer ≥ 0. A

callable term can be given if the strict iso Prolog flag is switched off (section 8.22.1, page 148)

predicate property a predicate property (section 8.8.2, page 75) read option a read option (section 8.14.1, page 104)

socket address  a term of the form 'AF UNIX'(A) or 'AF INET'(A,N) where A is an atom

and N an integer

socket domain  an atom in: 'AF UNIX' or 'AF INET'

source sink an atom identifying a source or a sink stream a stream-term: a term of the form '$stream'(N) where N is an integer ≥ 0 stream option a stream option (section 8.10.6, page 81) stream or alias a stream-term or an alias (atom)

stream position  a stream position: a term '$stream position'(I1, I2, I3, I4) where

I1, I2, I3 and I4 are integers

stream property a stream property (section 8.10.10, page 84)

stream seek method an atom in: bof, current or eof

term                 any term

var binding option a variable binding option (section 8.5.3, page 64) write option a write option (section 8.14.6, page 108)

\underline{6.3 Errors 39}

6.3 Errors

6.3.1 General format and error context

When an error occurs an exception of the form: error(ErrorTerm, Caller) is raised. ErrorTerm is a term specifying the error (detailed in next sections) and Caller is a term specifying the context of the error. The context is either the predicate indicator of the last invoked built-in predicate or an atom giving general context information.

Using exceptions allows the user both to recover an error using catch/3 (section 7.2.4, page 53) and to raise an error using throw/1 (section 7.2.4, page 53).

To illustrate how to write error cases, let us write a predicate my pred(X) where X must be an integer:

     my_pred(X) :-
             (   nonvar(X) ->
                 true
             ;   throw(error(instantiation_error, my_pred/1)),
             ),
             (   integer(X) ->
                 true
             ;   throw(error(type_error(integer, X), my_pred/1))
             ),
             ...

To help the user to write these error cases, a set of system predicates is provided to raise errors. These predicates are of the form '$pl err ...' and they all refer to the implicit error context. The predicates set bip name/2 (section 8.22.3, page 151) and current bip name/2 (section 8.22.4, page 151) are provided to set and recover the name and the arity associated with this context (an arity < 0 means that only the atom corresponding to the functor is significant). Using these system predicates the user could define the above predicate as follow:

     my_pred(X) :-
             set_bip_name(my_pred,1),
             (   nonvar(X) ->
                 true
             ;   '$pl_err_instantiation'
             ),
             (   integer(X) ->
                 true
             ;   '$pl_err_type'(integer, X)
             ),
             ...

The following sections detail each kind of errors (and associated system predicates).

6.3.2 Instantiation error

An instantiation error occurs when an argument or one of its components is variable while an instantiated argument was expected. ErrorTerm has the following form: instantiation error.

The system predicate '$pl err instantiation' raises this error in the current error context (section 6.3.1, page 39).

6.3.3 Uninstantiation error

An uninstantiation Error when an argument or one of its components is not a variable, and a variable or a component as variable is required. ErrorTerm has the following form: uninstantiation error(Culprit) where Culprit is the argument or one of its components which caused the error. The system predicate '$pl err uninstantiation'(Culprit) raises this error in the current error context (section 6.3.1, page 39).

6.3.4 Type error

A type error occurs when the type of an argument or one of its components is not the expected type (but not a variable). ErrorTerm has the following form: type error(Type, Culprit) where Type is the expected type and Culprit the argument which caused the error. Type is one of:

   • atom
   • atomic
   • boolean
   • byte
   • callable
   • character
   • compound

                                • evaluable
                                • fd bool evaluable
                                • fd evaluable
                                • fd variable
                                • float
                                • in byte
                                • in character

                                                              • integer
                                                              • list
                                                              • number
                                                              • pair
                                                              • predicate indicator

The system predicate '$pl err type'(Type, Culprit) raises this error in the current error context (section 6.3.1, page 39).

6.3.5 Domain error

A domain error occurs when the type of an argument is correct but its value is outside the expected

domain. ErrorTerm has the following form: domain error(Domain, Culprit) where Domain is the

expected domain and Culprit the argument which caused the error. Domain is one of:

   • atom property
   • buffering mode
   • character code list
   • close option
   • date time
   • eof action
   • fd labeling option
   • flag value
   • format control sequence

                                • g array index
                                • io mode
                                • non empty list
                                • not less than zero
                                • operator priority
                                • operator specifier
                                • order
                                • os file permission
                                • os file property

                                                              • os path
                                                              • predicate property
                                                              • prolog flag
                                                              • read option
                                                              • selectable item
                                                              • socket address
                                                              • socket domain
                                                              • source sink
                                                              • statistics key

   • statistics value
   • stream
   • stream option
   • stream or alias

                                • stream position
                                • stream property
                                • stream seek method
                                • stream type

                                                              • term stream or alias
                                                              • var binding option
                                                              • write option

The system predicate '$pl err domain'(Domain, Culprit) raises this error in the current error context (section 6.3.1, page 39).

6.3.6 Existence error

an existence error occurs when an object on which an operation is to be performed does not exist.

ErrorTerm has the following form: existence error(Object, Culprit) where Object is the type of

the object and Culprit the argument which caused the error. Object is one of:

   • procedure  • source sink  • stream

The system predicate '$pl err existence'(Object, Culprit) raises this error in the current error context (section 6.3.1, page 39).

6.3.7 Permission error

A permission error occurs when an attempt to perform a prohibited operation is made. ErrorTerm

has the following form: permission error(Operation, Permission, Culprit) where Operation is

the operation which caused the error, Permission the type of the tried permission and Culprit the argument which caused the error. Operation is one of:

   • access • add alias • close

                                • create • input • modify

                                                              • open
                                                              • output
                                                              • reposition

and Permission is one of:

   • binary stream
   • flag
   • operator

                                • past end of stream
                                • private procedure
                                • source sink

                                                              • static procedure
                                                              • stream
                                                              • text stream
The system predicate '$pl err permission'(Operation, Permission, Culprit) raises this error in

the current error context (section 6.3.1, page 39).

6.3.8 Representation error

A representation error occurs when an implementation limit has been breached. ErrorTerm has the following form: representation error(Limit) where Limit is the name of the reached limit. Limit is one of:

   • character
   • character code
   • in character code

                                • max arity
                                • max integer
                                • min integer

                                                              • integer 32bits
                                                              • too many variables

The errors max integer and min integer are not currently implemented.

The system predicate '$pl err representation'(Limit) raises this error in the current error context (section 6.3.1, page 39).

6.3.9 Evaluation error

An evaluation error occurs when an arithmetic expression gives rise to an exceptional value. ErrorTerm has the following form: evaluation error(Error) where Error is the name of the error. Error is one of:

   • float overflow
   • int overflow

                                • undefined • underflow

                                                              • zero divisor

The errors float overflow, int overflow, undefined and underflow are not currently implemented.

The system predicate '$pl err evaluation'(Error) raises this error in the current error context (section 6.3.1, page 39).

6.3.10 Resource error

A resource error occurs when GNU Prolog does not have enough resources. ErrorTerm has the following form: resource error(Resource) where Resource is the name of the resource. Resource is one of:

   • print object not linked  • too big fd constraint

The system predicate '$pl err resource'(Resource) raises this error in the current error context (section 6.3.1, page 39).

6.3.11 Syntax error

A syntax error occurs when a sequence of character does not conform to the syntax of terms. ErrorTerm has the following form: syntax error(Error) where Error is an atom explaining the error.

The system predicate '$pl err syntax'(Error) raises this error in the current error context (section 6.3.1, page 39).

6.3.12 System error

A system error can occur at any stage. A system error is generally associated with an external component (e.g. operating system). ErrorTerm has the following form: system error(Error) where Error is an atom explaining the error. This is an extension to ISO which only defines system error without arguments.

The system predicate '$pl err system'(Error) raises this error in the current error context (section 6.3.1, page 39).

44  6 FORMAT OF DEFINITIONS

7 Prolog directives and control constructs

7.1 Prolog directives

7.1.1 Introduction

Prolog directives are annotations inserted in Prolog source files for the compiler. A Prolog directive is used to specify:

• the properties of some procedures defined in the source file.

• the format and the syntax for read-terms in the source file (using changeable Prolog flags).

• included source files.

• a goal to be executed at run-time.

7.1.2 dynamic/1

Templates

     dynamic(+predicate indicator)
     dynamic(+predicate indicator list)
     dynamic(+predicate indicator sequence)

Description

dynamic(Pred) specifies that the procedure whose predicate indicator is Pred is a dynamic procedure. This directive makes it possible to alter the definition of Pred by adding or removing clauses. For more information refer to the section about dynamic clause management (section 8.7.1, page 70).

This directive shall precede the definition of Pred in the source file.

If there is no clause for Pred in the source file, Pred exists however as an empty predicate (this means

that current predicate(Pred) succeeds).

In order to allow multiple definitions, Pred can also be a list of predicate indicators or a sequence of predicate indicators using ','/2 as separator.

Portability

ISO directive.

7.1.3 public/1

Templates

     public(+predicate indicator)
     public(+predicate indicator list)
     public(+predicate indicator sequence)

Description

public(Pred) specifies that the procedure whose predicate indicator is Pred is a public procedure. This directive makes it possible to inspect the clauses of Pred. For more information refer to the section about dynamic clause management (section 8.7.1, page 70).

This directive shall precede the definition of Pred in the source file. Since a dynamic procedure is also public. It is useless (but correct) to define a public directive for a predicate already declared as dynamic.

In order to allow multiple definitions, Pred can also be a list of predicate indicators or a sequence of predicate indicators using ','/2 as separator.

Portability

GNU Prolog directive. The ISO reference does not define any directive to declare a predicate public but it does distinguish public predicates. It is worth noting that in most Prolog systems the public/1 directive is as a visibility declaration. Indeed, declaring a predicate as public makes it visible from any predicate defined in any other file (otherwise the predicate is only visible from predicates defined in the same source file as itself). When a module system is incorporated in GNU Prolog a more general visibility declaration shall be provided conforming to the ISO reference.

7.1.4 multifile/1

Templates

     multifile(+predicate indicator)
     multifile(+predicate indicator list)
     multifile(+predicate indicator sequence)

Description

multifile(Pred) specifies that the procedure whose predicate indicator is Pred is a multifle procedure (the clauses of Pred can reside in several source files). This directive is only supported by GNU Prolog since version 1.4.0.

The native compilation scheme of GNU Prolog requires that each Prolog source file refering to a multifile predicate Pred must include a multifile(Pred) directive even if no clause are defined in this file for Pred (i.e. Pred is only called by other predicates in this source file).

Portability

ISO directive.

7.1.5 discontiguous/1

Templates

     discontiguous(+predicate indicator)
     discontiguous(+predicate indicator list)
     discontiguous(+predicate indicator sequence)

discontiguous(Pred) specifies that the procedure whose predicate indicator is Pred is a discontiguous procedure. Namely, the clauses defining Pred are not restricted to be consecutive but can appear anywhere in the source file.

This directive shall precede the definition of Pred in the source file.

In order to allow multiple definitions, Pred can also be a list of predicate indicators or a sequence of predicate indicators using ','/2 as separator.

A multifile predicate (declared with a multifile/1 directive) cannot be directly called from a file where it is not declared as multifile (the native compiler must know the called predicate is multifile). Workarounds: either call it via a meta-call (e.g. using call/1) or declare it as multifile in the calling source file). A good habit is to encapsulate a multifile predicate in a monofile predicate which invokes it (external call only invoke the monofile wrapper predicate).

Portability

ISO directive.

7.1.6 ensure linked/1

Templates

     ensure linked(+predicate indicator)
     ensure linked(+predicate indicator list)
     ensure linked(+predicate indicator sequence)

Description

ensure linked(Pred) specifies that the procedure whose predicate indicator is Pred must be included by the linker. This directive is useful when compiling to native code to force the linker to include the code of a given predicate. Indeed, if the gplc is invoked with an option to reduce the size of the executable (section 4.4.3, page 23), the linker only includes the code of predicates that are statically referenced. However, the linker cannot detect dynamically referenced predicates (used as data passed to a meta-call predicate). The use of this directive prevents it to exclude the code of such predicates.

In order to allow multiple definitions, Pred can also be a list of predicate indicators or a sequence of predicate indicators using ','/2 as separator.

Portability

GNU Prolog directive.

7.1.7 built in/0, built in/1, built in fd/0, built in fd/1

Templates

     built in
     built in(+predicate indicator)
     built in(+predicate indicator list)
     built in(+predicate indicator sequence)
     built in fd

     built in fd(+predicate indicator)
     built in fd(+predicate indicator list)
     built in fd(+predicate indicator sequence)

Description

built in specifies that the procedures defined from now have the built in property (section 8.8.2, page 75).

built in(Pred) is similar to built in/0 but only affects the procedure whose predicate indicator is

Pred.

This directive shall precede the definition of Pred in the source file.

In order to allow multiple definitions, Pred can also be a list of predicate indicators or a sequence of predicate indicators using ','/2 as separator.

built in fd (resp. built in fd(Pred)) is similar to built in (resp. built in(Pred)) but sets the

built in fd predicate property (section 8.8.2, page 75).

Portability

GNU Prolog directives.

7.1.8 include/1

Templates

     include(+atom)

Description

include(File) specifies that the content of the Prolog source File shall be inserted. The resulting Prolog text is identical to the Prolog text obtained by replacing the directive by the content of the Prolog source File.

In case of File is a relative file name, it is searched in the current directory. If it is not found it is then searched in each directory of parent includers.

See absolute file name/2 for information about the syntax of File (section 8.26.1, page 158).

Portability

ISO directive.

7.1.9 if/1, else/0, endif/0, elif/1

Templates

     if(+callable term)
     else
     endif
     elif(+callable term)

Description These directives are for conditional compilation. if(Goal) compile subsequent code only if Goal succeeds. Goal is first processed by expand term/2 (section 8.17.2, page 124). If Goal raises an exception it is printed and Goal fails. else introduces the else part. endif terminates a conditional compilation part.

elif(Goal) is a shorthand for :- else.  :- if(Goal). . . . :- endif.

Portability GNU Prolog directive. Also in SWI and YAP.

7.1.10 ensure loaded/1

Templates

     ensure loaded(+atom)

Description ensure loaded(File) is not supported by GNU Prolog. When such a directive is encountered it is simply ignored. Portability ISO directive. Not supported.

7.1.11 op/3 Templates

     op(+integer, +operator specifier, +atom or atom list)

Description op(Priority, OpSpecifier, Operator) alters the operator table. This directive is executed as soon as it is encountered by calling the built-in predicate op/3 (section 8.14.10, page 113). A system directive is also generated to reflect the effect of this directive at run-time (section 4.4.4, page 27). Portability ISO directive.

7.1.12 char conversion/2

Templates

     char conversion(+character, +character)

Description char conversion(InChar, OutChar) alters the character-conversion mapping. This directive is executed as soon as it is encountered by a call to the built-in predicate char conversion/2 (section 8.14.12, page 116). A system directive is also generated to reflect the effect of this directive at run-time (section 4.4.4, page 27). Portability ISO directive.

7.1.13 set prolog flag/2

Templates

     set prolog flag(+flag, +term)

Description set prolog flag(Flag, Value) sets the value of the Prolog flag Flag to Value. This directive is executed as soon as it is encountered by a call to the built-in predicate set prolog flag/2 (section 8.22.1, page 148). A system directive is also generated to reflect the effect of this directive at run-time (section 4.4.4, page 27). Portability ISO directive.

7.1.14 initialization/1

Templates

     initialization(+callable term)

Description initialization(Goal) adds Goal to the set of goal which shall be executed at run-time. A user directive is generated to execute Goal at run-time. If several initialization directives appear in the same file they are executed in the order of appearance (section 4.4.4, page 27). Portability ISO directive.

7.1.15 foreign/2, foreign/1

Templates

     foreign(+callable term, +foreign option list)
     foreign(+callable term)

Description foreign(Template, Options) defines an interface predicate whose prototype is Template according to the options given by Options. Refer to the foreign code interface for more information (section 10.3, page 202).

foreign(Template) is equivalent to foreign(Template, []).

Portability GNU Prolog directive.

7.2 Prolog control constructs

GNU Prolog follows the ISO notion of control constructs.

7.2.1 true/0, fail/0, !/0

Templates

     true
     fail
     !

Description true always succeeds. fail always fails (enforces backtracking). ! always succeeds and the for side-effect of removing all choice-points created since the invocation of the predicate activating it. Errors None. Portability ISO control constructs.

7.2.2 (',')/2 - conjunction, (;)/2 - disjunction, (->)/2 - if-then, (*->)/2 - soft-cut (soft

if-then) Templates

     ','(+callable term, +callable term)
     ;(+callable term, +callable term)
     ->(+callable term, +callable term)
     *->(+callable term, +callable term)

Description

Goal1 , Goal2 executes Goal1 and, in case of success, executes Goal2.

Goal1 ; Goal2 first creates a choice-point and executes Goal1. On backtracking Goal2 is executed.

Goal1 -> Goal2 first executes Goal1 and, in case of success, removes all choice-points created by Goal1 and executes Goal2. This control construct acts like an if-then (Goal1 is the test part and Goal2 the then part). Note that if Goal1 fails ->/2 fails also. ->/2 is often combined with ;/2 to define an if-then-else as follows: Goal1 -> Goal2 ; Goal3. Note that Goal1 -> Goal2 is the first argument of the (;)/2 and Goal3 (the else part) is the second argument. Such an if-then-else control construct first creates a choice-point for the else-part (intuitively associated with ;/2) and then executes Goal1. In case of success, all choice-points created by Goal1 together with the choice-point for the else-part are removed and Goal2 is executed. If Goal1 fails then Goal3 is executed.

Goal1 *-> Goal2 ; Goal3 implements the so-called soft-cut. It acts as the above if-then-else except that if Goal1 succeeds only Goal3 is cut (the alternative solutions of Goal1 are preserved and can be found by backtracking). Note that Goal1 *-> Goal2 alone (i.e. without an else branch Goal3) is equivalent to

(Goal1 , Goal2).

',', ;, -> and *-> are predefined infix operators (section 8.14.10, page 113).

Errors

 Goal1 or Goal2 is a variable  instantiation error

Goal1 is neither a variable nor a callable term type error(callable, Goal1) Goal2 is neither a variable nor a callable term type error(callable, Goal2) The predicate indicator Pred of Goal1 or Goal2 does not correspond to an existing procedure and the value of the unknown Prolog flag is error (section 8.22.1, page 148)

                                             existence error(procedure, Pred)

Portability

ISO control constructs except (*->)/2 which is GNU Prolog specific.

7.2.3 call/1

Templates

     call(+callable term)

Description

call(Goal) executes Goal. call/1 succeeds if Goal represents a goal which is true. When Goal contains a cut symbol ! (section 7.2.1, page 51) as a subgoal, the effect of ! does not extend outside Goal.

 Goal is a variable  instantiation error

Goal is neither a variable nor a callable term type error(callable, Goal) The predicate indicator Pred of Goal does not correspond to an existing procedure and the value of the unknown Prolog flag is error (section 8.22.1, page 148)

                                             existence error(procedure, Pred)

Portability

ISO control construct.

7.2.4 catch/3, throw/1

Templates

     catch(?callable term, ?term, ?term)
     throw(+nonvar)

Description

catch(Goal, Catcher, Recovery) is similar to call(Goal) (section 7.2.3, page 52). If this succeeds or fails, so does the call to catch/3. If however, during the execution of Goal, there is a call to throw(Ball), the current flow of control is interrupted, and control returns to a call of catch/3 that is being executed. This can happen in one of two ways:

• implicitly, when an error condition for a built-in predicate is satisfied.

• explicitly, when the program executes a call of throw/1 because the program wishes to abandon the current processing, and instead to take an alternative action.

throw(Ball) causes the normal flow of control to be transferred back to an existing call of catch/3. When a call to throw(Ball) happens, Ball is copied and the stack is unwound back to the call to catch/3, whereupon the copy of Ball is unified with Catcher. If this unification succeeds, then catch/3 executes the goal Recovery using call/1 (section 7.2.3, page 52) in order to determine the success or failure of catch/3. Otherwise, in case the unification fails, the stack keeps unwinding, looking for an earlier invocation of catch/3. Ball may be any non-variable term.

Errors

 Ball is a variable  instantiation error

If Ball does not unify with the Catcher argument of any call of catch/3, a system error message is displayed and throw/1 fails.

When catch/3 calls Goal or Recovery it uses call/1 (section 7.2.3, page 52), an instantiation error,
a type error or an existence error can then occur depending on Goal or Recovery.

Portability

ISO control constructs.

54  7 PROLOG DIRECTIVES AND CONTROL CONSTRUCTS

8 Prolog built-in predicates

8.1 Type testing

8.1.1 var/1, nonvar/1, atom/1, integer/1, float/1, number/1, atomic/1,
       compound/1, callable/1, ground/1, is list/1, list/1,
       partial list/1, list or partial list/1

Templates

     var(?term)
     nonvar(?term)
     atom(?term)
     integer(?term)
     float(?term)
     number(?term)
     atomic(?term)

                                                 compound(?term)
                                                 callable(?term)
                                                 ground(?term)
                                                 is list(?term)
                                                 list(?term)
                                                 partial list(?term)
                                                 list or partial list(?term)

Description

var(Term) succeeds if Term is currently uninstantiated (which therefore has not been bound to anything, except possibly another uninstantiated variable).

nonvar(Term) succeeds if Term is currently instantiated (opposite of var/1).

atom(Term) succeeds if Term is currently instantiated to an atom.

integer(Term) succeeds if Term is currently instantiated to an integer.

float(Term) succeeds if Term is currently instantiated to a floating point number.

number(Term) succeeds if Term is currently instantiated to an integer or a floating point number.

atomic(Term) succeeds if Term is currently instantiated to an atom, an integer or a floating point number.

compound(Term) succeeds if Term is currently instantiated to a compound term, i.e. a term of arity > 0 (a list or a structure).

callable(Term) succeeds if Term is currently instantiated to a callable term, i.e. an atom or a compound term.

ground(Term) succeeds if Term is a ground term.

list(Term) succeeds if Term is currently instantiated to a list, i.e. the atom [] (empty list) or a term with principal functor '.'/2 and with second argument (the tail) a list.

is list(Term) behaves like list(Term) (for compatibility purpose).

partial list(Term) succeeds if Term is currently instantiated to a partial list, i.e. a variable or a term whose the main functor is '.'/2 and the second argument (the tail) is a partial list.

list or partial list(Term) succeeds if Term is currently instantiated to a list or a partial list.

Errors None. Portability

var/1, nonvar/1, atom/1, integer/1, float/1, number/1, atomic/1, compound/1 callable/1 and

ground/1 are ISO predicates.

list/1, partial list/1 and list or partial list/1 are GNU Prolog predicates.

8.2 Term unification

8.2.1 (=)/2 - Prolog unification Templates

     =(?term, ?term)

Description Term1 = Term2 unifies Term1 and Term2. No occurs check is done, i.e. this predicate does not check if a variable is unified with a compound term containing this variable (this can lead to an infinite loop). = is a predefined infix operator (section 8.14.10, page 113). Errors None. Portability ISO predicate.

8.2.2 unify with occurs check/2

Templates

     unify with occurs check(?term, ?term)

Description

unify with occurs check(Term1, Term2) unifies Term1 and Term2. The occurs check test is done (i.e.

the unification fails if a variable is unified with a compound term containing this variable). Errors None. Portability ISO predicate.

8.2.3 (\=)/2 - not Prolog unifiable

Templates

     \=(?term, ?term)

Description Term1 \= Term2 succeeds if Term1 and Term2 are not unifiable (no occurs check is done). \= is a predefined infix operator (section 8.14.10, page 113). Errors None. Portability ISO predicate.

8.3 Term comparison

8.3.1 Standard total ordering of terms

The built-in predicates described in this section allows the user to compare Prolog terms. Prolog terms are totally ordered according to the standard total ordering of terms which is as follows (from the smallest term to the greatest):

• variables, oldest first.

• finite domain variables (section 9.1.1, page 183), oldest first.

• floating point numbers, in numeric order.

• integers, in numeric order.

• atoms, in alphabetical (i.e. character code) order.

• compound terms, ordered first by arity, then by the name of the principal functor and by the

arguments in left-to-right order. A list is treated as a compound term (whose principal functor is '.'/2). The portability of the order of variables is not guaranteed (in the ISO reference the order of variables is system dependent).

8.3.2 (==)/2 - term identical, (\==)/2 - term not identical, (@<)/2 - term less than, (@=<)/2 - term less than or equal to, (@>)/2 - term greater than, (@>=)/2 - term greater than or equal to

     ==(?term, ?term)
     \==(?term, ?term)
     @<(?term, ?term)

                                                 @=<(?term, ?term)
                                                 @>(?term, ?term)
                                                 @>=(?term, ?term)

Description These predicates compare two terms according to the standard total ordering of terms (section 8.3.1, page 57).

Term1 == Term2 succeeds if Term1 and Term2 are equal.

Term1 \== Term2 succeeds if Term1 and Term2 are different.

Term1 @< Term2 succeeds if Term1 is less than Term2.

Term1 @=< Term2 succeeds if Term1 is less than or equal to Term2. Term1 @> Term2 succeeds if Term1 is greater than Term2. Term1 @>= Term2 succeeds if Term1 is greater than or equal to Term2. ==, \==, @<, @=<, @> and @>= are predefined infix operators (section 8.14.10, page 113). Errors None. Portability ISO predicates.

8.3.3 compare/3

Templates

     compare(?atom, +term, +term)

Description compare(Order, Term1, Term2) compares Term1 and Term2 according to the standard (section 8.3.1, page 57) and unifies Order with:

• the atom < if Term1 is less than Term2.

• the atom = if Term1 and Term2 are equal.

• the atom > if Term1 is greater than Term2. Errors Order is neither a variable nor an atom type error(atom, Order)

 Order is an atom but not <, = or >              domain error(order, Order)

Portability ISO predicate.

8.4 Term processing

8.4.1 functor/3

Templates

    functor(+nonvar, ?atomic, ?integer)
    functor(-nonvar, +atomic, +integer)

Description

functor(Term, Name, Arity) succeeds if the principal functor of Term is Name and its arity is Arity. This predicate can be used in two ways:

• Term is not a variable: extract the name (an atom or a number if Term is a number) and the arity of Term (if Term is atomic Arity = 0).

• Term is a variable: unify Term with a general term whose principal functor is given by Name and arity is given by Arity.

Errors

Term and Name are both variables  instantiation error
Term and Arity are both variables  instantiation error

Term is a variable and Name is neither a variable nor an atomic term

                                           type error(atomic, Name)

Term is a variable and Arity is neither a variable nor an integer

                                           type error(integer, Arity)

Term is a variable, Name is a constant but not an atom and Arity is an integer > 0

                                           type error(atom, Name)

Term is a variable and Arity is an integer > max arity flag (section 8.22.1, page 148)

                                           representation error(max arity)
Term is a variable and Arity is an integer < 0  domain error(not less than zero, Arity)

Portability

ISO predicate.

8.4.2 arg/3

Templates

    arg(+integer, +compound term, ?term)

Description

arg(N, Term, Arg) succeeds if the Nth argument of Term is Arg.

Errors

N is a variable  instantiation error
Term is a variable  instantiation error

N is neither a variable nor an integer type error(integer, N) Term is neither a variable nor a compound term type error(compound, Term)

N is an integer < 0  domain error(not less than zero, N)

Portability ISO predicate.

8.4.3 (=..)/2 - univ Templates

     =..(+nonvar, ?list)
     =..(-nonvar, +list)

Description Term =.. List succeeds if List is a list whose head is the atom corresponding to the principal functor of Term and whose tail is a list of the arguments of Term. =.. is a predefined infix operator (section 8.14.10, page 113). Errors Term is a variable and List is a partial list instantiation error List is neither a partial list nor a list type error(list, List) Term is a variable and List is a list whose head is a variable

                                             instantiation error

List is a list whose head H is neither an atom nor a variable and whose tail is not the empty list

                                             type error(atom, H)

List is a list whose head H is a compound term and whose tail is the empty list

                                             type error(atomic, H)
 Term is a variable and List is the empty list  domain error(non empty list, [])

Term is a variable and the tail of List has a length > max arity flag (section 8.22.1, page 148)

                                             representation error(max arity)

Portability ISO predicate.

8.4.4 copy term/2

Templates

     copy term(?term, ?term)

Description copy term(Term1, Term2) succeeds if Term2 unifies with a term T which is a renamed copy of Term1. Errors None. Portability ISO predicate.

8.4.5 term variables/2, term variables/3

Templates

     term variables(?term, ?list)
     term variables(?term, ?list, ?list)

Description term variables(Term, List) succeeds if List unifies with a list of variables (including FD variables), each sharing with a unique variable of Term. The variables in List are ordered in order of appearance traversing Term depth-first and left-to-right. term variables(Term, List, Tail) is a difference-list version of the above predicate, i.e. Tail is the tail of the variable-list List. Errors

 in term variables/2 List is neither a partial

list nor a list

                                             type error(list, List)

Portability term variables/2 is an ISO Predicate. term variables/3 is a GNU Prolog predicate.

8.4.6 subsumes term/2

Templates

     subsumes term(?term, ?term)

Description

subsumes term(General, Specific) succeeds if General can be made equivalent to Specific by bind-

ing variables in General leaving Specific unaffected. The current implementation performs the unification (with occurs check) and ensures that the variable set of Specific is not changed by the unification (which is then undone). Note that this predicate fails in the presence of FD variables in Specific. Errors None. Portability ISO predicate.

8.4.7 acyclic term/1

Templates

     acyclic term(?term)

Description acyclic term(Term) succeeds if Term does not contain a cyclic (sub-)term. In this case, Term may be processed safely. If acyclic term(Term) fails, Term contains a cycle and processing Term is not safe, because GNU Prolog does not support the unification of cyclic terms but permits their creation. Cycles can be safely undone by failing over their creation. The use of acyclic term/1 shall thus be reserved to protect critical predicates against cyclic terms. Errors None. Portability ISO predicate.

8.4.8 term hash/4, term hash/2

Templates

     term hash(?term, +integer, +integer, ?integer)
     term hash(?term, ?integer)

Description

term hash(Term, Depth, Range, Hash) succeeds if Hash is the hash code of Term. If Term is not ground

(see ground/1 (section 8.1.1, page 55)), the predicate simply succeeds (Hash is not unified). Depth is the depth limit to scan Term (starting from 1 for the top-level term). With Depth = 0 nothing is hashed, with 1 only atomic terms and the main functors/arity are hashed,... With Depth = -1 the full term is considered. The hash code is as follows: 0 ≤ Hash < Range. If Range = 0 then Hash is not restricted (currently it is < 268435456).

term hash(Term, Hash) is equivalent to term hash(Term, -1, 0, Hash).

NB: the computed hash code is independent of any runtime context (i.e. it is constant across different executions). It is also independent on the underlying machine. These predicates are useful to implement hash tables or argument indexing. Errors

 Depth is a variable  instantiation error

Depth is neither a variable nor an integer type error(integer, Depth)

 Range is a variable  instantiation error

Range is neither a variable nor an integer type error(integer, Range)

 Range is an integer < 0  domain error(not less than zero, Range)

Hash is neither a variable nor an integer type error(integer, Hash) Portability GNU Prolog predicate.

8.4.9 setarg/4, setarg/3

Templates

     setarg(+integer, +compound term, +term, +boolean)
     setarg(+integer, +compound term, +term)

Description

setarg(N, Term, NewValue, Undo) replaces destructively the Nth argument of Term with NewValue.

This assignment is undone on backtracking if Undo = true. This should only used if there is no further use of the old value of the replaced argument. If Undo = false then NewValue must be either an atom or an integer.

setarg(N, Term, NewValue) is equivalent to setarg(N, Term, NewValue, true).

Errors

 N is a variable  instantiation error

N is neither a variable nor an integer type error(integer, N)

 N is an integer < 0  domain error(not less than zero, N)
 Term is a variable  instantiation error

Term is neither a variable nor a compound term type error(compound, Term) NewValue is neither an atom nor an integer and

 Undo = false

                                             type error(atomic, NewValue)
 Undo is a variable  instantiation error

Undo is neither a variable nor a boolean type error(boolean, Undo) Portability GNU Prolog predicate.

8.5 Variable naming/numbering

8.5.1 name singleton vars/1

Templates

     name singleton vars(?term)

Description name singleton vars(Term) binds each singleton variable appearing in Term with a term of the form '$VARNAME'(' '). Such a term can be output by write term/3 as a variable name (section 8.14.6, page 108). Errors None. Portability GNU Prolog predicates.

8.5.2 name query vars/2

Templates

     name query vars(+list, ?list)

Description

name query vars(List, Rest) for each element of List of the form Name = Var where Name is an

atom and Var a variable, binds Var with the term '$VARNAME'(Name). Such a term can be output by write term/3 as a variable name (section 8.14.6, page 108). Rest is unified with the list of elements of List that have not given rise to a binding. This predicate is provided as a way to name the vari-

able lists obtained returned by read term/3 with variable names(List) or singletons(List) options

(section 8.14.1, page 104).

Errors

 List is a partial list  instantiation error

List is neither a partial list nor a list type error(list, List) Rest is neither a partial list nor a list type error(list, Rest)

Portability

GNU Prolog predicate.

8.5.3 bind variables/2, numbervars/3, numbervars/1

Templates

     bind variables(?term, +var binding option list)
     numbervars(?term, +integer, ?integer)
     numbervars(?term)

Description

bind variables(Term, Options) binds each variable appearing in Term according to the options given

by Options.

Variable binding options: Options is a list of variable binding options. If this list contains contradictory options, the rightmost option is the one which applies. Possible options are:

• numbervars: specifies that each variable appearing in Term should be bound to a term of the form '$VAR'(N) where N is an integer. Such a term can be output by write term/3 as a variable name (section 8.14.6, page 108). This is the default.

• namevars: specifies that each variables appearing in Term shall be bound to a term of the form '$VARNAME'(Name) where Name is the atom that would be output by write term/3 seeing a term of the '$VAR'(N) where N is an integer. Such a term can be output by write term/3 as a variable name (section 8.14.6, page 108). This is the alternative to numbervars.

• from(From): the first integer N to use for number/name variables of Term is From. The default value is 0.

• next(Next): when bind variables/2 succeeds, Next is unified with the (last integer N)+1 used to bind the variables of Term.

• exclude(List): collects all variable names appearing in List to avoid a clash when binding a

variable of Term. Precisely a number N ≥ From will not be used to bind a variable of Term if:

– there is a sub-term of List of the form '$VAR'(N) or '$VARNAME'(Name) where Name is the constant that would be output by write term/3 seeing a term of the '$VAR'(N).

– an element of List is of the form Name = Var where Name is an atom that would be output by write term/3 on seeing a term of the from '$VAR'(N). This case allows for lists returned

         by read term/3 (with variable names(List) or singletons(List) options) (section 8.14.1,

page 104) and by name query vars/2 (section 8.5.2, page 64).

numbervars(Term, From, Next) is equivalent to bind variables(Term, [from(From), next(Next)],

i.e. each variable of Term is bound to '$VAR'(N) where From ≤ N < Next.

numbervars(Term) is equivalent to numbervars(Term, 0,  ).

See also term variables (section 8.4.5, page 61) which returns the set of variables of a term.

Errors

Options is a partial list or a list with an element E which is a variable

                                             instantiation error
 Options is neither a partial list nor a list  type error(list, Options)

an element E of the Options list is neither a variable nor a variable binding option

                                             domain error(var binding option, E)
 From is a variable  instantiation error

From is neither a variable nor an integer type error(integer, From) Next is neither a variable nor an integer type error(integer, Next)

 List is a partial list  instantiation error

List is neither a partial list nor a list type error(list, List)

Portability

GNU Prolog predicates.

8.5.4 term ref/2

Templates

     term ref(+term, ?integer)
     term ref(?term, +integer)

Description

term ref(Term, Ref) succeeds if the internal reference of Term is Ref. This predicate can be used either to obtain the internal reference of a term or to obtain the term associated with a given reference. Note that two identical terms can have different internal references. A good way to use this predicate is to first record the internal reference of a given term and to later re-obtain the term via this reference.

Errors

 Term and Ref are both variables  instantiation error

Ref is neither a variable nor an integer type error(integer, Ref)

 Ref is an integer < 0  domain error(not less than zero, Ref)

GNU Prolog predicate.

8.6 Arithmetic

8.6.1 Evaluation of an arithmetic expression

An arithmetic expression is a Prolog term built from numbers, variables, and functors (or operators) that represent arithmetic functions. When an expression is evaluated each variable must be bound to a non-variable expression. An expression evaluates to a number, which may be an integer or a floating point number. The following table details the components of an arithmetic expression, how they are evaluated, the types expected/returned and if they are ISO or an extension:

Expression  Result = eval(Expression)  Signature ISO
a variable  bound to an expression E, result is eval(E)  IF → IF  Y
an integer number  this number  I  Y
a floating point number this number  F  Y
pi  the value of π = 3.141592...  F  Y
e  the value of e = 2.718281...  F  N
epsilon  difference between 1.0 and minimum float > 1.0  F  N
+ E                   eval(E)  IF → IF  Y
- E                   - eval(E)  IF → IF  Y
inc(E)                eval(E) + 1  IF → IF  N
dec(E)                eval(E) - 1  IF → IF  N
E1 + E2  eval(E1) + eval(E2)  IF, IF → IF Y
E1 - E2  eval(E1) - eval(E2)  IF, IF → IF Y
E1 * E2  eval(E1) * eval(E2)  IF, IF → IF Y
E1 / E2  eval(E1) / eval(E2)  IF, IF → F Y
E1 // E2  rnd(eval(E1) / eval(E2))  I, I → I  Y
E1 rem E2  eval(E1) - (rnd(eval(E1) / eval(E2)) * eval(E2))  I, I → I  Y
E1 div E2  ⌊(eval(E1) - eval(E1) mod eval(E2)) / eval(E2)⌋  I, I → I  Y
E1 mod E2  eval(E1) - (⌊eval(E1) / eval(E2)⌋ * eval(E2))  I, I → I  Y
E1 /\ E2  eval(E1) bitwise and eval(E2)  I, I → I  Y
E1 \/ E2  eval(E1) bitwise or eval(E2)  I, I → I  Y
xor(E1,E2)  eval(E1) bitwise xor eval(E2)  I, I → I  Y
\ E  bitwise not eval(E)  I → I  Y
E1 << E2  eval(E1) integer shift left eval(E2)  I, I → I  Y
E1 >> E2  eval(E1) integer shift right eval(E2)  I, I → I  Y
lsb(E)  least significant bit (from 0) of eval(E) or -1  I → I  N
msb(E)  most significant bit (from 0) of eval(E) or -1  I → I  N
popcount(E)  number of 1-bits in eval(E)  I → I  N
abs(E)  absolute value of eval(E)  IF → IF  Y
sign(E)  sign of eval(E) (-1 if < 0, 0 if = 0, +1 if > 0)  IF → IF  Y
min(E1,E2)  minimal value between eval(E1) and eval(E2)  IF, IF → ? Y
max(E1,E2)  maximal value between eval(E1) and eval(E2)  IF, IF → ? Y

gcd(E1,E2)

greatest common divisor of eval(E1) and eval(E2)

I, I → I N
E1 ^E2  eval(E1) raised to the power of eval(E2)  IF, IF → IF Y
E1 ** E2  eval(E1) raised to the power of eval(E2)  IF, IF → F Y
sqrt(E)  square root of eval(E)  IF → F  Y
tan(E)  tangent of eval(E)  IF → F  Y
atan(E)  arc tangent of eval(E)  IF → F  Y

atan2(Y,X)

principal value of arc tangent of eval(Y) / eval(X) using both signs for the quadrant

IF → F Y
cos(E)  cosine of eval(E)  IF → F  Y
acos(E)  arc cosine of eval(E)  IF, IF → F Y
sin(E)  sine of eval(E)  IF → F  Y
asin(E)  arc sine of eval(E)  IF → F  Y
tanh(E)  hyperbolic tangent of eval(E)  IF → F  N
atanh(E)  hyperbolic arc tangent of eval(E)  IF → F  N
cosh(E)  hyperbolic cosine of eval(E)  IF → F  N
acosh(E)  hyperbolic arc cosine of eval(E)  IF, IF → F N
sinh(E)  hyperbolic sine of eval(E)  IF → F  N
asinh(E)  hyperbolic arc sine of eval(E)  IF → F  N
exp(E)  e raised to the power of eval(E)  IF → F  Y

Expression  Result = eval(Expression)  Signature ISO
log(E)  natural logarithm of eval(E)  IF → F  Y
log10(E)  base 10 logarithm of eval(E)  IF → F  N
log(R, E)  base eval(R) logarithm of eval(E)  F, IF → F N
float(E)  the floating point number equal to eval(E)  IF → F  Y
ceiling(E)  rounds eval(E) upward to the nearest integer  F → I  Y
floor(E)  rounds eval(E) downward to the nearest integer  F → I  Y
round(E)  rounds eval(E) to the nearest integer  F → I  Y
truncate(E)  the integer value of eval(E)  F → I  Y
float fractional part(E) the float equal to the fractional part of eval(E)  F → F  Y
float integer part(E)  the float equal to the integer part of eval(E)  F → F  Y

The meaning of the signature field is as follows:

• I → I: unary function, the operand must be an integer and the result is an integer.

• F → F: unary function, the operand must be a floating point number and the result is a floating point number.

• F → I: unary function, the operand must be a floating point number and the result is an integer.

• IF → F: unary function, the operand can be an integer or a floating point number and the result is a floating point number.

• IF → IF: unary function, the operand can be an integer or a floating point number and the result has the same type as the operand.

• I, I → I: binary function: each operand must be an integer and the result is an integer.

• IF, IF → IF: binary function: each operand can be an integer or a floating point number and the result is a floating point number if at least one operand is a floating point number, an integer otherwise.

• IF, IF → ?: binary function: each operand can be an integer or a floating point number and the result has the same type as the selected operand. This is used for min and max. Note that in case of equality between an integer and a floating point number the result is an integer.

is, +, -, *, /, //, div, rem, mod, /\, \/, <<, >>, ** and ^are predefined infix operators. +, - and \, are predefined prefix operators (section 8.14.10, page 113).

Integer division rounding function: the integer division rounding function rnd(X) rounds the floating point number X to an integer. There are two possible definitions (depending on the target machine) for this function which differ on negative numbers:

• rnd(X) = integer part of X, e.g. rnd(-1.5) = -1 (round toward 0)

• rnd(X) = ⌊X⌋, e.g. rnd(-1.5) = -2 (round toward −∞)

The definition of this function determines the definition of the integer division and remainder ((//)/2

and (rem)/2). It is possible to test the value (toward zero or down) of the integer rounding function

Prolog flag to determine which function being used (section 8.22.1, page 148). Since rounding toward zero is the most common case, two additional evaluable functors ((div)/2 and (mod)/2) are available which consider rounding toward −∞.

Fast mathematical mode: in order to speed-up integer computations, the GNU Prolog compiler can generate faster code when invoked with the --fast-math option (section 4.4.3, page 23). In this mode only integer operations are allowed and a variable in an expression must be bound at evaluation time to an integer. No type checking is done.

Errors

a sub-expression E is a variable instantiation error a sub-expression E is neither a number nor an evaluable functor

                                             type error(evaluable, E)

a sub-expression E is a floating point number while an integer is expected

                                             type error(integer, E)

a sub-expression E is an integer while a floating point number is expected

                                             type error(float, E)
 a division by zero occurs  evaluation error(zero divisor)

Portability

Refer to the above table to determine which evaluable functors are ISO and which are GNU Prolog extensions. For efficiency reasons, GNU Prolog does not detect the following ISO arithmetic errors:

float overflow, int overflow, int underflow, and undefined.

8.6.2 (is)/2 - evaluate expression

Templates

     is(?term, +evaluable)

Description

Result is Expression succeeds if Result can be unified with eval(Expression). Refer to the evaluation of an arithmetic expression for the definition of the eval function (section 8.6.1, page 66).

is is a predefined infix operator (section 8.14.10, page 113).

Errors

Refer to the evaluation of an arithmetic expression for possible errors (section 8.6.1, page 66).

Portability

ISO predicate.

8.6.3 (=:=)/2 - arithmetic equal, (=\=)/2 - arithmetic not equal, (<)/2 - arithmetic less than, (=<)/2 - arithmetic less than or equal to, (>)/2 - arithmetic greater than, (>=)/2 - arithmetic greater than or equal to

Templates

     =:=(+evaluable, +evaluable)
     =\=(+evaluable, +evaluable)
     <(+evaluable, +evaluable)

                                                 =<(+evaluable, +evaluable)
                                                 >(+evaluable, +evaluable)
                                                 >=(+evaluable, +evaluable)

Description

Expr1 =:= Expr2 succeeds if eval(Expr1) = eval(Expr2).

Expr1 =\= Expr2 succeeds if eval(Expr1) ≠ eval(Expr2). Expr1 < Expr2 succeeds if eval(Expr1) < eval(Expr2). Expr1 =< Expr2 succeeds if eval(Expr1) ≤ eval(Expr2). Expr1 > Expr2 succeeds if eval(Expr1) > eval(Expr2). Expr1 >= Expr2 succeeds if eval(Expr1) ≥ eval(Expr2). Refer to the evaluation of an arithmetic expression for the definition of the eval function (section 8.6.1, page 66). =:=, =\=, <, =<, > and >= are predefined infix operators (section 8.14.10, page 113). Errors Refer to the evaluation of an arithmetic expression for possible errors (section 8.6.1, page 66). Portability ISO predicates.

8.6.4 succ/2

Templates

     succ(+integer, ?integer)
     succ(-integer, +integer)

Description succ(X, Y) is true iff Y is the successor of the non-negative integer X. Errors

 X and Y are both variables  instantiation error

X is neither a variable nor an integer type error(integer, X) Y is neither a variable nor an integer type error(integer, Y)

 X is an integer < 0  domain error(not less than zero, X)
 Y is an integer < 0  domain error(not less than zero, Y)

Portability GNU Prolog predicate.

8.7 Dynamic clause management

8.7.1 Introduction Static and dynamic procedures: a procedure is either dynamic or static. All built-in predicates are static. A user-defined procedure is static by default unless a dynamic/1 directive precedes its definition (section 7.1.2, page 45). Adding a clause to a non-existent procedure creates a dynamic procedure. The clauses of a dynamic procedure can be altered (e.g. using asserta/1), the clauses of a static procedure cannot be altered.

Private and public procedures: each procedure is either public or private. A dynamic procedure is always public. Each built-in predicate is private, and a static user-defined procedure is private by default unless a public/1 directive precedes its definition (section 7.1.3, page 45). If a dynamic declaration exists it is unnecessary to add a public declaration since a dynamic procedure is also public. A clause of a public procedure can be inspected (e.g. using clause/2), a clause of a private procedure cannot be inspected.

A logical database update view: any change in the database that occurs as the result of executing a goal (e.g. when a sub-goal is a call of assertz/1 or retract/1) only affects subsequent activations. The change does not affect any activation that is currently being executed. Thus the database is frozen during the execution of a goal, and the list of clauses defining a predication is fixed at the moment of its execution.

8.7.2 asserta/1, assertz/1

Templates

     asserta(+clause)
     assertz(+clause)

Description

asserta(Clause) first converts the term Clause to a clause and then adds it to the current internal database. The predicate concerned must be dynamic (section 8.7.1, page 70) or undefined and the clause is inserted before the first clause of the predicate. If the predicated is undefined it is created as a dynamic procedure.

assertz(Clause) acts like asserta/1 except that the clause is added at the end of all existing clauses of the concerned predicate.

Converting a term Clause to a clause Clause1:

• extract the head and the body of Clause: either Clause = (Head :- Body) or Clause = Head and

     Body = true.

• Head must be a callable term (or else the conversion fails).

• convert Body to a body clause (i.e. a goal) Body1.

• the converted clause Clause1 = (Head :- Body1).

Converting a term T to a goal:

• if T is a variable it is replaced by the term call(T).

• if T is a control construct (',')/2, (;)/2 or (->)/2 each argument of the control construct is recursively converted to a goal.

• if T is a callable term it remains unchanged.

• otherwise the conversion fails (T is neither a variable nor a callable term).

Errors

 Head is a variable  instantiation error

Head is neither a variable nor a callable term type error(callable, Head)

 Body cannot be converted to a goal  type error(callable, Body)

The predicate indicator Pred of Head is that of a static procedure

                                             permission error(modify,
                                             static procedure, Pred)

Portability

ISO predicates.

8.7.3 retract/1

Templates

     retract(+clause)

Description

retract(Clause) erases the first clause of the database that unifies with Clause. The concerned predicate must be a dynamic procedure (section 8.7.1, page 70). Removing all clauses of a procedure does not erase the procedure definition. To achieve this use abolish/1 (section 8.7.6, page 73). retract/1 is re-executable on backtracking.

Errors

 Head is a variable  instantiation error

Head is neither a variable nor a callable term type error(callable, Head) The predicate indicator Pred of Head is that of a static procedure

                                             permission error(modify,
                                             static procedure, Pred)

Portability

ISO predicate. In the ISO reference, the operation associated with the permission error is access while it is modify in GNU Prolog. This seems to be an error of the ISO reference since for asserta/1 (which is similar in spirit to retract/1) the operation is also modify.

8.7.4 retractall/1

Templates

     retractall(+head)

Description

retractall(Head) erases all clauses whose head unifies with Head. The concerned predicate must be a dynamic procedure (section 8.7.1, page 70). The procedure definition is not removed so that it is found by current predicate/1 (section 8.8.1, page 74). abolish/1 should be used to remove the procedure (section 8.7.6, page 73).

 Head is a variable  instantiation error
 Head is not a callable term  type error(callable, Head)

The predicate indicator Pred of Head is that of a static procedure

                                             permission error(modify,
                                             static procedure, Pred)

Portability ISO predicate.

8.7.5 clause/2

Templates

     clause(+head, ?callable term)

Description clause(Head, Body) succeeds if there exists a clause in the database that unifies with Head :- Body. The predicate in question must be a public procedure (section 8.7.1, page 70). Clauses are delivered from the first to the last. This predicate is re-executable on backtracking. Errors

 Head is a variable  instantiation error

Head is neither a variable nor a callable term type error(callable, Head) The predicate indicator Pred of Head is that of a private procedure

                                             permission error(access,
                                             private procedure, Pred)

Body is neither a variable nor a callable term type error(callable, Body) Portability ISO predicate.

8.7.6 abolish/1

Templates

     abolish(+predicate indicator)

Description abolish(Pred) removes from the database the procedure whose predicate indicator is Pred. The concerned predicate must be a dynamic procedure (section 8.7.1, page 70). Errors

Pred is a variable  instantiation error
Pred is a term Name/Arity and either Name or

Arity is a variable

                                           instantiation error

Pred is neither a variable nor a predicate indicator

                                           type error(predicate indicator, Pred)
Pred is a term Name/Arity and Arity is neither

a variable nor an integer

                                           type error(integer, Arity)

Pred is a term Name/Arity and Name is neither a variable nor an atom

                                           type error(atom, Name)
Pred is a term Name/Arity and Arity is an

integer < 0

                                           domain error(not less than zero, Arity)
Pred is a term Name/Arity and Arity is an

integer > max arity flag (section 8.22.1, page 148)

                                           representation error(max arity)

The predicate indicator Pred is that of a static procedure

                                           permission error(modify,
                                           static procedure, Pred)

Portability

ISO predicate.

8.8 Predicate information

8.8.1 current predicate/1

Templates

    current predicate(?predicate indicator)

Description

current predicate(Pred) succeeds if there exists a predicate indicator of a defined procedure that unifies with Pred. All user defined procedures are found, whether static or dynamic. Internal system procedures whose name begins with '$' are not found. A user-defined procedure is found even when it has no clauses. A user-defined procedure is not found if it has been abolished. To conform to the ISO reference, built-in predicates are not found except if the strict iso Prolog flag is switched off (section 8.22.1, page 148). This predicate is re-executable on backtracking.

Errors

Pred is neither a variable nor a predicate indicator

                                           type error(predicate indicator, Pred)
Pred is a term Name/Arity and Arity is neither

a variable nor an integer

                                           type error(integer, Arity)

Pred is a term Name/Arity and Name is neither a variable nor an atom

                                           type error(atom, Name)
Pred is a term Name/Arity and Arity is an

integer < 0

                                           domain error(not less than zero, Arity)
Pred is a term Name/Arity and Arity is an

integer > max arity flag (section 8.22.1, page 148)

                                           representation error(max arity)

ISO predicate.

8.8.2 predicate property/2

Templates

  predicate property(?callable, ?predicate property)

Description predicate property(Head, Property) succeeds if Head refers to a predicate that has a property Property. All user defined procedures and built-in predicates are found. Internal system procedures whose name begins with '$' are not found. This predicate is re-executable on backtracking. Since version 1.4.0, predicate property/2 no longer accepts a predicate indicator. Control constructs are now returned. Properties built in fd and control construct now imply the property built in. Predicate properties:

• static: if the procedure is static.

• dynamic: if the procedure is dynamic.

• private: if the procedure is private.

• public: if the procedure is public.

• monofile: if the procedure is monofile.

• multifile: if the procedure is multifile.

• user: if the procedure is a user-defined procedure.

• built in: if the procedure is a built-in predicate or a control construct.

• built in fd: if the procedure is an FD built-in predicate.

• control construct: if the procedure is a control construct (section 7.2, page 51).

• native code: if the procedure is compiled in native code.

• prolog file(File): source file from which the procedure has been read.

• prolog line(Line): line number of the source file.

• meta predicate(Head): if the procedure is a meta-predicate unify Head with the head-pattern. The head-pattern is a compound term with the same name and arity as the predicate where each argument of the term is a meta argument specifier as follows: integer N the argument is a term that is used to reference a predicate with N more arguments than

the given argument term (e.g. call(0)). : the argument is module sensitive, but does not directly refer to a predicate (e.g. consult(:)).

- the argument is not module sensitive and unbound on entry.

? the argument is not module sensitive and the mode is unspecified.

+ the argument is not module sensitive and bound (i.e., nonvar) on entry.

Errors

Head is neither a variable nor a callable term type error(callable, Head) Property is neither a variable nor a predicate property term

                                             domain error(predicate property,
                                             Property)
 Property = prolog file(File) and File is

neither a variable nor an atom

                                             type error(atom, File)
 Property = prolog line(Line) and Line is

neither a variable nor an integer

                                             type error(integer, Line)

Portability

GNU Prolog predicate.

8.9 All solutions

8.9.1 Introduction

It is sometimes useful to collect all solutions for a goal. This can be done by repeatedly backtracking and gradually building the list of solutions. The following built-in predicates are provided to automate this process.

The built-in predicates described in this section invoke call/1 (section 7.2.3, page 52) on the argument Goal. When efficiency is crucial and Goal is complex it is better to define an auxiliary predicate which can then be compiled, and have Goal call this predicate.

8.9.2 findall/4, findall/3

Templates

     findall(?term, +callable term, ?list, ?term) findall(?term, +callable term, ?list)

Description

findall(Template, Goal, Instances) succeeds if Instances unifies with the list of values to which a variable X not occurring in Template or Goal would be instantiated by successive re-executions of call(Goal), X = Template after systematic replacement of all variables in X by new variables. Thus, the order of the list Instances corresponds to the order in which the proofs are found.

findall(Template, Goal, Instances, Tail) is the difference list version of findall/3. The result
is the difference list Instances-Tail. Thus findall(Template, Goal, Instances) is equivalent to
findall(Template, Goal, Instances, []).

 Goal is a variable  instantiation error

Goal is neither a variable nor a callable term type error(callable, Goal) The predicate indicator Pred of Goal does not correspond to an existing procedure and the value of the unknown Prolog flag is error (section 8.22.1, page 148)

                                             existence error(procedure, Pred)

 Instances is neither a partial list nor a list  type error(list, Instances)

Tail is neither a partial list nor a list type error(list, Tail)

Portability

findall/3 is an ISO predicate. findall/4 is a GNU Prolog predicate.

8.9.3 bagof/3, setof/3

Templates

     bagof(?term, +callable term, ?list)
     setof(?term, +callable term, ?list)

Description

bagof(Template, Goal, Instances) assembles as a list the set of solutions of Goal for each different instantiation of the free variables in Goal. The elements of each list are in order of solution, but the order in which each list is found is undefined. This predicate is re-executable on backtracking.

Free variable set: bagof/3 groups the solutions of Goal according to the free variables in Goal. This set corresponds to all variables occurring in Goal but not in Template. It is sometimes useful to exclude some additional variables of Goal. For that, bagof/3 recognizes a goal of the form T^Goal and exclude all variables occurring in T from the free variable set. (^)/2 can be viewed as an existential quantifier (the logical reading of X^Goal being “there exists an X such that Goal is true”). The use of this existential qualifier is superfluous outside bagof/3 (and setof/3) and then is not recognized.

(^)/2 is a predefined infix operator (section 8.14.10, page 113).

setof(Template, Goal, Instances) is equivalent to bagof(Template,Goal,I), sort(I,Instances).

Each list is then a sorted list (duplicate elements are removed).

From the implementation point of view setof/3 is as fast as bagof/3. Both predicates use an in-place (i.e. destructive) sort (section 8.20.15, page 140) and require the same amount of memory.

Errors

 Goal is a variable  instantiation error

Goal is neither a variable nor a callable term type error(callable, Goal) The predicate indicator Pred of Goal does not correspond to an existing procedure and the value of the unknown Prolog flag is error (section 8.22.1, page 148)

                                             existence error(procedure, Pred)

 Instances is neither a partial list nor a list  type error(list, Instances)

Portability

ISO predicates.

8.10 Streams

8.10.1 Introduction

A stream provides a logical view of a source/sink.

Sources and sinks: a program can output results to a sink or input data from a source. A source/sink may be a file (regular file, terminal, device,…), a constant term, a pipe, a socket,…

Associating a stream to a source/sink: to manipulate a source/sink it must be associated with a stream. This provides a logical and uniform view of the source/sink whatever its type. Once this association has been established, i.e. a stream has been created, all subsequent references to the source/sink are made by referring the stream. A stream is unidirectional: it is either an input stream or an output stream. For a classical file, the association is done by opening the file (whose name is specified as an atom) with the open/4 (section 8.10.6, page 81). GNU Prolog makes it possible to treat a Prolog constant term as a source/sink and provides built-in predicates to associate a stream to such a term (section 8.11, page 94). GNU Prolog provides operating system interface predicates defining pipes between GNU Prolog and child processes with streams associated with these pipes, e.g. popen/3 (section 8.27.21, page 170). Similarly, socket interface predicates associate streams to a socket to allow the communication, e.g. socket connect/4 (section 8.28.5, page 178).

Stream-term: a stream-term identifies a stream during a call of an input/output built-in predicate. It is created as a result of associating a stream to a source/sink (section above). A stream-term is a compound term of the form '$stream'(I) where I is an integer.

Stream aliases: any stream may be associated with a stream alias which is an atom which may be used to refer to that stream. The association can be done at open time or using add stream alias/2 (section 8.10.20, page 90). Such an association automatically ends when the stream is closed. A particular alias only refers to at most one stream at any one time. However, more than one alias can be associated with a stream. Most built-in predicates which have a stream-term as an input argument also accept a stream alias as that argument. However, built-in predicates which return a stream-term do not accept a stream alias.

Standard streams: three streams are predefined and open during the execution of every goal: the standard input stream which has the alias user input, the standard output stream which has the alias user output and the standard error stream which has the alias user error. A goal which attempts to close either standard stream succeeds, but does not close the stream.

Current streams: during execution there is a current input stream and a current output stream. By default, the current input and output streams are the standard input and output streams, but the builtin predicates set input/1 (section 8.10.4, page 80) and set output/1 (section 8.10.5, page 80) can be used to change them. When the current input stream is closed, the standard input stream becomes the current input stream. When the current output stream is closed, the standard output stream becomes the current output stream.

Text streams and binary streams: a text stream is a sequence of characters. A text stream is also regarded as a sequence of lines where each line is a possibly empty sequence of characters followed by a new line character. GNU Prolog may add or remove space characters at the ends of lines in order to conform to the conventions for representing text streams in the operating system. A binary stream is a sequence of bytes. Only a few built-in predicates can deal with binary streams, e.g. get byte/2 (section 8.13, page 101).

\underline{8.10 Streams 79}

Stream positions: the stream position of a stream identifies an absolute position of the source/sink to which the stream is connected and defines where in the source/sink the next input or output will take place. A stream position is a ground term of the form '$stream position'(I1, I2, I3, I4) where I1, I2, I3 and I4 are integers. Stream positions are used to reposition a stream (when possible) using for instance set stream position/2 (section 8.10.13, page 86).

The position end of stream: when all data of a stream S has been input S has a stream position end-of-stream. At this stream position a goal to input more data will return a specific value to indicate that end of stream has been reached (e.g. -1 for get code/2 or end of file for get char/2,…). When this terminating value has been input, the stream has a stream position past-end-of-stream.

Buffering mode: input/output on a stream can be buffered (line-buffered or block-buffered) or not buffered at all. The buffering mode can be specified at open time or using set stream buffering/2 (section 8.10.27, page 93). Line buffering is used on output streams, output data are only written to the sink when a new-line character is output (or at the close time). Block buffering is used on input or output. On input streams, when an input is requested on the source, if the buffer is empty, all available characters are read (within the limits of the size of the buffer), subsequent reads will first use the characters in the buffer. On output streams, output data are stored in the buffer and only when the buffer is full is it physically written on the sink. Thus, an output to a buffered stream may not be sent immediately to the sink connected to that stream. When it is necessary to be certain that output has been delivered, the built-in predicate flush output/1 (section 8.10.8, page 83) should be used. Finally, it is also possible to use non-buffered streams, in that case input/output are directly done on the connected source/sink. This can be useful for communication purposes (e.g. sockets) or when a precise control is needed, e.g. select/5 (section 8.27.25, page 173).

Stream mirrors: any stream may be associated with mirror streams specified at open time or using add stream mirror/2 (section 8.10.22, page 91). Then, all characters/bytes read from/written to the stream are also written on each mirror stream. The association automatically ends when either the stream or the mirror stream is closed. It is also possible to explicitly remove a mirror stream using remove stream mirror/2 (section 8.10.23, page 91).

8.10.2 current input/1

Templates

     current input(?stream)

Description

current input(Stream) unifies Stream with the stream-term identifying the current input stream.

Errors

 Stream is neither a variable nor a stream  domain error(stream, Stream)

Portability

ISO predicate.

8.10.3 current output/1

Templates

     current output(?stream)

Description current output(Stream) unifies Stream with the stream-term identifying the current output stream. Errors

 Stream is neither a variable nor a stream  domain error(stream, Stream)

Portability ISO predicate.

8.10.4 set input/1

Templates

     set input(+stream or alias)

Description set input(SorA) sets the current input stream to be the stream associated with the stream-term or alias

SorA.

Errors

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)

Portability ISO predicate.

8.10.5 set output/1

Templates

     set output(+stream or alias)

Description set output(SorA) sets the current output stream to be the stream associated with the stream-term or alias SorA. Errors

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an input stream  permission error(output, stream, SorA)

Portability

ISO predicate.

8.10.6 open/4, open/3

Templates

     open(+source sink, +io mode, -stream, +stream option list)
     open(+source sink, +io mode, -stream)

Description

open(SourceSink, Mode, Stream, Options) opens the source/sink SourceSink for input or output

as indicated by Mode and the list of stream-options Options and unifies Stream with the stream-term which is associated with this stream. See absolute file name/2 for information about the syntax of SourceSink (section 8.26.1, page 158).

Input/output modes: Mode is an atom which defines the input/output operations that may be performed the stream. Possible modes are:

• read: the source/sink is a source and must already exist. Input starts at the beginning of the source.

• write: the source/sink is a sink. If the sink already exists then it is emptied else an empty sink is created. Output starts at the beginning of that sink.

• append: the source/sink is a sink. If the sink does not exist it is created. Output starts at the end of that sink.

Stream options: Options is a list of stream options. If this list contains contradictory options, the rightmost option is the one which applies. Possible options are:

• type(text/binary): specifies whether the stream is a text stream or a binary stream. The default value is text.

• reposition(true/false): specifies whether it is possible to reposition the stream. The default value is true except if the stream cannot be repositioned (e.g. a terminal).

• eof action(error/eof code/reset): specifies the effect of attempting to input from a stream

whose stream position is past-end-of-stream:

– error: a permission error is raised signifying that no more input exists in this stream.

– eof code: the result of input is as if the stream position is end-of-stream.

– reset: the stream position is reset so that it is not past-end-of-stream, and another attempt

is made to input from it (this is useful when inputting from a terminal). The default value is eof code.

• alias(Alias): specifies that the atom Alias is to be an alias for the stream. By default no alias is attached to the stream. Several aliases can be defined for a same stream.

• mirror(Mirror): specifies the stream associated with the stream-term or alias Mirror is a mirror for the stream. By default no mirror is attached to the stream. Several mirrors can be defined for a same stream.

• buffering(none/line/block): specifies which type of buffering is used by input/output operations

on this stream:

– none: no buffering.

– line: output operations buffer data emitted until a new-line occurs

– block: input/output operations buffer data until a given number (implementation dependant)

of characters/bytes have been treated. The default value is line for a terminal (TTY), block otherwise.

open(SourceSink, Mode, Stream) is equivalent to open(SourceSink, Mode, Stream, []).

Errors

SourceSink is a variable  instantiation error
Mode is a variable  instantiation error

Options is a partial list or a list with an element E which is a variable

                                           instantiation error

Mode is neither a variable nor an atom type error(atom, Mode)

Options is neither a partial list nor a list  type error(list, Options)
Stream is not a variable  uninstantiation error(Stream)

SourceSink is neither a variable nor a source/sink

                                           domain error(source sink, SourceSink)

Mode is an atom but not an input/output mode domain error(io mode, Mode) an element E of the Options list is neither a variable nor a stream-option

                                           domain error(stream option, E)

the source/sink specified by SourceSink does not exist

                                           existence error(source sink, SourceSink)

the source/sink specified by SourceSink cannot be opened

                                           permission error(open, source sink,
                                           SourceSink)

an element E of the Options list is alias(A) and A is already associated with an open stream

                                           permission error(open, source sink,
                                           alias(A))

an element E of the Options list is mirror(M) and M is not associated with an open stream

                                           existence error(stream, M)

an element E of the Options list is mirror(M) and M is an input stream

                                           permission error(output, stream, M)

an element E of the Options list is reposition(true) and it is not possible to reposition this stream

                                           permission error(open, source sink,
                                           reposition(true))

Portability

ISO predicates. The mirror and buffering stream options are GNU Prolog extensions.

8.10.7 close/2, close/1

Templates

    close(+stream or alias, +close option list)
    close(+stream or alias)

Description

close(SorA, Options) closes the stream associated with the stream-term or alias SorA. If SorA is the standard input stream or the standard output stream close/2 simply succeeds else the associated source/sink is physically closed. If SorA is the current input stream the current input stream becomes the standard input stream user input. If SorA is the current output stream the current output stream becomes the standard output stream user output.

Close options: Options is a list of close options. For the moment only one option is available:

• force(true/false): with false, if an error occurs when trying to close the source/sink, the stream is not closed and an error (system error or resource error) is raised (but close/2 succeeds). With true, if an error occurs it is ignored and the stream is closed. The purpose of force/1 option is to allow an error handling routine to do its best to reclaim resources. The default value is false.

close(SorA) is equivalent to close(SorA, []).

Errors

 SorA is a variable  instantiation error

Options is a partial list or a list with an element E which is a variable

                                             instantiation error
 Options is neither a partial list nor a list  type error(list, Options)

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)

an element E of the Options list is neither a variable nor a close-option

                                             domain error(close option, E)
 SorA is not associated with an open stream  existence error(stream, SorA)

SorA needs a special close (section 8.11, page 94) system error(needs special close)

Portability

ISO predicates. The system error(needs special close) is a GNU Prolog extension.

8.10.8 flush output/1, flush output/0

Templates

     flush output(+stream or alias)
     flush output

Description

flush output(SorA) sends any buffered output characters/bytes to the stream.

flush output/0 applies to the current output stream.

Errors

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an input stream  permission error(output, stream, SorA)

Portability

ISO predicates.

8.10.9 current stream/1

     current stream(?stream)

Description current stream(Stream) succeeds if there exists a stream-term that unifies with Stream. This predicate is re-executable on backtracking. Errors Stream is neither a variable nor a stream-term domain error(stream, Stream) Portability GNU Prolog predicate.

8.10.10 stream property/2

Templates

     stream property(?stream, ?stream property)

Description

stream property(Stream, Property) succeeds if current stream(Stream) succeeds (section 8.10.9,

page 83) and if Property unifies with one of the properties of the stream. This predicate is re-executable on backtracking. Stream properties:

• file name(F): the name of the connected source/sink.

   • mode(M): M is the open mode (read, write, append).

• input: if it is an input stream.

• output: if it is an output stream.

• alias(A): A is an alias of the stream.

• mirror(M): M is a mirror stream of the stream.

• type(T): T is the type of the stream (text, binary).

• reposition(R): R is the reposition boolean (true, false).

   • eof action(A): A is the end-of-file action (error, eof code, reset).
   • buffering(B): B is the buffering mode (none, line, block).

• end of stream(E): E is the current end-of-stream status (not, at, past). If the stream position is end-of-stream then E is unified with at else if the stream position is past-end-of-stream then E is unified with past else E is unified with not.

• position(P): P is the stream-position term associated with the current position.

Errors

 Stream is a variable  instantiation error

Stream is neither a variable nor a stream-term domain error(stream, Stream) Property is neither a variable nor a stream property

                                             domain error(stream property, Property)
 Property = file name(E), mode(E), alias(E),
 end of stream(E), eof action(E),
 reposition(E), type(E) or buffering(E) and

E is neither a variable nor an atom

                                             type error(atom, E)

Portability

ISO predicate. The buffering/1 property is a GNU Prolog extension.

8.10.11 at end of stream/1, at end of stream/0

Templates

     at end of stream(+stream or alias)
     at end of stream

Description

at end of stream(SorA) succeeds if the stream associated with stream-term or alias SorA has a stream position end-of-stream or past-end-of-stream. This predicate can be defined using stream property/2 (section 8.10.10, page 84).

at end of stream/0 applies to the current input stream.

Errors

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)

Portability

ISO predicates. The permission error(input, stream, SorA) is a GNU Prolog extension.

8.10.12 stream position/2

Templates

     stream position(+stream or alias, ?stream position)

Description

stream position(SorA, Position) succeeds unifying Position with the stream-position term associated with the current position of the stream-term or alias SorA. This predicate can be defined using stream property/2 (section 8.10.10, page 84).

Errors

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)

Position is neither a variable nor a stream-position term

                                             domain error(stream position, Position)
 SorA is not associated with an open stream  existence error(stream, SorA)

Portability

GNU Prolog predicate.

8.10.13 set stream position/2

Templates

     set stream position(+stream or alias, +stream position)

Description

set stream position(SorA, Position) sets the position of the stream associated with the streamterm or alias SorA to Position. Position should have previously been returned by stream property/2 (section 8.10.10, page 84) or by stream position/2 (section 8.10.12, page 85).

Errors

 SorA is a variable  instantiation error
 Position is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)

Position is neither a variable nor a stream-position term

                                             domain error(stream position, Position)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA has stream property reposition(false)

                                             permission error(reposition, stream,
                                             SorA)

Portability

ISO predicate.

8.10.14 seek/4

Templates

     seek(+stream or alias, +stream seek method, +integer, ?integer)

Description

seek(SorA, Whence, Offset, NewOffset) sets the position of the stream associated with the streamterm or alias SorA to Offset according to Whence and unifies NewOffset with the new offset from the beginning of the file. seek/4 can only be used on binary streams. Whence is an atom from:

• bof: the position is set relatively to the begin of the file (Offset should be ≥ 0).

• current: the position is set relatively to the current position (Offset can be ≥ 0 or ≤ 0).

• eof: the position is set relatively to the end of the file (Offset should be ≤ 0).

This predicate is an interface to the C Unix function lseek(2).

Errors

SorA is a variable  instantiation error
Whence is a variable  instantiation error
Offset is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
Whence is neither a variable nor an atom  type error(atom, Whence)

Whence is an atom but not a valid stream seek method

                                           domain error(stream seek method, Whence)
Offset is neither a variable nor an integer  type error(integer, Offset)
NewOffset is neither a variable nor an integer  type error(integer, NewOffset)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA has stream property reposition(false)

                                           permission error(reposition, stream,
                                           SorA)

SorA is associated with a text stream

                                           permission error(reposition, text stream,
                                           SorA)

Portability

GNU Prolog predicate.

8.10.15 character count/2

Templates

    character count(+stream or alias, ?integer)

Description

character count(SorA, Count) unifies Count with the number of characters/bytes read/written on the stream associated with stream-term or alias SorA.

Errors

SorA is a variable  instantiation error

Count is neither a variable nor an integer type error(integer, Count) SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)

Portability

8.10.16 line count/2

Templates

     line count(+stream or alias, ?integer)

Description line count(SorA, Count) unifies Count with the number of lines read/written on the stream associated with the stream-term or alias SorA. This predicate can only be used on text streams. Errors

 SorA is a variable  instantiation error

Count is neither a variable nor an integer type error(integer, Count) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)

SorA is associated with a binary stream

                                             permission error(access, binary stream,
                                             SorA)

Portability GNU Prolog predicate.

8.10.17 line position/2

Templates

     line position(+stream or alias, ?integer)

Description line position(SorA, Count) unifies Count with the number of characters read/written on the current line of the stream associated with the stream-term or alias SorA. This predicate can only be used on text streams. Errors

 SorA is a variable  instantiation error

Count is neither a variable nor an integer type error(integer, Count) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)

SorA is associated with a binary stream

                                             permission error(access, binary stream,
                                             SorA)

Portability GNU Prolog predicate.

8.10.18 stream line column/3

Templates

     stream line column(+stream or alias, ?integer, ?integer)

Description

stream line column(SorA, Line, Column) unifies Line (resp. Column) with the current line number

(resp. column number) of the stream associated with the stream-term or alias SorA. This predicate can only be used on text streams. Note that Line corresponds to the value returned by line count/2 + 1 (section 8.10.16, page 88) and Column to the value returned by line position/2 + 1 (section 8.10.17, page 88).

Errors

 SorA is a variable  instantiation error

Line is neither a variable nor an integer type error(integer, Line)

 Column is neither a variable nor an integer  type error(integer, Column)

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)

SorA is associated with a binary stream

                                             permission error(access, binary stream,
                                             SorA)

Portability

GNU Prolog predicate.

8.10.19 set stream line column/3

Templates

     set stream line column(+stream or alias, +integer, +integer)

Description

set stream line column(SorA, Line, Column) sets the stream position of the stream associated with the stream-term or alias SorA according to the line number Line and the column number Column. This predicate can only be used on text streams. It first repositions the stream to the beginning of the file and then reads character by character until the required position is reached.

SorA is a variable  instantiation error
Line is a variable  instantiation error
Column is a variable  instantiation error

Line is neither a variable nor an integer type error(integer, Line)

Column is neither a variable nor an integer  type error(integer, Column)

SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)

SorA is associated with a binary stream

                                           permission error(reposition,
                                           binary stream, SorA)
SorA has stream property reposition(false)

                                           permission error(reposition, stream,
                                           SorA)

Portability

GNU Prolog predicate.

8.10.20 add stream alias/2

Templates

    add stream alias(+stream or alias, +atom)

Description

add stream alias(SorA, Alias) adds Alias as a new alias to the stream associated with the streamterm or alias SorA.

Errors

SorA is a variable  instantiation error
Alias is a variable  instantiation error

Alias is neither a variable nor an atom type error(atom, Alias) SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)

Alias is already associated with an open stream

                                           permission error(add alias, source sink,
                                           alias(Alias))

Portability

GNU Prolog predicate.

8.10.21 current alias/2

Templates

    current alias(?stream, ?atom)

Description

current alias(Stream, Alias) succeeds if current stream(Stream) succeeds (section 8.10.9, page 83)

and if Alias unifies with one of the aliases of the stream. It can be defined using stream property/2 (section 8.10.10, page 84). This predicate is re-executable on backtracking.

Errors

Stream is neither a variable nor a stream-term domain error(stream, Stream) Alias is neither a variable nor an atom type error(atom, Alias)

Portability

GNU Prolog predicate.

8.10.22 add stream mirror/2

Templates

     add stream mirror(+stream or alias, +stream or alias)

Description

add stream mirror(SorA, Mirror) adds the stream associated with the stream-term or alias Mirror as a new mirror to the stream associated with the stream-term or alias SorA. After this, all characters (or bytes) read from (or written to) SorA are also written to Mirror. This mirroring occurs until Mirror is explicitly removed using remove stream mirror/2 (section 8.10.23, page 91) or implicitly when Mirror is closed. Several mirror streams can be associated with a same stream. If Mirror represents the same stream as SorA or if Mirror is already a mirror for SorA, no mirror is added.

Errors

 SorA is a variable  instantiation error
 Mirror is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)

Mirror is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, Mirror)
 SorA is not associated with an open stream  existence error(stream, SorA)
 Mirror is not associated with an open stream  existence error(stream, Mirror)
 Mirror is an input stream  permission error(output, stream, Mirror)

Portability

GNU Prolog predicate.

8.10.23 remove stream mirror/2

Templates

     remove stream mirror(+stream or alias, +stream or alias)

Description

remove stream mirror(SorA, Mirror) removes the stream associated with the stream-term or alias Mirror from the list of mirrors of the stream associated with the stream-term or alias SorA. This predicate fails if Mirror is not a mirror stream for SorA.

Errors

 SorA is a variable  instantiation error
 Mirror is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)

Mirror is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, Mirror)
 SorA is not associated with an open stream  existence error(stream, SorA)
 Mirror is not associated with an open stream  existence error(stream, Mirror)

Portability GNU Prolog predicate.

8.10.24 current mirror/2

Templates

     current mirror(?stream, ?stream)

Description

current mirror(Stream, M) succeeds if current stream(Stream) succeeds (section 8.10.9, page 83)

and if M unifies with one of the mirrors of the stream. It can be defined using stream property/2 (section 8.10.10, page 84). This predicate is re-executable on backtracking. Errors Stream is neither a variable nor a stream-term domain error(stream, Stream) M is neither a variable nor a stream-term domain error(stream, M) Portability GNU Prolog predicate.

8.10.25 set stream type/2

Templates

     set stream type(+stream or alias, +atom)

Description set stream type(SorA, Type) updates the type associated with stream-term or alias SorA. The value of Type is an atom in text or binary as for open/4 (section 8.10.6, page 81). The type of a stream can only be changed before any input/output operation is executed. Errors

 SorA is a variable  instantiation error
 Type is a variable  instantiation error
 Type is neither a variable nor a valid type  domain error(stream type, Type)

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)

An I/O operation has already been executed on

 SorA

                                             permission error(modify, stream, SorA)

Portability

GNU Prolog predicate.

8.10.26 set stream eof action/2

Templates

     set stream eof action(+stream or alias, +atom)

Description

set stream eof action(SorA, Action) updates the eof action option associated with the stream-

term or alias SorA. The value of Action is one of the atoms error, eof code, reset as for open/4 (section 8.10.6, page 81).

Errors

 SorA is a variable  instantiation error
 Action is a variable  instantiation error

Action is neither a variable nor a valid eof action domain error(eof action, Action) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(modify, stream, SorA)

Portability

GNU Prolog predicate.

8.10.27 set stream buffering/2

Templates

     set stream buffering(+stream or alias, +atom)

Description

set stream buffering(SorA, Buffering) updates the buffering mode associated with the stream-term or alias SorA. The value of Buffering is one of the atoms none, line or block as for open/4 (section 8.10.6, page 81). This predicate may only be used after opening a stream and before any other operations have been performed on it.

 SorA is a variable  instantiation error
 Buffering is a variable  instantiation error

Buffering is neither a variable nor a valid buffering mode

                                             domain error(buffering mode, Buffering)

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)

Portability

GNU Prolog predicate.

8.11 Constant term streams

8.11.1 Introduction

Constant term streams allow the user to consider a constant term (atom, character list or character code list) as a source/sink by associating to them a stream. Reading from a constant term stream will deliver the characters of the constant term as if they had been read from a standard file. Characters written on a constant term stream are stored to form the final constant term when the stream is closed. The built-in predicates described in this section allow the user to open and close a constant term stream for input or output. However, very often, a constant term stream is created to be only read or written once and then closed. To avoid the creation and the destruction of such a stream, GNU Prolog offers several built-in predicates to perform single input/output from/to constant terms (section 8.15, page 117).

8.11.2 open input atom stream/2, open input chars stream/2,
        open input codes stream/2

Templates

     open input atom stream(+atom, -stream)
     open input chars stream(+character list, -stream)
     open input codes stream(+character code list, -stream)

Description

open input atom stream(Atom, Stream) unifies Stream with the stream-term which is associated with a new input text-stream whose data are the characters of Atom.

open input chars stream(Chars, Stream) is similar to open input atom stream/2 except that data

are the content of the character list Chars.

open input codes stream(Codes, Stream) is similar to open input atom stream/2 except that data

are the content of the character code list Codes.

Stream is not a variable  uninstantiation error(Stream)
Atom is a variable  instantiation error

Chars is a partial list or a list with an element E which is a variable

                                           instantiation error

Codes is a partial list or a list with an element E which is a variable

                                           instantiation error

Atom is neither a variable nor a an atom type error(atom, Atom) Chars is neither a partial list nor a list type error(list, Chars) Codes is neither a partial list nor a list type error(list, Codes) an element E of the Chars list is neither a variable nor a character

                                           type error(character, E)

an element E of the Codes list is neither a variable nor an integer

                                           type error(integer, E)

an element E of the Codes list is an integer but not a character code

                                           representation error(character code)

Portability

GNU Prolog predicates.

8.11.3 close input atom stream/1, close input chars stream/1,
       close input codes stream/1

Templates

    close input atom stream(+stream or alias)
    close input chars stream(+stream or alias)
    close input codes stream(+stream or alias)

Description

close input atom stream(SorA) closes the constant term stream associated with the stream-term or alias SorA. SorA must a stream open with open input atom stream/2 (section 8.11.1, page 94).

close input chars stream(SorA) acts similarly for a character list stream.

close input codes stream(SorA) acts similarly for a character code list stream.

Errors

SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an output stream  permission error(close, stream, SorA)

SorA is a stream-term or alias but does not refer to a constant term stream.

                                           domain error(term stream or alias, SorA)

Portability

8.11.4 open output atom stream/1, open output chars stream/1,
        open output codes stream/1

Templates

     open output atom stream(-stream)
     open output chars stream(-stream)
     open output codes stream(-stream)

Description

open output atom stream(Stream) unifies Stream with the stream-term which is associated with a new output text-stream. All characters written to this stream are collected and will be returned as an atom when the stream is closed by close output atom stream/2 (section 8.11.5, page 96).

open output chars stream(Stream) is similar to open output atom stream/1 except that the result

will be a character list.

open output codes stream(Stream) is similar to open output atom stream/1 except that the result

will be a character code list.

Errors

 Stream is not a variable  uninstantiation error(Stream)

Portability

GNU Prolog predicates.

8.11.5 close output atom stream/2, close output chars stream/2,
        close output codes stream/2

Templates

     close output atom stream(+stream or alias, ?atom)
     close output chars stream(+stream or alias, ?character list)
     close output codes stream(+stream or alias, ?character code list)

Description

close output atom stream(SorA, Atom) closes the constant term stream associated with the streamterm or alias SorA. SorA must be associated with a stream open with open output atom stream/1 (section 8.11.4, page 96). Atom is unified with an atom formed with all characters written on the stream.

close output chars stream(SorA, Chars) acts similarly for a character list stream.

close output codes stream(SorA, Codes) acts similarly for a character code list stream.

SorA is a variable  instantiation error

Atom is neither a variable nor an atom type error(atom, Atom) Chars is neither a partial list nor a list type error(list, Chars) Codes is neither a partial list nor a list type error(list, Codes) an element E of the Chars list is neither a variable nor a character

                                           type error(character, E)

an element E of the Codes list is neither a variable nor an integer

                                           type error(integer, E)

an element E of the Codes list is an integer but not a character code

                                           representation error(character code)

SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an input stream  permission error(close, stream, SorA)

SorA is a stream-term or alias but does not refer to a constant term stream

                                           domain error(term stream or alias, SorA)

Portability

GNU Prolog predicates.

8.12 Character input/output

These built-in predicates enable a single character or character code to be input from and output to a text stream. The atom end of file is returned as character to indicate the end-of-file. -1 is returned as character code to indicate the end-of-file.

8.12.1 get char/2, get char/1, get code/1, get code/2

Templates

    get char(+stream or alias, ?in character)
    get char(?in character)
    get code(+stream or alias, ?in character code)
    get code(?in character code)

Description

get char(SorA, Char) succeeds if Char unifies with the next character read from the stream associated with the stream-term or alias SorA.

get code/2 is similar to get char/2 but deals with character codes.

get char/1 and get code/1 apply to the current input stream.

 SorA is a variable  instantiation error

Char is neither a variable nor an in-character type error(in character, Char) Code is neither a variable nor an integer type error(integer, Code) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)

SorA is associated with a binary stream

                                             permission error(input, binary stream,
                                             SorA)

SorA has stream properties

 end of stream(past) and eof action(error)

                                             permission error(input,
                                             past end of stream, SorA)

The entity input from the stream is not a character

                                             representation error(character)
 Code is an integer but not an in-character code  representation error(in character code)

Portability

ISO predicates.

8.12.2 get key/2, get key/1 get key no echo/2, get key no echo/1

Templates

     get key(+stream or alias, ?integer)
     get key(?integer)
     get key no echo(+stream or alias, ?integer)
     get key no echo(?integer)

Description

get key(SorA, Code) succeeds if Code unifies with the character code of the next key read from the stream associated with the stream-term or alias SorA. It is intended to read a single key from the keyboard (thus SorA should refer to current input stream). No buffering is performed (a character is read as soon as available) and function keys can also be read (in that case, Code is an integer > 255). The read character is echoed if it is printable.

This facility is only possible if the linedit facility has been installed (section 4.2.6, page 18) otherwise get key/2 behaves similarly to get code/2 (section 8.12.1, page 97) (the code of the first character is returned) but also pumps remaining characters until a character < space (0x20) is read (in particular RETURN). The same behavior occurs if SorA does not refer to the current input stream or if this stream is not attached to a terminal.

get key no echo/2 behaves similarly to get key/2 except that the read character is not echoed.

get key/1 and get key no echo/1 apply to the current input stream.

SorA is a variable  instantiation error

Code is neither a variable nor an integer type error(integer, Code) SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an output stream  permission error(input, stream, SorA)

SorA is associated with a binary stream

                                           permission error(input, binary stream,
                                           SorA)

SorA has stream properties

end of stream(past) and eof action(error)

                                           permission error(input,
                                           past end of stream, SorA)

Portability

GNU Prolog predicates.

8.12.3 peek char/2, peek char/1, peek code/1, peek code/2

Templates

    peek char(+stream or alias, ?in character)
    peek char(?in character)
    peek code(+stream or alias, ?in character code)
    peek code(?in character code)

Description

peek char(SorA, Char) succeeds if Char unifies with the next character that will be read from the stream associated with the stream-term or alias SorA. The character is not read.

peek code/2 is similar to peek char/2 but deals with character codes.

peek char/1 and peek code/1 apply to the current input stream.

Errors

SorA is a variable  instantiation error

Char is neither a variable nor an in-character type error(in character, Char) Code is neither a variable nor an integer type error(integer, Code) SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an output stream  permission error(input, stream, SorA)

SorA is associated with a binary stream

                                           permission error(input, binary stream,
                                           SorA)

SorA has stream properties

end of stream(past) and eof action(error)

                                           permission error(input,
                                           past end of stream, SorA)

The entity input from the stream is not a character

                                           representation error(character)
Code is an integer but not an in-character code  representation error(in character code)

Portability

ISO predicates.

8.12.4 unget char/2, unget char/1, unget code/2, unget code/1

Templates

     unget char(+stream or alias, +character)
     unget char(+character)
     unget code(+stream or alias, +character code)
     unget code(+character code)

Description

unget char(SorA, Char) pushes back Char onto the stream associated with the stream-term or alias SorA. Char will be the next character read by get char/2. The maximum number of characters that can be cumulatively pushed back is given by the max unget Prolog flag (section 8.22.1, page 148).

unget code/2 is similar to unget char/2 but deals with character codes.

unget char/1 and unget code/1 apply to the current input stream.

Errors

 SorA is a variable  instantiation error
 Char is a variable  instantiation error
 Code is a variable  instantiation error

Char is neither a variable nor a character type error(character, Char) Code is neither a variable nor an integer type error(integer, Code) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)

SorA is associated with a binary stream

                                             permission error(input, binary stream,
                                             SorA)
 Code is an integer but not a character code  representation error(character code)

Portability

GNU Prolog predicates.

8.12.5 put char/2, put char/1, put code/1, put code/2, nl/1, nl/0

Templates

     put char(+stream or alias, +character)
     put char(+character)
     put code(+stream or alias, +character code)
     put code(+character code)
     nl(+stream or alias)
     nl

Description

put char(SorA, Char) writes Char onto the stream associated with the stream-term or alias SorA.

put code/2 is similar to put char/2 but deals with character codes. nl(SorA) writes a new-line character onto the stream associated with the stream-term or alias SorA. This is equivalent to put char(SorA, '\n'). put char/1, put code/1 and nl/0 apply to the current output stream. Errors

 SorA is a variable  instantiation error
 Char is a variable  instantiation error
 Code is a variable  instantiation error

Char is neither a variable nor a character type error(character, Char) Code is neither a variable nor an integer type error(integer, Code) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an input stream  permission error(output, stream, SorA)

SorA is associated with a binary stream

                                             permission error(output, binary stream,
                                             SorA)
 Code is an integer but not a character code  representation error(character code)

Portability ISO predicates.

8.13 Byte input/output

These built-in predicates enable a single byte to be input from and output to a binary stream. -1 is returned to indicate the end-of-file.

8.13.1 get byte/2, get byte/1

Templates

     get byte(+stream or alias, ?in byte)
     get byte(?in byte)

Description get byte(SorA, Byte) succeeds if Byte unifies with the next byte read from the stream associated with the stream-term or alias SorA. get byte/1 applies to the current input stream. Errors

SorA is a variable  instantiation error

Byte is neither a variable nor an in-byte type error(in byte, Byte) SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an output stream  permission error(input, stream, SorA)

SorA is associated with a text stream

                                           permission error(input, text stream,
                                           SorA)

SorA has stream properties

end of stream(past) and eof action(error)

                                           permission error(input,
                                           past end of stream, SorA)

Portability

ISO predicates.

8.13.2 peek byte/2, peek byte/1

Templates

    peek byte(+stream or alias, ?in byte)
    peek byte(?in byte)

Description

peek byte(SorA, Byte) succeeds if Byte unifies with the next byte that will be read from the stream associated with the stream-term or alias SorA. The byte is not read.

peek byte/1 applies to the current input stream.

Errors

SorA is a variable  instantiation error

Byte is neither a variable nor an in-byte type error(in byte, Byte) SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an output stream  permission error(input, stream, SorA)

SorA is associated with a text stream

                                           permission error(input, text stream,
                                           SorA)

SorA has stream properties

end of stream(past) and eof action(error)

                                           permission error(input,
                                           past end of stream, SorA)

Portability

ISO predicates.

8.13.3 unget byte/2, unget byte/1

Templates

    unget byte(+stream or alias, +byte)
    unget byte(+byte)

Description unget byte(SorA, Byte) pushes back Byte onto the stream associated with the stream-term or alias SorA. Byte will be the next byte read by get byte/2. The maximum number of bytes that can be successively pushed back is given by the max unget Prolog flag (section 8.22.1, page 148). unget byte/1 applies to the current input stream. Errors

 SorA is a variable  instantiation error
 Byte is a variable  instantiation error

Byte is neither a variable nor a byte type error(byte, Byte) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)

SorA is associated with a text stream

                                             permission error(input, text stream,
                                             SorA)

Portability GNU Prolog predicates.

8.13.4 put byte/2, put byte/1

Templates

     put byte(+stream or alias, +byte)
     put byte(+byte)

Description put byte(SorA, Byte) writes Byte onto the stream associated with the stream-term or alias SorA. put byte/1 applies to the current output stream. Errors

 SorA is a variable  instantiation error
 Byte is a variable  instantiation error

Byte is neither a variable nor a byte type error(byte, Byte) SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(output, stream, SorA)

SorA is associated with a text stream

                                             permission error(output, text stream,
                                             SorA)

Portability GNU Prolog predicates.

8.14 Term input/output

These built-in predicates enable a Prolog term to be input from or output to a text stream. The atom end of file is returned as term to indicate the end-of-file. The syntax of such terms can also be altered by changing the operators (section 8.14.10, page 113), and making some characters equivalent to others (section 8.14.12, page 116) if the char conversion Prolog flag is on (section 8.22.1, page 148). Double quoted tokens will be returned as an atom or a character list or a character code list depending on the value of the double quotes Prolog flag (section 8.22.1, page 148). Similarly, back quoted tokens are returned depending on the value of the back quotes Prolog flag.

8.14.1 read term/3, read term/2, read/2, read/1

Templates

     read term(+stream or alias, ?term, +read option list)
     read term(?term, +read option list)
     read(+stream or alias, ?term)
     read(?term)

Description

read term(SorA, Term, Options) is true if Term unifies with the next term read from the stream associated with the stream-term or alias SorA according to the options given by Options.

Read options: Options is a list of read options. If this list contains contradictory options, the rightmost option is the one which applies. Possible options are:

• variables(VL): VL is unified with the list of all variables of the input term, in left-to-right traversal order. Anonymous variables are included in the list VL.

• variable names(VNL): VNL is unified with the list of pairs Name = Var where Var is a named variable of the term and Name is the atom associated with the name of Var. Anonymous variables are not included in the list VNL. The pairs appear in left-to-right traversal order of their Var in the term.

• singletons(SL): SL is unified with the list of pairs Name = Var where Var is a named variable which occurs only once in the term and Name is the atom associated to the name of Var. Anonymous variables are not included in the list SL.

• syntax error(error/warning/fail): specifies the effect of a syntax error:

       – error: a syntax error is raised.

– warning: a warning message is displayed and the predicate fails.

– fail: the predicate quietly fails. The default value is the value of the syntax error Prolog flag (section 8.22.1, page 148).

• end of term(dot/eof): specifies the end-of-term delimiter: dot is the classical full-stop delimiter (a dot followed with a layout character), eof is the end-of-file delimiter. This option is useful for predicates like read term from atom/3 (section 8.15.1, page 117) to avoid to add a terminal dot at the end of the atom. The default value is dot.

read(SorA, Term) is equivalent to read term(SorA, Term, []).

read term/2 and read/1 apply to the current input stream.

 SorA is a variable  instantiation error

Options is a partial list or a list with an element E which is a variable

                                             instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 Options is neither a partial list nor a list  type error(list, Options)

an element E of the Options list is neither a variable nor a valid read option

                                             domain error(read option, E)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)

SorA is associated with a binary stream

                                             permission error(input, binary stream,
                                             SorA)

SorA has stream properties

 end of stream(past) and eof action(error)

                                             permission error(input,
                                             past end of stream, SorA)

a syntax error occurs and the value of the

 syntax error Prolog flag is error

(section 8.22.1, page 148)

                                             syntax error(atom explaining the error)

Portability

ISO predicates. The ISO reference raises a representation error(Flag) where Flag is max arity, max integer, or min integer when the read term breaches an implementation defined limit specified by Flag. GNU Prolog detects neither min integer nor max integer violation and treats a max arity violation as a syntax error. The read options syntax error and end of term are GNU Prolog extensions.

8.14.2 read atom/2, read atom/1, read integer/2, read integer/1,
        read number/2, read number/1

Templates

     read atom(+stream or alias, ?atom)
     read atom(?atom)
     read integer(+stream or alias, ?integer)
     read integer(?integer)
     read number(+stream or alias, ?number)
     read number(?number)

Description

read atom(SorA, Atom) succeeds if Atom unifies with the next atom read from the stream associated with the stream-term or alias SorA.

read integer(SorA, Integer) succeeds if Integer unifies with the next integer read from the stream associated with the stream-term or alias SorA.

read number(SorA, Number) succeeds if Number unifies with the next number (integer or floating point number) read from the stream associated with the stream-term or alias SorA.

read atom/1, read integer/1 and read number/1 apply to the current input stream.

SorA is a variable  instantiation error

Atom is neither a variable nor an atom type error(atom, Atom)

Integer is neither a variable nor an integer  type error(integer, Integer)
Number is neither a variable nor a number  type error(number, Number)

SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an output stream  permission error(input, stream, SorA)

SorA is associated with a binary stream

                                           permission error(input, binary stream,
                                           SorA)

SorA has stream properties

end of stream(past) and eof action(error)

                                           permission error(input,
                                           past end of stream, SorA)

a syntax error occurs and the value of the

syntax error Prolog flag is error

(section 8.22.1, page 148)

                                           syntax error(atom explaining the error)

Portability

GNU Prolog predicates.

8.14.3 read token/2, read token/1

Templates

    read token(+stream or alias, ?nonvar)
    read token(?nonvar)

Description

read token(SorA, Token) succeeds if Token unifies with the encoding of the next Prolog token read from the stream associated with stream-term or alias SorA.

Token encoding:

• var(A): a variable is read whose name is the atom A.

• an atom A: an atom A is read.

• integer N: an integer N is read.

• floating point number N: a floating point number N is read.

• string(A): a string (double quoted item) is read whose characters forms the atom A.

• punct(P): a punctuation character P is read (P is a one-character atom in ()[]{|}, the atom

    full stop or the atom end of file).

• back quotes(A): a back quoted item is read whose characters forms the atom A.

• extended(A): an extended character A (an atom) is read.

As for read term/3, the behavior of read token/2 can be affected by some Prolog flags (section 8.14, page 104).

read token/1 applies to the current input stream.

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an output stream  permission error(input, stream, SorA)

SorA is associated with a binary stream

                                             permission error(input, binary stream,
                                             SorA)

SorA has stream properties

 end of stream(past) and eof action(error)

                                             permission error(input,
                                             past end of stream, SorA)

a syntax error occurs and the value of the

 syntax error Prolog flag is error

(section 8.22.1, page 148)

                                             syntax error(atom explaining the error)

Portability

GNU Prolog predicates.

8.14.4 syntax error info/4

Templates

     syntax error info(?atom, ?integer, ?integer, ?atom)

Description

syntax error info(FileName, Line, Column, Error) returns the information associated with the last

syntax error. Line is the line number of the error, Column is the column number of the error and Error is an atom explaining the error.

Errors

 FileName is neither a variable nor an atom  type error(atom, FileName)

Line is neither a variable nor an integer type error(integer, Line)

 Column is neither a variable nor an integer  type error(integer, Column)

Error is neither a variable nor an atom type error(atom, Error)

Portability

GNU Prolog predicate.

8.14.5 last read start line column/2

Templates

     last read start line column(?integer, ?integer)

Description

last read start line column(Line, Column) unifies Line and Column with the line number and the

column number associated with the start of the last read predicate. This predicate can be used after calling one of the following predicates: read term/3, read term/2, read/2, read/1 (section 8.14.1, page 104),

read atom/2, read atom/1, read integer/2, read integer/1, read number/2, read number/1 (sec-

tion 8.14.2, page 105) or read token/2, read token/1 (section 8.14.3, page 106).

Errors

Line is neither a variable nor an integer type error(integer, Line)

Column is neither a variable nor an integer  type error(integer, Column)

Portability

GNU Prolog predicate.

8.14.6 write term/3, write term/2, write/2, write/1, writeq/2, writeq/1,
   write canonical/2, write canonical/1, display/2, display/1, print/2,
   print/1

Templates

write term(+stream or alias, ?term, +write option list)
write term(?term, +write option list)
write(+stream or alias, ?term)
write(?term)
writeq(+stream or alias, ?term)
writeq(?term)
write canonical(+stream or alias, ?term)
write canonical(?term)
display(+stream or alias, ?term)
display(?term)
print(+stream or alias, ?term)
print(?term)

Description

write term(SorA, Term, Options) writes Term to the stream associated with the stream-term or alias SorA according to the options given by Options.

Write options: Options is a list of write options. If this list contains contradictory options, the rightmost option is the one which applies. Possible options are:

• quoted(true/false): if true each atom and functor is quoted if this would be necessary for the term to be input by read term/3. If false no extra quotes are written. The default value is false.

• ignore ops(true/false): if true each compound term is output in functional notation (neither operator notation nor list notation is used). If false operator and list notations are used. The default value is false.

• numbervars(true/false): if true a term of the form '$VAR'(N), where N is an integer, is output as a variable name (see below). If false such a term is output normally (according to the other options). The default value is false.

• namevars(true/false): if true a term of the form '$VARNAME'(Name), where Name is an atom

respecting the syntax of variable names, is output as a variable name (see below). If false such a term is output normally (according to the other options). The default value is false.

• variable names(VNL): VNL is a list of pairs Name = Var where Var is a variable and Name is the atom associated with the name of Var. Each variable Var is written as the atom Name (with quoted(false)) iff a term Name = Var is an element of the list VNL. If several pairs exist for the same variable name the first one applies.

• space args(true/false): if true an extra space character is emitted after each comma separating the arguments of a compound term in functional notation or of a list. If false no extra space is emitted. The default value is false.

   • portrayed(true/false): if true and if there exists a predicate portray/1, write term/3 acts

as follows: if Term is a variable it is simply written. If Term is non-variable then it is passed to portray/1. If this succeeds then it is assumed that Term has been output. Otherwise write term/3 outputs the principal functor of Term (Term itself if it is atomic) according to other options and recursively calls portray/1 on the components of Term (if it is a compound term). With ignore ops(false) a list is first passed to portray/1 and only if this call fails each element of the list is passed to portray/1 (thus every sub-list is not passed). The default value is false.

• max depth(N): controls the depth of output for compound terms. N is an integer specifying the depth. The output of a term whose depth is greater than N gives rise to the output of ... (3 dots). By default there is no depth limit.

• priority(N): specifies the starting priority to output the term. This option controls if Term should be enclosed in brackets. N is a positive integer ≤ 1200. By default N = 1200.

Variable numbering: when the numbervars(true) option is passed to write term/3 any term of the form '$VAR'(N) where N is an integer is output as a variable name consisting of a capital letter possibly followed by an integer. The capital letter is the (I+1)th letter of the alphabet and the integer is J, where I = N mod 26 and J = N // 26. The integer J is omitted if it is zero. For example:

'$VAR'(0) is written as A
'$VAR'(1) is written as B

     ...

'$VAR'(25) is written as Z
'$VAR'(26) is written as A1
'$VAR'(27) is written as B1

Variable naming: when the namevars(true) option is passed to write term/3 any term of the form '$VARNAME'(Name) where Name is an atom is output as a variable name consisting of the characters Name. For example: '$VARNAME'('A') is written as A (even in the presence of the quoted(true) option).

write(SorA, Term) is equivalent to write term(SorA, Term, [numbervars(true),
namevars(true)]).
writeq(SorA, Term) is equivalent to write term(SorA, Term, [quoted(true),
numbervars(true), namevars(true)]).
write canonical(SorA, Term) is equivalent to write term(SorA, Term, [quoted(true),
ignore ops(true), numbervars(false), namevars(false)]).
display(SorA, Term) is equivalent to write term(SorA, Term, [ignore ops(true),
numbervars(false), namevars(false)]).
print(SorA, Term) is equivalent to write term(SorA, Term, [numbervars(false),
portrayed(true)]).
write term/2, write/1, writeq/1, write canonical/1, display/1 and print/1 apply to the current

output stream.

 SorA is a variable  instantiation error

Options is a partial list or a list with an element E which is a variable

                                             instantiation error
 Options is neither a partial list nor a list  type error(list, Options)

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)

an element E of the Options list is neither a variable nor a valid write-option

                                             domain error(write option, E)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an input stream  permission error(output, stream, SorA)

SorA is associated with a binary stream

                                             permission error(output, binary stream,
                                             SorA)

Portability

ISO predicates except display/1-2 and print/1-2 that are GNU Prolog predicates. namevars, variable names space args, portrayed, max depth and priority options are GNU Prolog extensions.

8.14.7 format/3, format/2

Templates

     format(+stream or alias, +character code list or atom, +list)
     format(+character code list or atom, +list)

Description

format(SorA, Format, Arguments) writes the Format string replacing each format control sequence F by the corresponding element of Arguments (formatted according to F) to the stream associated with the stream-term or alias SorA.

Format control sequences: the general format of a control sequence is '~NC'. The character C determines the type of the control sequence. N is an optional numeric argument. An alternative form of N is '*'. '*' implies that the next argument Arg in Arguments should be used as a numeric argument in the control sequence. The use of C printf() formatting sequence (beginning by the character %) is also allowed. The following control sequences are available:

Format sequence type of the argument

Description

~Na

atom print the atom without quoting. N is minimal number of characters to print using spaces on the right if needed (default: the length of the atom)

~Nc

character code print the character associated with the code. N is the number of times to print the character (default: 1)

~Nf ~Ne ~NE ~Ng ~NG

float expression pass the argument Arg and N to the C printf() function as:

if N is not specified printf("%f",Arg) else printf("%.Nf",Arg).

Similarly for ~Ne, ~NE, ~Ng and ~NG

~Nd

integer expression print the argument. N is the number of digits after the decimal point.

If N is 0 no decimal point is printed (default: 0)

~ND

integer expression identical to ~Nd except that ',' separates groups of three digits to

the left of the decimal point

~Nr

integer expression print the argument according to the radix N. 2 ≤ N ≤ 36 (default: 8).

The letters a-z denote digits > 9

~NR

integer expression identical to ~Nr except that the letters A-Z denote digits > 9

~Ns

character code list print exactly N characters (default: the length of the list)

~NS

character list print exactly N characters (default: the length of the list)

~i

term

ignore the current argument

~k

term

pass the argument to write canonical/1 (section 8.14.6, page 108)

~p

term

pass the argument to print/1 (section 8.14.6, page 108)

~q

term

pass the argument to writeq/1 (section 8.14.6, page 108)

~w

term

pass the argument to write/1 (section 8.14.6, page 108)

~~

none

print the character '~'

~Nn

none

print N new-line characters (default: 1)

~N

none

print a new-line character if not at the beginning of a line

~?

atom

use the argument as a nested format string

%F

atom, integer or float expression interface to the C function printf(3) for outputting atoms (C string), integers and floating point numbers. * are also allowed.

format/2 applies to the current output stream.

SorA is a variable  instantiation error

Format is a partial list or a list with an element E which is a variable

                                           instantiation error
Arguments is a partial list  instantiation error

Format is neither a partial list nor a list or an atom

                                           type error(list, Format)
Arguments is neither a partial list nor a list  type error(list, Arguments)

an element E of the Format list is neither a variable nor a character code

                                           representation error(character code, E)

SorA is neither a variable nor a stream-term or alias

                                           domain error(stream or alias, SorA)

an element E of Format is not a valid format control sequence

                                           domain error(format control sequence, E)

the Arguments list does not contain sufficient elements

                                           domain error(non empty list, [])

an element E of the Arguments list is a variable while a non-variable term was expected

                                           instantiation error

an element E of the Arguments list is neither variable nor an atom while an atom was expected

                                           type error(atom, E)

an element E of the Arguments cannot be evaluated as an arithmetic expression while an integer or a floating point number was expected

an arithmetic error (section 8.6.1, page 66)

an element E of the Arguments list is neither variable nor character code while a character code was expected

                                           representation error(character code, E)
SorA is not associated with an open stream  existence error(stream, SorA)
SorA is an input stream  permission error(output, stream, SorA)

SorA is associated with a binary stream

                                           permission error(output, binary stream,
                                           SorA)

Portability

GNU Prolog predicates.

8.14.8 portray clause/2, portray clause/1

Templates

    portray clause(+stream or alias, +clause)
    portray clause(+clause)

Description

portray clause(SorA, Clause) pretty prints Clause to the stream associated with the stream-term or alias SorA. portray clause/2 uses the variable binding predicates name singleton vars/1 (section 8.5.1, page 63) and numbervars/1 (section 8.5.3, page 64). This predicate is used by listing/1 (section 8.23.3, page 154).

portray clause/1 applies to the current output stream.

 Clause is a variable  instantiation error

Clause is neither a variable nor a callable term type error(callable, Clause)

 SorA is a variable  instantiation error

SorA is neither a variable nor a stream-term or alias

                                             domain error(stream or alias, SorA)
 SorA is not associated with an open stream  existence error(stream, SorA)
 SorA is an input stream  permission error(output, stream, SorA)

SorA is associated with a binary stream

                                             permission error(output, binary stream,
                                             SorA)

Portability

GNU Prolog predicates.

8.14.9 get print stream/1

Templates

     get print stream(?stream)

Description

get print stream(Stream) unifies Stream with the stream-term associated with the output stream used by print/2 (section 8.14.6, page 108). The purpose of this predicate is to allow a user-defined portray/1 predicate to identify the output stream in use.

Errors

Stream is neither a variable nor a stream-term domain error(stream, Stream)

Portability

GNU Prolog predicate.

8.14.10 op/3

Templates

     op(+integer, +operator specifier, +atom or atom list)

Description

op(Priority, OpSpecifier, Operator) alters the operator table. Operator is declared as an operator with properties defined by specifier OpSpecifier and Priority. Priority must be an integer ≥ 0 and ≤ 1200. If Priority is 0 then the operator properties of Operator (if any) are canceled. Operator may also be a list of atoms in which case all of them are declared to be operators. In general, operators can be removed from the operator table and their priority or specifier can be changed. However, it is an error to attempt to change the ',' operator from its initial status. An atom can have multiple operator definitions (e.g. prefix and infix like +) however an atom cannot have both an infix and a postfix operator definitions.

Operator specifiers: the following specifiers are available:

Specifier Type Associativity
fx  prefix  no
fy  prefix  yes
xf  postfix  no
yf  postfix  yes
xfx  infix  no
yfx  infix  left
xfy  infix  right

Prolog predefined operators:

Priority Specifier Operators

1200    xfx    :-  -->
1200     fx     :-
1105    xfy    |
1100    xfy    ;
1050    xfy    -> *->
1000    xfy    ,
900     fy     \+
700    xfx    {=  \=  =..   ==  \==  @<  @=<  @>  @>=  is  =:=  =\=  <  =< \atop >  >=}
600    xfy    :
500    yfx    +  -  /\  \/

400    yfx    *  /  //  rem  mod  div  <<  >>

200    xfx    **  ^
200     fy     +  -  \

FD predefined operators:

Priority Specifier Operators

750    xfy    #<=>  #\<=>
740    xfy    #==>  #\==>
730    xfy    ##  #\/  #\\/
720    yfx    #/\  #\/\
710     fy     #\
700    xfx    #=  #\=  #<  #=<  #>  #>=  #=#  #\=#  #<#  #=<#  #>#  #>=#
500    yfx    +  -
400    yfx    *  /  //  rem
200    xfy    **
200     fy     +  -

Priority is a variable  instantiation error
OpSpecifier is a variable  instantiation error

Operator is a partial list or a list with an element E which is a variable

                                           instantiation error
Priority is neither a variable nor an integer  type error(integer, Priority)
OpSpecifier is neither a variable nor an atom  type error(atom, OpSpecifier)

Operator is neither a partial list nor a list nor an atom

                                           type error(list, Operator)

an element E of the Operator list is neither a variable nor an atom

                                           type error(atom, E)
Priority is an integer not ≥ 0 and ≤ 1200  domain error(operator priority, Priority)

OpSpecifier is not a valid operator specifier

                                           domain error(operator specifier,
                                           OpSpecifier)

Operator (or an element of the Operator list) is

','

                                           permission error(modify, operator, ',')
OpSpecifier is a specifier such that Operator

would have a postfix and an infix definition.

                                           permission error(create, operator,
                                           Operator)

Operator (or an element of the Operator list) is | and it would have a prefix or a postfix definition or its Priority would be ≤ 1100.

                                           permission error(create, operator, '|')

Operator (or an element of the Operator list) is

[] or {}.

                                           permission error(create, operator,
                                           Operator)

Portability

ISO predicate.

The ISO reference implies that if a program calls current op/3, then modifies an operator definition by calling op/3 and backtracks into the call to current op/3, then the changes are guaranteed not to affect that current op/3 goal. This is not guaranteed by GNU Prolog.

8.14.11 current op/3

Templates

    current op(?integer, ?operator specifier, ?atom)

Description

current op(Priority, OpSpecifier, Operator) succeeds if Operator is an operator with properties

defined by specifier OpSpecifier and Priority. This predicate is re-executable on backtracking.

Errors

Priority is neither a variable nor an operator priority

                                           domain error(operator priority, Priority)

OpSpecifier is neither a variable nor an operator specifier

                                           domain error(operator specifier,
                                           OpSpecifier)
Operator is neither a variable nor an atom  type error(atom, Operator)

Portability

ISO predicate.

8.14.12 char conversion/2

Templates

     char conversion(+character, +character)

Description

char conversion(InChar, OutChar) alters the character-conversion mapping. This mapping is used by the following read predicates: read term/3 (section 8.14.1, page 104), read atom/2, read integer/2, read number/2 (section 8.14.2, page 105) and read token/2 (section 8.14.3, page 106) to replace any occurrence of a character InChar by OutChar. However the conversion mechanism should have been previously activated by switching on the char conversion Prolog flag (section 8.22.1, page 148). When InChar and OutChar are the same, the effect is to remove any conversion of a character InChar.

Note that the single character read predicates (e.g. get char/2) never do character conversion. If such behavior is required, it must be explicitly done using current char conversion/2 (section 8.14.13, page 116).

Errors

 InChar is a variable  instantiation error
 OutChar is a variable  instantiation error
 InChar is neither a variable nor a character  type error(character, InChar)
 OutChar is neither a variable nor a character  type error(character, OutChar)

Portability

ISO predicate. The type error(character,...) is a GNU Prolog behavior, the ISO reference instead defines a representation error(character) in this case. This seems to be an error of the ISO reference since, for many other built-in predicates accepting a character (e.g. char code/2, put char/2), a

type error is raised.

The ISO reference implies that if a program calls current char conversion/2, then modifies the character mapping by calling char conversion/2, and backtracks into the call to current char conversion/2 then the changes are guaranteed not to affect that current char conversion/2 goal. This is not guaranteed by GNU Prolog.

8.14.13 current char conversion/2

Templates

     current char conversion(?character, ?character)

Description

current char conversion(InChar, OutChar) succeeds if the conversion of InChar is OutChar according

to the character-conversion mapping. In that case, InChar and OutChar are different. This predicate is re-executable on backtracking.

Errors

 InChar is neither a variable nor a character  type error(character, InChar)
 OutChar is neither a variable nor a character  type error(character, OutChar)

Portability

ISO predicate. Same remark as for char conversion/2 (section 8.14.12, page 116).

8.15 Input/output from/to constant terms

These built-in predicates enable a Prolog term to be input from or output to a Prolog constant term (atom, character list or character code list). All these predicates can be defined using constant term streams (section 8.11, page 94). They are however simpler to use.

8.15.1 read term from atom/3, read from atom/2, read token from atom/2

Templates

     read term from atom(+atom ?term, +read option list)
     read from atom(+atom, ?term)
     read token from atom(+atom, ?nonvar)

Description

Like read term/3, read/2 (section 8.14.1, page 104) and read token/2 (section 8.14.3, page 106) except that characters are not read from a text-stream but from Atom; the atom given as first argument.

Errors

 Atom is a variable  instantiation error

Atom is neither a variable nor an atom type error(atom, Atom)

see associated predicate errors (section 8.14.1, page 104) and (section 8.14.3, page 106)

Portability

GNU Prolog predicates.

8.15.2 read term from chars/3, read from chars/2, read token from chars/2

Templates

     read term from chars(+character list ?term, +read option list)
     read from chars(+character list, ?term)
     read token from chars(+character list, ?nonvar)

Description

Like read term/3, read/2 (section 8.14.1, page 104) and read token/2 (section 8.14.3, page 106) except that characters are not read from a text-stream but from Chars; the character list given as first argument.

Chars is a partial list or a list with an element E which is a variable

                                             instantiation error

Chars is neither a partial list nor a list type error(list, Chars) an element E of the Chars list is neither a variable nor a character

                                             type error(character, E)

see associated predicate errors (section 8.14.1, page 104) and (section 8.14.3, page 106)

Portability

GNU Prolog predicates.

8.15.3 read term from codes/3, read from codes/2, read token from codes/2

Templates

     read term from codes(+character code list ?term, +read option list)
     read from codes(+character code list, ?term)
     read token from codes(+character code list, ?nonvar)

Description

Like read term/3, read/2 (section 8.14.1, page 104) and read token/2 (section 8.14.3, page 106) except that characters are not read from a text-stream but from Codes; the character code list given as first argument.

Errors

Codes is a partial list or a list with an element E which is a variable

                                             instantiation error

Codes is neither a partial list nor a list type error(list, Codes) an element E of the Codes list is neither a variable nor an integer

                                             type error(integer, E)

an element E of the Codes list is an integer but not a character code

                                             representation error(character code, E)

see associated predicate errors (section 8.14.1, page 104) and (section 8.14.3, page 106)

Portability

GNU Prolog predicates.

8.15.4 write term to atom/3, write to atom/2, writeq to atom/2,
        write canonical to atom/2, display to atom/2, print to atom/2,
        format to atom/3

Templates

     write term to atom(?atom, ?term, +write option list)
     write to atom(?atom, ?term)
     writeq to atom(?atom, ?term)
     write canonical to atom(?atom, ?term)

     display to atom(?atom, ?term)
     print to atom(?atom, ?term)
     format to atom(?atom, +character code list or atom, +list)

Description

Similar to write term/3, write/2, writeq/2, write canonical/2, display/2, print/2 (section 8.14.6,

page 108) and format/3 (section 8.14.7, page 110) except that characters are not written onto a textstream but are collected as an atom which is then unified with the first argument Atom.

Errors

Atom is neither a variable nor an atom type error(atom, Atom)

see associated predicate errors (section 8.14.6, page 108) and (section 8.14.7, page 110)

Portability

GNU Prolog predicates.

8.15.5 write term to chars/3, write to chars/2, writeq to chars/2,
        write canonical to chars/2, display to chars/2, print to chars/2,
        format to chars/3

Templates

     write term to chars(?character list, ?term, +write option list)
     write to chars(?character list, ?term)
     writeq to chars(?character list, ?term)
     write canonical to chars(?character list, ?term)
     display to chars(?character list, ?term)
     print to chars(?character list, ?term)
     format to chars(?character list, +character code list or atom, +list)

Description

Similar to write term/3, write/2, writeq/2, write canonical/2, display/2, print/2 (section 8.14.6,

page 108) and format/3 (section 8.14.7, page 110) except that characters are not written onto a textstream but are collected as a character list which is then unified with the first argument Chars.

Errors

Chars is neither a partial list nor a list type error(list, Chars) An element E of the list Chars is neither a variable nor a one-char atom

                                             type error(character, E)

see associated predicate errors (section 8.14.6, page 108) and (section 8.14.7, page 110)

Portability

8.15.6 write term to codes/3, write to codes/2, writeq to codes/2,
        write canonical to codes/2, display to codes/2, print to codes/2,
        format to codes/3

Templates

     write term to codes(?character code list, ?term, +write option list)
     write to codes(?character code list, ?term)
     writeq to codes(?character code list, ?term)
     write canonical to codes(?character code list, ?term)
     display to codes(?character code list, ?term)
     print to codes(?character code list, ?term)
     format to codes(?character code list, +character code list or atom, +list)

Description

Similar to write term/3, write/2, writeq/2, write canonical/2, display/2, print/2 (section 8.14.6,

page 108) and format/3 (section 8.14.7, page 110) except that characters are not written onto a textstream but are collected as a character code list which is then unified with the first argument Codes.

Errors

Codes is neither a partial list nor a list type error(list, Codes) An element E of the list Codes is neither a variable nor an integer

                                             type error(integer, E)

An element E of the list Codes is an integer but not a character code

                                             representation error(character code)

see associated predicate errors (section 8.14.6, page 108) and (section 8.14.7, page 110)

Portability

GNU Prolog predicates.

8.16 DEC-10 compatibility input/output

8.16.1 Introduction

The DEC-10 Prolog I/O predicates manipulate streams implicitly since they only refer to current input/output streams (section 8.10.1, page 78). The current input and output streams are initially set to user input and user output respectively. The predicate see/1 (resp. tell/1, append/1) can be used for setting the current input (resp. output) stream to newly opened streams for particular files. The predicate seen/0 (resp. told/0) close the current input (resp. output) stream, and resets it to the standard input (resp. output). The predicate seeing/1 (resp. telling/1) is used for retrieving the file name associated with the current input (resp. output) stream. The file name user stands for the standard input or output, depending on context (user input and user output can also be used). The DEC-10 Prolog I/O predicates are only provided for compatibility, they are now obsolete and their use is discouraged. The predicates for explicit stream manipulation should be used instead (section 8.10, page 78).

8.16.2 see/1, tell/1, append/1

Templates

     see(+source sink)
     see(+stream)
     tell(+source sink)
     tell(+stream)
     append(+source sink)
     append(+stream)

Description

see(FileName) sets the current input stream to FileName. If there is a stream opened by see/1 associated with the same FileName already, then it becomes the current input stream. Otherwise, FileName is opened for reading and becomes the current input stream.

tell(FileName) sets the current output stream to FileName. If there is a stream opened by tell/1 associated with the same FileName already, then it becomes the current output stream. Otherwise, FileName is opened for writing and becomes the current output stream.

append(FileName) like tell/1 but FileName is opened for writing + append.

A stream-term (obtained with any other built-in predicate) can also be provided as FileName to these predicates.

Errors

See errors associated with open/4 (section 8.10.6, page 81).

Portability

GNU Prolog predicates. Deprecated.

8.16.3 seeing/1, telling/1

Templates

     seeing(?source sink)
     telling(?source sink)

Description

seeing(FileName) succeeds if FileName unifies with the name of the current input file, if it was opened by see/1; else with the current input stream-term, if this is not user input, otherwise with user.

telling(FileName) succeeds if FileName unifies with the name of the current output file, if it was opened by tell/1 or append/1; else with the current output stream-term, if this is not user output, otherwise with user.

Errors

None.

Portability GNU Prolog predicates. Deprecated.

8.16.4 seen/0, told/0

Templates

     seen
     told

Description seen closes the current input, and resets it to user input. told closes the current output, and resets it to user output. Errors None. Portability GNU Prolog predicates. Deprecated.

8.16.5 get0/1, get/1, skip/1

Templates

     get0(?in character code)
     get(?in character code)
     skip(+character code)

Description get0(Code) succeeds if Code unifies with the next character code read from the current input stream. Thus it is equivalent to get code(Code) (section 8.12.1, page 97). get(Code) succeeds if Code unifies with the next character code read from the current input stream that is not a layout character. skip(Code) skips just past the next character code Code from the current input stream. Errors See errors for get code/2 (section 8.12.1, page 97). Portability GNU Prolog predicates. Deprecated.

8.16.6 put/1, tab/1

Templates

     put(+character code)
     tab(+evaluable)

Description

put(Code) writes the character whose code is Code onto the current output stream. It is equivalent to put code(Code) (section 8.12.5, page 100).

tab(N) writes N spaces onto the current output stream. N may be an arithmetic expression.

Errors

See errors for put code/2 (section 8.12.5, page 100) and for arithmetic expressions (section 8.6.1, page 66).

Portability

GNU Prolog predicates. Deprecated.

8.17 Term expansion

8.17.1 Definite clause grammars

Definite clause grammars are a useful notation to express grammar rules. However the ISO reference does not include them, so they should be considered as a system dependent feature. Definite clause grammars are an extension of context-free grammars. A grammar rule is of the form:

head --> body.

--> is a predefined infix operator (section 8.14.10, page 113).

Here are some features of definite clause grammars:

• a non-terminal symbol may be any callable term.

• a terminal symbol may be any Prolog term and is written as a list. The empty list represents an empty sequence of terminals.

• a sequence is expressed using the Prolog conjunction operator ((’,’)/2).

• the head of a grammar rule consists of a non-terminal optionally followed by a sequence of terminals (i.e. a Prolog list).

• the body of a grammar rule consists of a sequence of non-terminals, terminals, predicate call, disjunction (using ;/2), if-then (using (->)/2) or cut (using !).

• a predicate call must be enclosed in curly brackets (using {}/1). This makes it possible to express an extra condition.

A grammar rule is nothing but a “syntactic sugar” for a Prolog clause. Each grammar rule accepts as input a list of terminals (tokens), parses a prefix of this list and gives as output the rest of this list (possibly enlarged). This rest is generally parsed later. So, each a grammar rule is translated into a Prolog clause that explicitly the manages the list. Two arguments are then added: the input list (Start) and the output list (End). For instance:

p --> q.

is translated into:

p(Start, End) :- q(Start, End).

Extra arguments can be provided and the body of the rule can contain several non-terminals. Example:

p(X, Y) -->
q(X),
r(X, Y),
s(Y).

is translated into:

p(X, Y, Start, End) :-
q(X, Start, A),
r(X, Y, A, B),
s(Y, B, End).

Terminals are translated using unification:

assign(X,Y) --> left(X), [:=], right(Y), [;].

is translated into:

assign(X,Y,Start,End) :-
left(X, Start, A),
A=[:=|B],
right(Y, B, C),
C=[;|End].

Terminals appearing on the left-hand side of a rule are connected to the output argument of the head.

It is possible to include a call to a prolog predicate enclosing it in curly brackets (to distinguish them from non-terminals):

assign(X,Y) --> left(X), [:=], right(Y0), {Y is Y0 }, [;].

is translated into:

assign(X,Y,Start,End) :-
left(X, Start, A),
A=[:=|B],
right(Y0, B, C),
Y is Y0,
C=[;|End].

Cut, disjunction and if-then(-else) are translated literally (and do not need to be enclosed in curly brackets).

8.17.2 expand term/2, term expansion/2

Templates

expand term(?term, ?term)
term expansion(?term, ?term)

Description expand term(Term1, Term2) succeeds if Term2 is a transformation of Term1. The transformation steps are as follows:

• if Term1 is a variable, it is unified with Term2

• if term expansion(Term1, Term2) succeeds Term2 is assumed to be the transformation of Term1.

• if Term1 is a DCG then Term2 is its translation (section 8.17.1, page 123).

• otherwise Term2 is unified with Term1. term expansion(Term1, Term2) is a hook predicate allowing the user to define a specific transformation. The GNU Prolog compiler (section 4.4, page 21) automatically calls expand term/2 on each Term1 read in. However, in the current release, only DCG transformation are done by the compiler (i.e.

term expansion/2 cannot be used). To use term expansion/2, it is necessary to call expand term/2

explicitly. Errors None. Portability GNU Prolog predicate.

8.17.3 phrase/3, phrase/2

Templates

     phrase(?term, ?list, ?list)
     phrase(?term, ?list)

Description phrase(Phrase, List, Remainder) succeeds if the list List is in the language defined by the grammar rule body Phrase. Remainder is what remains of the list after a phrase has been found.

phrase(Phrase, List) is equivalent to phrase(Phrase, List, []).

Errors

 Phrase is a variable  instantiation error

Phrase is neither a variable nor a callable term type error(callable, Phrase) List is neither a list nor a partial list type error(list, List)

 Remainder is neither a list nor a partial list  type error(list, Remainder)

Portability GNU Prolog predicates.

8.18 Logic, control and exceptions

8.18.1 abort/0, stop/0, top level/0, break/0, halt/1, halt/0

Templates

     abort
     stop
     top level
     break
     halt(+integer)
     halt

Description

abort aborts the current execution. If this execution was initiated under a top-level the control is given back to the top-level and the message {execution aborted} is displayed. Otherwise, e.g. execution started by a initialization/1 directive (section 7.1.14, page 50), abort/0 is equivalent to halt(1) (see below).

stop stops the current execution. If this execution was initiated under a top-level the control is given back to the top-level. Otherwise, stop/0 is equivalent to halt(0) (see below).

top level starts a new recursive top-level (including the banner display). To end this new top-level simply type the end-of-file key sequence (Ctl-D) or its term representation: end of file.

break invokes a recursive top-level (no banner is displayed). To end this new level simply type the end-of-file key sequence (Ctl-D) or its term representation: end of file.

halt(Status) causes the GNU Prolog process to immediately exit back to the shell with the return code

Status.

halt is equivalent to halt(0).

Errors

 Status is a variable  instantiation error
 Status is neither a variable nor an integer  type error(integer, Status)

Portability

halt/1 and halt/0 are ISO predicates. abort/0, stop/0, top level/0 and break/0 are GNU Prolog

predicates.

8.18.2 false/0, once/1, (\+)/1 - not provable, call/2-11, call with args/1-11, call det/2,
        forall/2

Templates

     false
     once(+callable term)
     \+(+callable term)
     call(+callable term, +term,..., +term)

     call with args(+atom, +term,..., +term)
     call det(+callable term, ?boolean)
     forall(+callable term, +callable term)

Description

false always fails and enforces backtracking. It is equivalent to the fail/0 control construct (section 7.2.1, page 51).

once(Goal) succeeds if call(Goal) succeeds. However once/1 is not re-executable on backtracking since all alternatives of Goal are cut. once(Goal) is equivalent to call(Goal), !.

\+ Goal succeeds if call(Goal) fails and fails otherwise. This built-in predicate gives negation by failure.

call(Closure, Arg1,..., ArgN) calls the goal call(Goal) where Goal is constructed by appending

Arg1,..., ArgN (1 ≤ N ≤ 10) additional arguments to the arguments (if any) of Closure.

call with args(Functor, Arg1,..., ArgN) calls the goal whose functor is Functor and whose argu-

ments are Arg1,…, ArgN (0 ≤ N ≤ 10).

call det(Goal, Deterministic) succeeds if call(Goal) succeeds and unifies Deterministic with

true if Goal has not created any choice-points, with false otherwise.

forall(Condition, Action) succeeds if for all alternative bindings of Condition, Action can be proven.

It is equivalent to \+ (Condition, \+ Action).

\+ is a predefined prefix operator (section 8.14.10, page 113).

Errors

 Goal (or Condition or Action) is a variable  instantiation error
 Goal (or Condition or Action) is neither a

variable nor a callable term

                                             type error(callable, Goal)

The predicate indicator Pred of Goal does not correspond to an existing procedure and the value of the unknown Prolog flag is error (section 8.22.1, page 148)

                                             existence error(procedure, Pred)

 Functor is a variable  instantiation error
 Functor is neither a variable nor an atom  type error(atom, Functor)

Deterministic is neither a variable nor a boolean

                                             type error(boolean, Deterministic)

for call/2-11 the resulting arity of Goal (arity of Closure + N) is an integer > max arity flag (section 8.22.1, page 148)

                                             representation error(max arity)

Portability

false/0, call/2-8, once/1 and (\+)/1 are ISO predicates. call/9-11, call with args/1-11, call det/2

and forall/2 are GNU Prolog predicates.

8.18.3 repeat/0

     repeat

Description repeat generates an infinite sequence of backtracking choices. The purpose is to repeatedly perform some action on elements which are somehow generated, e.g. by reading them from a stream, until some test becomes true. Repeat loops cannot contribute to the logic of the program. They are only meaningful if the action involves side-effects. The only reason for using repeat loops instead of a more natural tailrecursive formulation is efficiency: when the test fails back, the Prolog engine immediately reclaims any working storage consumed since the call to repeat/0. Errors None. Portability ISO predicate.

8.18.4 between/3, for/3

Templates

     between(+integer, +integer, ?integer)
     for(?integer, +integer, +integer)

Description between(Lower, Upper, Counter) generates an sequence of backtracking choices instantiating Counter to the values Lower, Lower+1,…, Upper.

for(Counter, Lower, Upper) is equivalent to between(Lower, Upper, Counter). This predicate is

deprecated and new code should use between/3. Errors

 Counter is neither a variable nor an integer  type error(integer, Counter)
 Lower is a variable  instantiation error

Lower is neither a variable nor an integer type error(integer, Lower)

 Upper is a variable  instantiation error

Upper is neither a variable nor an integer type error(integer, Upper) Portability GNU Prolog predicate.

8.19 Atomic term processing

These built-in predicates enable atomic terms to be processed as a sequence of characters and character codes. Facilities exist to split and join atoms, to convert a single character to and from the corresponding character code, and to convert a number to and from a list of characters and character codes.

8.19.1 atom length/2

Templates

     atom length(+atom, ?integer)

Description atom length(Atom, Length) succeeds if Length unifies with the number of characters of the name of

Atom.

Errors

 Atom is a variable  instantiation error

Atom is neither a variable nor an atom type error(atom, Atom)

 Length is neither a variable nor an integer  type error(integer, Length)
 Length is an integer < 0  domain error(not less than zero, Length)

Portability ISO predicate.

8.19.2 atom concat/3

Templates

     atom concat(+atom, +atom, ?atom)
     atom concat(?atom, ?atom, +atom)

Description atom concat(Atom1, Atom2, Atom12) succeeds if the name of Atom12 is the concatenation of the name of Atom1 with the name of Atom1. This predicate is re-executable on backtracking (e.g. if Atom12 is instantiated and both Atom1 and Atom2 are variables). Errors

 Atom1 and Atom12 are variables  instantiation error
 Atom2 and Atom12 are variables  instantiation error

Atom1 is neither a variable nor an atom type error(atom, Atom1) Atom2 is neither a variable nor an atom type error(atom, Atom2)

 Atom12 is neither a variable nor an atom  type error(atom, Atom12)

Portability ISO predicate.

8.19.3 sub atom/5

Templates

     sub atom(+atom, ?integer, ?integer, ?integer, ?atom)

Description

sub atom(Atom, Before, Length, After, SubAtom) succeeds if atom Atom can be split into three

atoms, AtomL, SubAtom and AtomR such that Before is the number of characters of the name of AtomL, Length is the number of characters of the name of SubAtom and After is the number of characters of the name of AtomR. This predicate is re-executable on backtracking.

Errors

 Atom is a variable  instantiation error

Atom is neither a variable nor an atom type error(atom, Atom)

 SubAtom is neither a variable nor an atom  type error(atom, SubAtom)
 Before is neither a variable nor an integer  type error(integer, Before)
 Length is neither a variable nor an integer  type error(integer, Length)

After is neither a variable nor an integer type error(integer, After)

 Before is an integer < 0  domain error(not less than zero, Before)
 Length is an integer < 0  domain error(not less than zero, Length)
 After is an integer < 0  domain error(not less than zero, After)

Portability

ISO predicate.

8.19.4 char code/2

Templates

     char code(+character, ?character code)
     char code(-character, +character code)

Description

char code(Char, Code) succeeds if the character code for the one-char atom Char is Code.

Errors

 Char and Code are variables  instantiation error

Char is neither a variable nor a one-char atom type error(character, Char) Code is neither a variable nor an integer type error(integer, Code)

 Code is an integer but not a character code  representation error(character code)

Portability

ISO predicate.

8.19.5 lower upper/2

Templates

     lower upper(+character, ?character)
     lower upper(-character, +character)

lower upper(Char1, Char2) succeeds if Char1 and Char2 are one-char atoms and if Char2 is the upper conversion of Char1. If Char1 (resp. Char2) is a character that is not a lower (resp. upper) letter then

Char2 is equal to Char1.

Errors

 Char1 and Char2 are variables  instantiation error

Char1 is neither a variable nor a one-char atom type error(character, Char1) Char2 is neither a variable nor a one-char atom type error(character, Char2)

Portability

GNU Prolog predicate.

8.19.6 atom chars/2, atom codes/2

Templates

     atom chars(+atom, ?character list)
     atom chars(-atom, +character list)
     atom codes(+atom, ?character code list)
     atom codes(-atom, +character code list)

Description

atom chars(Atom, Chars) succeeds if Chars is the list of one-char atoms whose names are the successive characters of the name of Atom.

atom codes(Atom, Codes) is similar to atom chars/2 but deals with a list of character codes.

Errors

Atom is a variable and Chars (or Codes) is a partial list or a list with an element which is a variable

                                             instantiation error

Atom is neither a variable nor an atom type error(atom, Atom) Chars is neither a list nor a partial list type error(list, Chars) Codes is neither a list nor a partial list type error(list, Codes) An element E of the list Chars is neither a variable nor a one-char atom

                                             type error(character, E)

An element E of the list Codes is neither a variable nor an integer

                                             type error(integer, E)

An element E of the list Codes is an integer but not a character code

                                             representation error(character code)

Portability

ISO predicates. The ISO reference only causes a type error(list, Chars) if Atom is a variable and Chars is neither a list nor a partial list. GNU Prolog always checks if Chars is a list. Similarly for Codes. The type error(integer, E) when an element E of the Codes is not an integer is a GNU Prolog extension. This seems to be an omission in the ISO reference since this error is detected for many other built-in predicates accepting a character code (e.g. char code/2, put code/2).

8.19.7 number atom/2, number chars/2, number codes/2

Templates

    number atom(+number, ?atom)
    number atom(-number, +atom)
    number chars(+number, ?character list)
    number chars(-number, +character list)
    number codes(+number, ?character code list)
    number codes(-number, +character code list)

Description

number atom(Number, Atom) succeeds if Atom is an atom whose name corresponds to the characters of

Number.

number chars(Number, Chars) is similar to number atom/2 but deals with a list of characters.

number codes(Number, Codes) is similar to number atom/2 but deals with a list of character codes.

Errors

Number and Atom are variables  instantiation error

Number is a variable and Chars (or Codes) is a partial list or a list with an element which is a variable

                                           instantiation error
Number is neither a variable nor an number  type error(number, Number)

Atom is neither a variable nor an atom type error(atom, Atom) Chars is neither a list nor a partial list type error(list, Chars) Codes is neither a list nor a partial list type error(list, Codes) An element E of the list Chars is neither a variable nor a one-char atom

                                           type error(character, E)

An element E of the list Codes is neither a variable nor an integer

                                           type error(integer, E)

An element E of the list Codes is an integer but not a character code

                                           representation error(character code)
Number is a variable, Atom (or Chars or Codes)

cannot be parsed as a number and the value of

the syntax error Prolog flag is error

(section 8.22.1, page 148)

                                           syntax error(atom explaining the error)

Portability

number atom/2 is a GNU Prolog predicate. number chars/2 and number codes/2 are ISO predicates.

GNU Prolog only raises an error about an element E of the Chars (or Codes) list when Number is a variable while the ISO reference always check this. This seems an error since the list itself is only checked if Number is a variable.

The type error(integer, E) when an element E of the Codes is not an integer is a GNU Prolog extension. This seems to be an omission in the ISO reference since this error is detected for many other built-in predicates accepting a character code (e.g. char code/2, put code/2).

8.19.8 name/2

Templates

     name(+atomic, ?character code list)
     name(-atomic, +character code list)

Description

name(Constant, Codes) succeeds if Codes is a list whose elements are the character codes corresponding to the successive characters of Constant (a number or an atom). However, there atoms are for which name(Constant, Codes) is true, but which will not be constructed if name/2 is called with Constant uninstantiated, e.g. the atom '1024'. For this reason the use of name/2 is discouraged and should be limited to compatibility purposes. It is preferable to use atom codes/2 (section 8.19.6, page 131) or number chars/2 (section 8.19.7, page 132).

Errors

Constant is a variable and Codes is a partial list or a list with an element which is a variable

                                             instantiation error

Constant is neither a variable nor an atomic term

                                             type error(atomic, Constant)

Constant is a variable and Codes is neither a list nor a partial list

                                             type error(list, Codes)

Constant is a variable and an element E of the list Codes is neither a variable nor an integer

                                             type error(integer, E)

Constant is a variable and an element E of the list Codes is an integer but not a character code

                                             representation error(character code)

Portability

GNU Prolog predicate.

8.19.9 new atom/2, new atom/1

Templates

     new atom(+atom, -atom)
     new atom(-atom)

Description

new atom(Prefix, Atom) unifies Atom with a new atom whose name begins with the characters of the name of Prefix. This predicate is then a symbol generator. It is guaranteed that Atom does not exist before the invocation of new atom/3. The characters appended to Prefix to form Atom are in: A-Z (capital letter), a-z (small letter) and 0-9 (digit).

new atom/1 is similar to new atom(term , Atom), i.e. the generated atom begins with term .

Errors

 Prefix is a variable  instantiation error
 Prefix is neither a variable nor an atom  type error(atom, Prefix)
 Atom is not a variable  uninstantiation error(Atom)

Portability GNU Prolog predicate.

8.19.10 current atom/1

Templates

     current atom(?atom)

Description current atom(Atom) succeeds if there exists an atom that unifies with Atom. All atoms are found except those beginning with a '$' (system atoms). This predicate is re-executable on backtracking. Errors Atom is neither a variable nor an atom type error(atom, Atom) Portability GNU Prolog predicate.

8.19.11 atom property/2

Templates

     atom property(?atom, ?atom property)

Description

atom property(Atom, Property) succeeds if current atom(Atom) succeeds (section 8.19.10, page 134)

and if Property unifies with one of the properties of the atom. This predicate is re-executable on backtracking. Atom properties:

• length(Length): Length is the length of the name of the atom.

• hash(Hash): Hash is the hash code of the atom, see also term hash/2 (section 8.4.8, page 62).

• prefix op: if there is a prefix operator currently defined with this name.

• infix op: if there is an infix operator currently defined with this name.

• postfix op: if there is a postfix operator currently defined with this name.

• needs quotes: if the atom must be quoted to be read later.

• needs scan: if the atom must be scanned when output to be read later (e.g. contains special

characters that must be output with a \ escape sequence). Errors Atom is neither a variable nor an atom type error(atom, Atom) Property is neither a variable nor a n atom property term

                                             domain error(atom property, Property)
 Property = length(E) or hash(E) and E is

neither a variable nor an integer

                                             type error(integer, E)

Portability GNU Prolog predicate.

8.20 List processing

These predicates manipulate lists. They are bootstrapped predicates (i.e. written in Prolog) and no error cases are tested (for the moment). However, since they are written in Prolog using other built-in predicates, some errors can occur due to those built-in predicates.

8.20.1 append/3

Templates

     append(?list, ?list, ?list)

Description append(List1, List2, List12) succeeds if the concatenation of the list List1 and the list List2 is the list List12. This predicate is re-executable on backtracking (e.g. if List12 is instantiated and both List1 and List2 are variable). Errors None. Portability GNU Prolog predicate.

8.20.2 member/2, memberchk/2

Templates

     member(?term, ?list)
     memberchk(?term, ?list)

Description member(Element, List) succeeds if Element belongs to the List. This predicate is re-executable on backtracking and can be thus used to enumerate the elements of List. memberchk/2 is similar to member/2 but only succeeds once. Errors None. Portability GNU Prolog predicate.

8.20.3 reverse/2

Templates

     reverse(?list, ?list)

Description reverse(List1, List2) succeeds if List2 unifies with the list List1 in reverse order. Errors None. Portability GNU Prolog predicate.

8.20.4 delete/3, select/3

Templates

     delete(?list, ?term, ?list)
     select(?term, ?list, ?list)

Description

delete(List1, Element, List2) removes all occurrences of Element in List1 to provide List2. A

strict term equality is required, cf. (==)/2 (section 8.3.2, page 57).

select(Element, List1, List2) removes one occurrence of Element in List1 to provide List2. This

predicate is re-executable on backtracking. Errors None. Portability GNU Prolog predicate.

8.20.5 subtract/3

Templates

     subtract(+list, +list, ?list)

Description

subtract(List1, List2, List3) removes all elements in List2 from List1 to provide List3. Member-

ship is tested using memberchk/2 (section 8.20.2, page 135). The predicate runs in O(|List2| × |List1|). Errors None. Portability GNU Prolog predicate.

8.20.6 permutation/2

Templates

     permutation(?list, ?list)

Description permutation(List1, List2) succeeds if List2 is a permutation of the elements of List1. This predicate is re-executable on backtracking. Errors None. Portability GNU Prolog predicate.

8.20.7 prefix/2, suffix/2

Templates

     prefix(?list, ?list)
     suffix(?list, ?list)

Description prefix(Prefix, List) succeeds if Prefix is a prefix of List. This predicate is re-executable on backtracking. suffix(Suffix, List) succeeds if Suffix is a suffix of List. This predicate is re-executable on backtracking. Errors None. Portability GNU Prolog predicate.

8.20.8 sublist/2

Templates

     sublist(?list, ?list)

Description sublist(List1, List2) succeeds if all elements of List1 appear in List2 in the same order. This predicate is re-executable on backtracking. Errors None. Portability GNU Prolog predicate.

8.20.9 last/2

Templates

     last(?list, ?term)

Description

last(List, Element) succeeds if Element is the last element of List.

Errors None. Portability GNU Prolog predicate.

8.20.10 flatten/2

Templates

     flatten(?term, ?list)

Description flat(List1, List2) succeeds if List2 is the flatten version of List1. Errors None. Portability GNU Prolog predicate.

8.20.11 length/2

Templates

     length(?list, ?integer)

Description

length(List, Length) succeeds if Length is the length of List.

Errors

 Length is an integer < 0  domain error(not less than zero, Length)

GNU Prolog predicate.

8.20.12 nth/3 Templates

     nth(?integer, ?list, ?term)

Description

nth(N, List, Element) succeeds if the Nth argument of List is Element.

Errors None. Portability GNU Prolog predicate.

8.20.13 max list/2, min list/2, sum list/2

Templates

     min list(+list, ?number)
     max list(+list, ?number)
     sum list(+list, ?number)

Description min list(List, Min) succeeds if Min is the smallest number in List. max list(List, Max) succeeds if Max is the largest number in List. sum list(List, Sum) succeeds if Sum is the sum of all the elements in List. List must be a list of arithmetic evaluable terms (section 8.6.1, page 66). Errors None. Portability GNU Prolog predicate.

8.20.14 maplist/2-8

Templates

     maplist(+callable term, +list, ..., +list)

Description maplist(Goal, List) succeeds if Goal can succesfully be applied on all elements of List. maplist(Goal, List1, List2) succeeds if Goal can succesfully be applied to all pairs of elements of

List1 and List2.

maplist(Goal, List1, List2, List3) succeeds if Goal can succesfully be applied to all triples of elements of List1..List3. maplist(Goal, List1, List2, ..., ListN) succeeds if Goal can succesfully be applied to all N-uples (N ≤ 8) of elements of List1..ListN. Errors an error occurs executing a directive see call/1 errors (section 7.2.3, page 52) Portability GNU Prolog predicate.

8.20.15 sort/2, msort/2, keysort/2 sort/1, msort/1, keysort/1

Templates

     sort(+list, ?list)
     msort(+list, ?list)
     keysort(+list, ?list)
     sort(+list)
     msort(+list)
     keysort(+list)

Description sort(List1, List2) succeeds if List2 is the sorted list corresponding to List1 where duplicate elements are merged. msort/2 is similar to sort/2 except that duplicate elements are not merged. keysort(List1, List2) succeeds if List2 is the sorted list of List1 according to the keys. The list List1 consists of pairs (items of the form Key-Value). These items are sorted according to the value of Key yielding the List2. Duplicate keys are not merged. This predicate is stable, i.e. if K-A occurs before K-B in the input, then K-A will occur before K-B in the output.

sort/1, msort/1 and keysort/1 are similar to sort/2, msort/2 and keysort/2 but achieve a sort in-

place destructing the original List1 (this in-place assignment is not undone at backtracking). The sorted list occupies the same memory space as the original list (saving thus memory consumption).

The time complexity of these sorts is O(N log N), N being the length of the list to sort.

These predicates refer to the standard ordering of terms (section 8.3.1, page 57).

Errors

 List1 is a partial list  instantiation error

List1 is neither a partial list nor a list type error(list, List1) List2 is neither a partial list nor a list type error(list, List2)

 for keysort/2: an element of List1 is a variable instantiation error

for keysort/2: an element E of List1 is neither a variable nor a pair

                                             type error(pair, E)

for keysort/2: an element E of List2 is neither a variable nor a pair

                                             type error(pair, E)

Portability

sort/2 and keysort/2 are ISO predicates.

sort/1, keysort/1 and msort/1-2 are GNU Prolog predicates.

8.21 Global variables

8.21.1 Introduction

GNU Prolog provides a simple and powerful way to assign and read global variables. A global variable is associated with each atom, its initial value is the integer 0. A global variable can store 3 kinds of objects:

• a copy of a term (the assignment can be made backtrackable or not).

• a link to a term (the assignment is always backtrackable).

• an array of objects (recursively).

The space necessary for copies and arrays is dynamically allocated and recovered as soon as possible. For instance, when an atom is associated with a global variable whose current value is an array, the space for this array is recovered (unless the assignment is to be undone when backtracking occurs).

When a link to a term is associated with a global variable, the reference to this term is stored and thus the original term is returned when the content of the variable is read.

Global variable naming convention: a global variable is referenced by an atom.

If the variable contains an array, an index (ranging from 0) can be provided using a compound term whose principal functor is the corresponding atom and the argument is the index. In case of a multi-dimensional array, each index is given as the arguments of the compound term.

If the variable contains a term (link or copy), it is possible to only reference a sub-term by giving its argument number (also called argument selector). Such a sub-term is specified using a compound term whose principal functor is -/2 and whose first argument is a global variable name and the second argument is the argument number (from 1). This can be applied recursively to specify a sub-term of any depth. In case of a list, a argument number I represents the Ith element of the list. In the rest of this section we use the operator notation since - is a predefined infix operator (section 8.14.10, page 113).

In the following, GVarName represents a reference to a global variable and its syntax is as follows:

GVarName ::= atom  whole content of a variable

                    atom(Integer,. . . ,Integer) element of an array

GVarName-Integer  sub-term selection
Integer ::= integer  immediate value
GVarName  indirect value

When a GVarName is used as an index or an argument number (i.e. indirection), the value of this variable must be an integer.

Here are some examples of the naming convention:

a  the content of variable associated with a (any kind)
t(1) the 2nd element of the array associated with t
t(k) if the value associated with k is I, the Ith element of the array associated with t
a-1-2 if the value associated with a is f(g(a,b,c),2), the sub-term b

Here are the errors associated with global variable names and common to all predicates.

 GVarName is a variable  instantiation error
 GVarName is neither a variable nor a callable term type error(callable, GVarName)

GVarName contains an invalid argument number (or GVarName is an array)

                                             domain error(g argument selector,
                                             GVarName)

GVarName contains an invalid index (or GVarName is not an array)

                                             domain error(g array index, GVarName)

GVarName is used as an indirect index or argument selector and is not an integer

                                             type error(integer, GVarName)

Arrays: the predicates g assign/2, g assignb/2 and g link/2 (section 8.21.2, page 143) can be used to create an array. They recognize some terms as values. For instance, a compound term with principal functor g array is used to define an array of fixed size. There are 3 forms for the term g array:

• g array(Size): if Size is an integer > 0 then defines an array of Size elements which are all initialized with the integer 0.

• g array(Size, Initial): as above but the elements are initialized with the term Initial instead of 0. Initial can contain other array definitions allowing thus for multi-dimensional arrays.

• g array(List): as above if List is a list of length Size except that the elements of the array are initialized according to the elements of List (which can contain other array definitions).

An array can be extended explicitly using a compound term with principal functor g array extend which accept the same 3 forms detailed above. In that case, the existing elements of the array are not initialized. If g array extend is used with an object which is not an array it is similar to g array.

Finally, an array can be automatically expanded when needed. The programmer does not need to explicitly control the expansion of an automatic array. An array is expanded as soon as an index is outside the current size of this array. Such an array is defined using a compound term with principal

functor g array auto:

• g array auto(Size): if Size is an integer > 0 then defines an automatic array whose initial size is Size. All elements are initialized with the integer 0. Elements created during implicit expansions will be initialized with 0.

• g array auto(Size, Initial): as above but the elements are initialized with the term Initial instead of 0. Initial can contain other array definitions allowing thus for multi-dimensional arrays. Elements created during implicit expansions will be initialized with Initial.

• g array auto(List): as above if List is a list of length Size except that the elements of the array are initialized according to the elements of List (which can contain other array definitions). Elements created during implicit expansions will be initialized with 0.

In any case, when an array is read, a term of the form g array([Elem0,..., ElemSize-1]) is returned.

Some examples using global variables are presented later (section 8.21.7, page 146).

8.21.2 g assign/2, g assignb/2, g link/2

Templates

     g assign(+callable term, ?term)
     g assignb(+callable term, ?term)
     g link(+callable term, ?term)

Description

g assign(GVarName, Value) assigns a copy of the term Value to GVarName. This assignment is not undone when backtracking occurs.

g assignb/2 is similar to g assign/2 but the assignment is undone at backtracking.

g link(GVarName, Value) makes a link between GVarName to the term Value. This allows the user to give a name to any Prolog term (in particular non-ground terms). Such an assignment is always undone when backtracking occurs (since the term may no longer exist). If Value is an atom or an integer, g link/2 and g assignb/2 have the same behavior. Since g link/2 only handles links to existing terms it does not require extra memory space and is not expensive in terms of execution time.

NB: argument selectors can only be used with g assign/2 (i.e. when using an argument selector inside an assignment, this one must not be backtrackable).

Errors

See common errors detailed in the introduction (section 8.21.1, page 141)

GVarName contains an argument selector and the assignment is backtrackable

                                             domain error(g argument selector,
                                             GVarName)

Portability

GNU Prolog predicates.

8.21.3 g read/2

     g read(+callable term, ?term)

Description

g read(GVarName, Value) unifies Value with the term assigned to GVarName.

Errors See common errors detailed in the introduction (section 8.21.1, page 141) Portability GNU Prolog predicate.

8.21.4 g array size/2

Templates

     g array size(+callable term, ?integer)

Description g array size(GVarName, Value) unifies Size with the dimension (an integer > 0) of the array assigned to GVarName. Fails if GVarName is not an array. Errors See common errors detailed in the introduction (section 8.21.1, page 141) Size is neither a variable nor an integer type error(integer, Size) Portability GNU Prolog predicate.

8.21.5 g inc/3, g inc/2, g inco/2, g inc/1, g dec/3, g dec/2, g deco/2, g dec/1

Templates

     g inc(+callable term, ?integer, ?integer)
     g inc(+callable term, ?integer)
     g inco(+callable term, ?integer)
     g inc(+callable term)
     g dec(+callable term, ?integer, ?integer)
     g dec(+callable term, ?integer)
     g deco(+callable term, ?integer)
     g dec(+callable term)

Description

g inc(GVarName, Old, New) unifies Old with the integer assigned to GVarName, increments GVarName

and then unifies New with the incremented value.

g inc(GVarName, New) is equivalent to g inc(GVarName,  , New).
g inco(GVarName, Old) is equivalent to g inc(GVarName, Old,  ).
g inc(GVarName) is equivalent to g inc(GVarName,  ,  ).

Predicates g dec are similar but decrement the content of GVarName instead. Errors See common errors detailed in the introduction (section 8.21.1, page 141) Old is neither a variable nor an integer type error(integer, Old) New is neither a variable nor an integer type error(integer, New)

 GVarName stores an array  type error(integer, g array)

GVarName stores a term T which is not an integer type error(integer, T) Portability GNU Prolog predicates.

8.21.6 g set bit/2, g reset bit/2, g test set bit/2, g test reset bit/2

Templates

     g set bit(+callable term, +integer)
     g reset bit(+callable term, +integer)
     g test set bit(+callable term, +integer)
     g test reset bit(+callable term, +integer)

Description g set bit(GVarName, Bit) sets to 1 the bit number specified by Bit of the integer assigned to GVarName to 1. Bit numbers range from 0 to the maximum number allowed for integers (this is architecture dependent). If Bit is greater than this limit, the modulo with this limit is taken. g reset bit(GVarName, Bit) is similar to g set bit/2 but sets the specified bit to 0. g test set bit/2 succeeds if the specified bit is set to 1. g test reset bit/2 succeeds if the specified bit is set to 0. Errors See common errors detailed in the introduction (section 8.21.1, page 141)

 Bit is a variable  instantiation error

Bit is neither a variable nor an integer type error(integer, Bit)

 Bit is an integer < 0  domain error(not less than zero, Bit)
 GVarName stores an array  type error(integer, g array)

GVarName stores a term T which is not an integer type error(integer, T) Portability GNU Prolog predicates.

8.21.7 Examples

Simulating g inc/3: this predicate behaves like: global variable:

     my_g_inc(Var, Old, New) :-
             g_read(Var, Old),
             N is Value + 1,
             g_assign(Var, X),
     New = N.

The query: my g inc(c, X, ) will succeed unifying X with 0, another call to my g inc(a, Y, ) will then unify Y with 1, and so on.

Difference between g assign/2 and g assignb/2: g assign/2 does not undo its assignment when backtracking occurs whereas g assignb/2 undoes it.

     test(Old) :-                       testb(Old) :-
             g assign(x,1),                     g assign(x,1),
             (   g read(x, Old),                (   g read(x, Old),
                 g assign(x, 2)                    g assignb(x, 2)
             ;   g read(x, Old),                ;   g read(x, Old),
                 g assign(x, 3)                     g assign(x, 3)
             ).                                 ).

The query test(Old) will succeed unifying Old with 1 and on backtracking with 2 (i.e. the assignment of the value 2 has not been undone). The query testb(Old) will succeed unifying Old with 1 and on backtracking with 1 (i.e. the assignment of the value 2 has been undone).

Difference between g assign/2 and g link/2: g assign/2 (and g assignb/2) creates a copy of the term whereas g link/2 does not. g link/2 can be used to avoid passing big data structures (e.g. dictionaries,…) as arguments to predicates.

     test(B) :-                         test(B) :-
             g assign(b, f(X)),                g link(b, f(X)),
             X = 12,                            X = 12,
             g read(b, B).                      g read(b, B).

The query test(B) will succeed unifying B with f( ) (g assign/2 assigns a copy of the value). The query test(B) will succeed unifying B with f(12) (g link/2 assigns a pointer to the term).

Simple array definition: here are some queries to show how arrays can be handled:

     | ?- g_assign(w, g_array(3)), g_read(w, X).
     X = g_array([0,0,0])
     | ?- g_assign(w(0), 16), g_assign(w(1), 32), g_assign(w(2), 64), g_read(w, X).
     X = g_array([16,32,64])

this is equivalent to:

     | ?- g_assign(k, g_array([16,32,64])), g_read(k, X).

X = g_array([16,32,64])
| ?- g_assign(k, g_array(3,null)), g_read(k, X), g_array_size(k, S).
S = 3
X = g_array([null,null,null])

2-D array definition:

| ?- g_assign(w, g_array(2, g_array(3))), g_read(w, X).
X = g_array([g_array([0,0,0]),g_array([0,0,0])])
| ?- (   for(I,0,1), for(J,0,2), K is I*3+J, g_assign(w(I,J), K),
         fail
     ;   g_read(w, X)
     ).
X = g_array([g_array([0,1,2]),g_array([3,4,5])])
| ?- g_read(w(1),X).
X = g_array([3,4,5])

Hybrid array:

| ?- g_assign(w,g_array([1,2,g_array([a,b,c]), g_array(2,z),5])), g_read(w, X).
X = g_array([1,2,g_array([a,b,c]), g_array([z,z]),5])
| ?- g_read(w(1), X), g_read(w(2,1), Y), g_read(w(3,1), Z).
X = 2
Y = b
Z = z
| ?- g_read(w(1,2),X).
uncaught exception: error(domain_error(g_array_index,w(1,2)),g_read/2)

Array extension:

| ?- g_assign(a, g_array([10,20,30])), g_read(a, X).
X = g_array([10,20,30])
| ?- g_assign(a, g_array_extend(5,null)), g_read(a, X).
X = g_array([10,20,30,null,null])
| ?- g_assign(a, g_array([10,20,30])), g_read(a, X).
X = g_array([10,20,30])
| ?- g_assign(a, g_array_extend([1,2,3,4,5,6])), g_read(a, X).
X = g_array([10,20,30,4,5,6])

Automatic array:

| ?- g_assign(t, g_array_auto(3)), g_assign(t(1), foo), g_read(t,X).
X = g_array([0,foo,0])
| ?- g_assign(t(5), bar), g_read(t,X).
X = g_array([0,foo,0,0,0,bar,0,0])
| ?- g_assign(t, g_array_auto(2, g_array(2))), g_assign(t(1,1), foo),
     g_read(t,X).
X = g_array([g_array([0,0]),g_array([0,foo])])
| ?- g_assign(t(3,0), bar), g_read(t,X).
X = g_array([g_array([0,0]),g_array([0,foo]),g_array([0,0]),g_array([bar,0])])
| ?- g_assign(t(3,4), bar), g_read(t,X).
uncaught exception: error(domain_error(g_array_index,t(3,4)),g_assign/2)
| ?- g_assign(t, g_array_auto(2, g_array_auto(2))), g_assign(t(1,1), foo),
     g_read(t,X).
X = g_array([g_array([0,0]),g_array([0,foo])])
| ?- g_assign(t(3,3), bar), g_read(t,X).
X = g_array([g_array([0,0]),g_array([0,foo]),g_array([0,0]),
    g_array([0,0,0,bar])])
| ?- g_assign(t, g_array_auto(2, g_array_auto(2, null))), g_read(t(2,3), U),
     g_read(t, X).
U = null
X = g_array([g_array([null,null]),g_array([null,null]),
             g_array([null,null,null,null]),g_array([null,null])])

8.22 Prolog state

8.22.1 set prolog flag/2

Templates

set prolog flag(+flag, +term)

Description

set prolog flag(Flag, Value) sets the value of the Prolog flag Flag to Value.

Prolog flags: a Prolog flag is an atom which is associated with a value that is either implementation defined or defined by the user. Each flag has a permitted range of values; any other value is a domain error. The following two tables present available flags, the possible values, a description and if they are ISO or an extension. The first table presents unchangeable flags while the second one the changeable flags. For flags whose default values is machine independent, this value is \underline{underlined}.

Unchangeable flags:

Flag  Values  Description  ISO

prolog name

an atom name of the Prolog system

N

prolog version

an atom version number of the Prolog system

N

prolog date

an atom date of the Prolog system

N

prolog copyright

an atom copyright message of the Prolog system

N

dialect

an atom

                                         fixed to gprolog

N

version                      an integer  Major ∗ 10000 + Minor ∗ 100 + P atch  N

version data

a structure

                                         gprolog(Major,Minor,Patch,Extra)

N

bounded

                           \underline{true} / false

are integers bounded ?

Y

max integer

an integer greatest integer

Y

min integer

an integer smallest integer

Y

integer rounding function

                           toward zero down

rnd(X) = integer part of X

rnd(X) = ⌊X⌋ (section 8.6.1, page 66)

Y

max arity

an integer maximum arity for compound terms (255)

Y

max atom

an integer maximum number of atoms

N

max unget

an integer maximum number of successive ungets

N

home

an atom GNU Prolog home directory

N

host os

an atom Operating System identifier

N

host vendor

an atom Operating System vendor

N

host cpu

an atom processor identifier

N

host

an atom a combination of the OS-vendor-cpu

N

arch

an atom a combination of the OS-cpu

N

address bits

an integer address size of the machine (32 or 64)

N

unix

                              on/off

is the architecture an Unix-like OS ?

N

compiled at

an atom compilation date using DATE and TIME C compiler macros

N

c cc

an atom C compiler used to compile GNU Prolog

                                         (gcc, cc, clang, cl,...)

N

c cc version data

a structure

                                         c cc(Major,Minor,Patch,Extra)

N

c cflags

an atom CFLAGS used to compile GNU Prolog

N

c ldflags

an atom LDFLAGS used to compile GNU Prolog

N

argv

a list of atoms list of command-line arguments

N

Changeable flags:

Flag  Values  Description  ISO

char conversion

on / \underline{off}

is character conversion activated ?

Y

singleton warning

\underline{on} / off

warn about named singleton variables ?

N

suspicious warning

\underline{on} / off

warn about suspicious predicate ?

N

multifile warning

\underline{on} / off

warn about unsupported multifile directive ?

N

strict iso

\underline{on} / off

strict ISO behavior ?

N

debug

on / \underline{off}

is the debugger activated ?

Y

double quotes

atom
chars
\underline{codes}
atom no escape
chars no escape
codes no escape

a double quoted constant is returned as: an atom a list of characters a list of character codes as atom but ignore escape sequences as chars but ignore escape sequences as code but ignore escape sequences

{Y \atop N}

back quotes

atom
chars
codes
atom no escape
chars no escape
codes no escape

a back quoted constant is returned as: an atom a list of characters a list of character codes as atom but ignore escape sequences as chars but ignore escape sequences as code but ignore escape sequences

N

unknown

\underline{error} warning fail

a predicate calls an unknown procedure:

an existence error is raised

a message is displayed then fails quietly fails

Y

syntax error

\underline{error} warning fail

a predicate causes a syntax error:

a syntax error is raised

a message is displayed then fails quietly fails

N

os error

\underline{error} warning fail

a predicate causes an O.S. error:

a system error is raised

a message is displayed then fails quietly fails

N

The strict iso flag is introduced to allow a compatibility with other Prolog systems. When turned off the following relaxations apply:

• built-in predicates are found by current predicate/1 (section 8.8.1, page 74).

• the term parser (read/1 and friends) is more indulgent, e.g. 0'' is accepted and returns 39, the escape sequence \s (space) and \e (escape) are accepted.

• the following arithmetic rounding functions: ceiling, floor, round, truncate also accept integers (section 8.6.1, page 66).

Errors

Flag is a variable  instantiation error
Value is a variable  instantiation error

Flag is neither a variable nor an atom type error(atom, Flag)

Flag is an atom but not a valid flag  domain error(prolog flag, Flag)
Value is inappropriate for Flag  domain error(flag value, Flag+Value)

Value is appropriate for Flag but flag Flag is not modifiable

permission error(modify, flag, Flag)

Portability ISO predicate. All ISO flags are implemented.

8.22.2 current prolog flag/2

Templates

     current prolog flag(?flag, ?term)

Description current prolog flag(Flag, Value) succeeds if there exists a Prolog flag that unifies with Flag and whose value unifies with Value. This predicate is re-executable on backtracking. Errors Flag is neither a variable nor an atom type error(atom, Flag)

 Flag is an atom but not a valid flag  domain error(prolog flag, Flag)

Portability ISO predicate.

8.22.3 set bip name/2

Templates

     set bip name(+atom, +arity)

Description set bip name(Functor, Arity) initializes the context of the error (section 6.3.1, page 39) with Functor and Arity (if Arity < 0 only Functor is significant). Errors

 Functor is a variable  instantiation error
 Arity is a variable  instantiation error
 Functor is neither a variable nor an atom  type error(atom, Functor)

Arity is neither a variable nor an integer type error(integer, Arity) Portability GNU Prolog predicate.

8.22.4 current bip name/2

Templates

     current bip name(?atom, ?arity)

Description

current bip name(Functor, Arity) succeeds if Functor and Arity correspond to the context of the

error (section 6.3.1, page 39) (if Arity < 0 only Functor is significant).

Errors

 Functor is neither a variable nor an atom  type error(atom, Functor)

Arity is neither a variable nor an integer type error(integer, Arity)

Portability

GNU Prolog predicate.

8.22.5 write pl state file/1, read pl state file/1

Templates

     write pl state file(+source sink)
     read pl state file(+source sink)

Description

write pl state file(FileName) writes onto FileName all information that influences the parsing of a term (section 8.14, page 104). This allows a sub-process written in Prolog to read this file and then process any Prolog term as done by the parent process. This file can also be passed as argument of the --pl-state option when invoking gplc (section 4.4.3, page 23). More precisely the following elements are saved:

• all operator definitions (section 8.14.10, page 113).

• the character conversion table (section 8.14.12, page 116).

   • the value of char conversion, double quotes, back quotes, singleton warning, suspicious warning

and multifile warning Prolog flags (section 8.22.1, page 148).

read pl state file(FileName) reads (restores) from FileName all information previously saved by

write pl state file/1.

Errors

 FileName is a variable  instantiation error
 FileName is neither a variable nor an atom  type error(atom, FileName)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

8.23 Program state

8.23.1 consult/1, '.'/2 - program consult

Templates

     consult(+atom or atom list)
     '.'(+atom, +atom list)

Description

consult(Files) compiles and loads into memory each file of the list Files. Each file is compiled for byte-code using the GNU Prolog compiler (section 4.4, page 21) then loaded using load/1 (section 8.23.2, page 153). It is possible to specify user as a file name to directly enter the program from the terminal. Files can be also a single file name (i.e. an atom). Refer to the section concerning the consult of a Prolog program for more information (section 4.2.3, page 17).

The final file name of a file is computed using the predicates prolog file name/2 (section 8.26.4, page 159) and absolute file name/2 (section 8.26.1, page 158).

[ File | Files ], i.e. '.'(File, Files) is equivalent to consult([ File | Files ]).

Since version 1.4.0, with the introduction of shebang support, consult/1 ignores the first line of a Prolog source file which directly begins with #. See (section 4.2.4, page 18) for more information about shebang support and PrologScript.

Errors

Files is a partial list or a list with an element E which is a variable

                                             instantiation error

Files is neither a partial list nor a list nor an atom

                                             type error(list, Files)

an element E of the Files list is neither a variable nor an atom

                                             type error(atom, E)

an element E of the Files list is an atom but not a valid pathname

                                             domain error(os path, E)

an element E of the Files list is a valid pathname but does not correspond to an existing source

                                             existence error(source sink, E)

an error occurs executing a directive see call/1 errors (section 7.2.3, page 52)

Portability

GNU Prolog predicates.

8.23.2 load/1

Templates

     load(+atom or atom list)

load(Files) loads into memory each file of the list Files. Each file must have been previously compiled for byte-code using the GNU Prolog compiler (section 4.4, page 21). Files can be also a single file name (i.e. an atom).

The final file name of a file is computed using the predicates absolute file name/2 (section 8.26.1, page 158). If no suffix is given '.wbc' is appended to the file name.

Errors

Files is a partial list or a list with an element E which is a variable

                                             instantiation error

Files is neither a partial list nor a list nor an atom

                                             type error(list, Files)

an element E of the Files list is neither a variable nor an atom

                                             type error(atom, E)

an element E of the Files list is an atom but not a valid pathname

                                             domain error(os path, E)

an element E of the Files list is a valid pathname but does not correspond to an existing source

                                             existence error(source sink, E)

an error occurs executing a directive see call/1 errors (section 7.2.3, page 52)

Portability

GNU Prolog predicate.

8.23.3 listing/1, listing/0

Templates

     listing(+predicate indicator)
     listing(+atom)
     listing

Description

listing(Pred) lists the clauses of the consulted predicate whose predicate indicator is Pred. Pred can also be a single atom in which case all predicates whose name is Pred are listed (of any arity). This predicate uses portray clause/2 (section 8.14.8, page 112) to output the clauses.

listing lists all clauses of all consulted predicates.

Errors

 Pred is a variable  instantiation error

Pred is neither a variable nor predicate indicator or an atom

                                             type error(predicate indicator, Pred)

Portability

8.24 System statistics

8.24.1 statistics/0, statistics/2

Templates

     statistics
     statistics(?atom, ?list)

Description

statistics displays statistics about memory usage and run times.

statistics(Key, Value) unifies Value with the current value of the statistics key Key. Value a list of two elements. Times are in milliseconds, sizes of areas in bytes.

Key  Description  Value

 user time  user CPU time  [SinceStart, SinceLast]
 system time system CPU time  [SinceStart, SinceLast]
 cpu time  total CPU time (user + system)  [SinceStart, SinceLast]
 real time  absolute time  [SinceStart, SinceLast]

local stack local stack sizes (control, environments, choices) [UsedSize, FreeSize]
global stack global stack sizes (compound terms)  [UsedSize, FreeSize]
trail stack trail stack sizes (variable bindings to undo)  [UsedSize, FreeSize]
cstr stack  constraint trail sizes (finite domain constraints) [UsedSize, FreeSize]

 atoms  atom table  [NumberOfAtoms, FreeNumberOfAtoms]

Note that the key runtime is recognized as user time for compatibility purpose.

Errors

 Key is neither a variable nor a valid key  domain error(statistics key, Key)

Value is neither a variable nor a list of two elements

                                             domain error(statistics value, Value)

Value is a list of two elements and an element E is neither a variable nor an integer

                                             type error(integer, E)

Portability

GNU Prolog predicates.

8.24.2 user time/1, system time/1, cpu time/1, real time/1

Templates

     user time(?integer)
     system time(?integer)
     cpu time(?integer)
     real time(?integer)

Description

user time(Time) unifies Time with the user CPU time elapsed since the start of Prolog. system time(Time) unifies Time with the system CPU time elapsed since the start of Prolog. cpu time(Time) unifies Time with the CPU time (user + system) elapsed since the start of Prolog. real time(Time) unifies Time with the absolute time elapsed since the start of Prolog. Errors Time is neither a variable nor an integer type error(integer, Time) Portability GNU Prolog predicates.

8.25 Random number generator

8.25.1 set seed/1, randomize/0

Templates

     set seed(+integer)
     randomize

Description set seed(Seed) reinitializes the random number generator seed with Seed. randomize reinitializes the random number generator. This predicates calls set seed/1 with a random value depending on the absolute time. Errors

 Seed is a variable  instantiation error

Seed is neither a variable nor an integer type error(integer, Seed)

 Seed is an integer < 0  domain error(not less than zero, Seed)

Portability GNU Prolog predicates.

8.25.2 get seed/1

Templates

     get seed(?integer)

Description get seed(Seed) unifies Seed with the current random number generator seed. Errors Seed is neither a variable nor an integer type error(integer, Seed)

 Seed is an integer < 0  domain error(not less than zero, Seed)

Portability GNU Prolog predicate.

8.25.3 random/1

Templates

     random(-float)

Description random(Number) unifies Number with a random floating point number such that 0.0 ≤ Number < 1.0. Errors

 Number is not a variable  uninstantiation error(Number)

Portability GNU Prolog predicate.

8.25.4 random/3

Templates

     random(+number, +number, -number)

Description

random(Base, Max, Number) unifies Number with a random number such that Base ≤ Number < Max.

If both Base and Max are integers Number will be an integer, otherwise Number will be a floating point number. Errors

 Base is a variable  instantiation error

Base is neither a variable nor a number type error(number, Base)

 Max is a variable  instantiation error

Max is neither a variable nor a number type error(number, Max)

 Number is not a variable  uninstantiation error(Number)

Portability GNU Prolog predicate.

8.26 File name processing

8.26.1 absolute file name/2

Templates

     absolute file name(+atom, atom)

Description

absolute file name(File1, File2) succeeds if File2 is the absolute pathname associated with the relative file name File1. File1 can contain $VAR NAME sub-strings. When such a sub-string is encountered, it is expanded with the value of the environment variable whose name is VAR NAME if exists (otherwise no expansion is done). File1 can also begin with a sub-string ~USER NAME/, this is expanded as the home directory of the user USER NAME . If USER NAME does not exist File1 is an invalid pathname. If no USER NAME is given (i.e. File1 begins with ~/) the ~ character is expanded as the value of the environment variable HOME. If the HOME variable is not defined File1 is an invalid pathname. Relative references to the current directory (/./ sub-string) and to the parent directory (/../ sub-strings) are removed and no longer appear in File2. File1 is also invalid if it contains too many /../ consecutive sub-strings (i.e. parent directory relative references). Finally if File1 is user then File2 is also unified with user to allow this predicate to be called on Prolog file names (since user in DEC-10 input/output predicates denotes the current input/output stream).

Under Windows the following applies:

• an alternate recognized form for $VAR NAME is %VAR NAME%.

• when ~ is expanded, if the HOME variable is not defined, ~ is expanded using HOMEDIR and HOMEPATH (if HOMEPATH is not defined then File1 is an invalid pathname).

Most predicates using a file name implicitly call this predicate to obtain the desired file, e.g. open/4.

Errors

 File1 is a variable  instantiation error

File1 is neither a variable nor an atom type error(atom, File1) File2 is neither a variable nor an atom type error(atom, File2)

 File1 is an atom but not a valid pathname  domain error(os path, File1)

Portability

GNU Prolog predicate.

8.26.2 is absolute file name/1, is relative file name/1

Templates

     is absolute file name(+atom)
     is relative file name(+atom)

Description

is absolute file name(PathName) succeeds if PathName is an absolute file name.

Conversely, is relative file name(PathName) succeeds if PathName is not an absolute file name. See absolute file name/2 for information about the syntax of PathName (section 8.26.1, page 158). The current implementation does not check the validity of PathName. If PathName starts with a / (slash) it is considered as absolute. Under Windows, PathName can also start with a \ (backslash) or a drive specification. Errors

 PathName is a variable  instantiation error
 PathName is neither a variable nor an atom  type error(atom, PathName)

Portability GNU Prolog predicates.

8.26.3 decompose file name/4

Templates

     decompose file name(+atom, ?atom, ?atom, ?atom)

Description

decompose file name(File, Directory, Prefix, Suffix) decomposes the pathname File and ex-

tracts the Directory part (characters before the last /), the Prefix part (characters after the last / and before the last . or until the end if there is no suffix) and the Suffix part (characters from the last . to the end of the string). Errors

 File is a variable  instantiation error

File is neither a variable nor an atom type error(atom, File)

 Directory is neither a variable nor an atom  type error(atom, Directory)
 Prefix is neither a variable nor an atom  type error(atom, Prefix)
 Suffix is neither a variable nor an atom  type error(atom, Suffix)

Portability GNU Prolog predicate.

8.26.4 prolog file name/2

Templates

     prolog file name(+atom, ?atom)

Description prolog file name(File1, File2) unifies File2 with the Prolog file name associated with File1. More precisely File2 is computed as follows:

• if File1 has a suffix or if it is user then File2 is unified with File1.

• else if the file whose name is File1 + '.pl' exists then File2 is unified with this name.

• else if the file whose name is File1 + '.pro' exists then File2 is unified with this name.

• else if the file whose name is File1 + '.prolog' exists then File2 is unified with this name.

• else File2 is unified with the name File1 + '.pl'. This predicate uses absolute file name/2 to check the existence of a file (section 8.26.1, page 158). Errors

 File1 is a variable  instantiation error

File1 is neither a variable nor an atom type error(atom, File1) File2 is neither a variable nor an atom type error(atom, File2)

 File1 is an atom but not a valid pathname  domain error(os path, File1)

Portability GNU Prolog predicate.

8.27 Operating system interface

8.27.1 argument counter/1

Templates

     argument counter(?integer)

Description argument counter(Counter) succeeds if Counter is the number of arguments of the command-line. Since the first argument is always the name of the running program, Counter is always ≥ 1. See (section 4.2, page 13) for more information about command-line arguments retrieved under the top level. Errors

 Counter is neither a variable nor an integer  type error(integer, Counter)

Portability GNU Prolog predicate.

8.27.2 argument value/2

Templates

     argument value(+integer, ?atom)

Description argument value(N, Arg) succeeds if the Nth argument on the command-line unifies with Arg. The first argument is always the name of the running program and its number is 0. The number of arguments on the command-line can be obtained using argument counter/1 (section 8.27.1, page 160).

Errors

 N is a variable  instantiation error

N is neither a variable nor an integer type error(integer, N)

 N is an integer < 0  domain error(not less than zero, N)

Arg is neither a variable nor an atom type error(atom, Arg) Portability GNU Prolog predicate.

8.27.3 argument list/1

Templates

     argument list(?list)

Description argument list(Args) succeeds if Args unifies with the list of atoms associated with each argument on the command-line other than the first argument (the name of the running program). Errors Args is neither a partial list nor a list type error(list, Args) Portability GNU Prolog predicate.

8.27.4 environ/2

Templates

     environ(?atom, ?atom)

Description environ(Name, Value) succeeds if Name is the name of an environment variable whose value is Value. This predicate is re-executable on backtracking. Errors Name is neither a variable nor an atom type error(atom, Name) Value is neither a variable nor an atom type error(atom, Value) Portability GNU Prolog predicate.

8.27.5 make directory/1, delete directory/1, change directory/1

Templates

     make directory(+atom)
     delete directory(+atom)
     change directory(+atom)

Description make directory(PathName) creates the directory whose pathname is PathName. delete directory(PathName) removes the directory whose pathname is PathName. change directory(PathName) sets the current directory to the directory whose pathname is PathName. See absolute file name/2 for information about the syntax of PathName (section 8.26.1, page 158). Errors

 PathName is a variable  instantiation error
 PathName is neither a variable nor an atom  type error(atom, PathName)
 PathName is an atom but not a valid pathname  domain error(os path, PathName)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability GNU Prolog predicates.

8.27.6 working directory/1

Templates

     working directory(?atom)

Description working directory(PathName) succeeds if PathName is the pathname of the current directory. Errors

 PathName is neither a variable nor an atom  type error(atom, PathName)

Portability GNU Prolog predicate.

8.27.7 directory files/2

Templates

     directory files(+atom, ?list)

Description

directory files(PathName, Files) succeeds if Files is the list of all entries (files, sub-directories,…) in the directory whose pathname is PathName. See absolute file name/2 for information about the syntax of PathName (section 8.26.1, page 158).

Errors

 PathName is a variable  instantiation error
 PathName is neither a variable nor an atom  type error(atom, PathName)
 PathName is an atom but not a valid pathname  domain error(os path, PathName)

Files is neither a partial list nor a list type error(list, Files) an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.27.8 rename file/2

Templates

     rename file(+atom, +atom)

Description

rename file(PathName1, PathName2) renames the file or directory whose pathname is PathName1 to

PathName2. See absolute file name/2 for information about the syntax of PathName1 and PathName2

(section 8.26.1, page 158).

Errors

 PathName1 is a variable  instantiation error
 PathName1 is neither a variable nor an atom  type error(atom, PathName1)
 PathName1 is an atom but not a valid pathname domain error(os path, PathName1)
 PathName2 is a variable  instantiation error
 PathName2 is neither a variable nor an atom  type error(atom, PathName2)
 PathName2 is an atom but not a valid pathname domain error(os path, PathName2)

an operating system error occurs and value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.27.9 delete file/1, unlink/1

     delete file(PathName)
     unlink(PathName)

Description

delete file(PathName) removes the existing file whose pathname is PathName.

unlink/1 is similar to delete file/1 except that it never causes a system error (e.g. if PathName does not refer to an existing file).

See absolute file name/2 for information about the syntax of PathName (section 8.26.1, page 158).

Errors

 PathName is a variable  instantiation error
 PathName is neither a variable nor an atom  type error(atom, PathName)
 PathName is an atom but not a valid pathname  domain error(os path, PathName)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicates.

8.27.10 file permission/2, file exists/1

Templates

     file permission(+atom, +atom)
     file permission(+atom, +atom list)
     file exists(+atom)

Description

file permission(PathName, Permission) succeeds if PathName is the pathname of an existing file (or

directory) whose permissions include Permission.

File permissions: Permission can be a single permission or a list of permissions. A permission is an atom among:

• read: the file or directory can be read.

• write: the file or directory can be written.

• execute: the file can be executed.

• search: the directory can be searched.

If PathName does not exists or if its permissions do not include Permission this predicate fails.

file exists(PathName) is equivalent to file permission(PathName, []), i.e. it succeeds if PathName

is the pathname of an existing file (or directory).

See absolute file name/2 for information about the syntax of PathName (section 8.26.1, page 158).

Errors

PathName is a variable  instantiation error
PathName is neither a variable nor an atom  type error(atom, PathName)
PathName is an atom but not a valid pathname  domain error(os path, PathName)

Permission is a partial list or a list with an element which is a variable

                                           instantiation error

Permission is neither an atom nor partial list or a list

                                           type error(list, Permission)

an element E of the Permission list is neither a variable nor an atom

                                           type error(atom, E)

an element E of the Permission is an atom but not a valid permission

                                           domain error(os file permission,
                                           Permission)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                           system error(atom explaining the error)

Portability

GNU Prolog predicates.

8.27.11 file property/2

Templates

    file property(+atom, ?os file property)

Description

file property(PathName, Property) succeeds if PathName is the pathname of an existing file (or directory) and if Property unifies with one of the properties of the file. This predicate is re-executable on backtracking.

File properties:

• absolute file name(File): File is the absolute file name of PathName (section 8.26.1, page 158).

• real file name(File): File is the real file name of PathName (follows symbolic links).

  • type(Type): Type is the type of PathName. Possible values are: regular, directory, fifo, socket,
    character device, block device or unknown.

• size(Size): Size is the size (in bytes) of PathName.

• permission(Permission): Permission is a permission of PathName (section 8.27.10, page 164).

• last modification(DT): DT is the last modification date and time (section 8.27.14, page 167).

See absolute file name/2 for information about the syntax of PathName (section 8.26.1, page 158).

PathName is a variable  instantiation error
PathName is neither a variable nor an atom  type error(atom, PathName)
PathName is an atom but not a valid pathname  domain error(os path, PathName)

Property is neither a variable nor a file property term

                                           domain error(os file property, Property)
Property = absolute file name(E),
real file name(E), type(E) or permission(E)

and E is neither a variable nor an atom

                                           type error(atom, E)
Property = last modification(DateTime)

and DateTime is neither a variable nor a compound term

                                           type error(compound, DateTime)
Property = last modification(DateTime)

and DateTime is a compound term but not a structure dt/6

                                           domain error(date time, DateTime)
Property = size(E) or
last modification(DateTime) and DateTime is

a structure dt/6 but an element E is neither a variable nor an integer

                                           type error(integer, E)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                           system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.27.12 temporary name/2

Templates

    temporary name(+atom, ?atom)

Description

temporary name(Template, PathName) creates a unique file name PathName whose pathname begins by Template. Template should contain a pathname with six trailing Xs. PathName is Template with the six Xs replaced with a letter and the process identifier. This predicate is an interface to the C Unix function

mktemp(3).

See absolute file name/2 for information about the syntax of Template (section 8.26.1, page 158).

Errors

Template is a variable  instantiation error
Template is neither a variable nor an atom  type error(atom, Template)
Template is an atom but not a valid pathname  domain error(os path, Template)
PathName is neither a variable nor an atom  type error(atom, PathName)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                           system error(atom explaining the error)

Portability

8.27.13 temporary file/3

Templates

     temporary file(+atom, +atom, ?atom)

Description

temporary file(Directory, Prefix, PathName) creates a unique file name PathName whose pathname

begins by Directory/Prefix. If Directory is the empty atom '' a standard temporary directory will be used (e.g. /tmp). Prefix can be the empty atom ''. This predicate is an interface to the C Unix

function tempnam(3).

See absolute file name/2 for information about the syntax of Directory (section 8.26.1, page 158).

Errors

 Directory is a variable  instantiation error
 Directory is neither a variable nor an atom  type error(atom, Directory)
 Directory is an atom but not a valid pathname domain error(os path, Directory)
 Prefix is a variable  instantiation error
 Prefix is neither a variable nor an atom  type error(atom, Prefix)
 PathName is neither a variable nor an atom  type error(atom, PathName)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.27.14 date time/1

Templates

     date time(?compound)

Description

date time(DateTime) unifies DateTime with a compound term containing the current date and time.

DateTime is a structure dt(Year, Month, Day, Hour, Minute, Second). Each sub-argument of the

term dt/6 is an integer.

Errors

DateTime is neither a variable nor a compound term

                                             type error(compound, DateTime)

DateTime is a compound term but not a structure dt/6

                                             domain error(date time, DateTime)

DateTime is a structure dt/6 and an element E is neither a variable nor an integer

                                             type error(integer, E)

Portability

8.27.15 host name/1

Templates

     host name(?atom)

Description

host name(HostName) unifies HostName with the name of the host machine executing the current GNU Prolog process. If the sockets are available (section 8.28.1, page 176), the name returned will be fully qualified. In that case, host name/1 will also succeed if HostName is instantiated to the unqualified name (or an alias) of the machine.

Errors

 HostName is neither a variable nor an atom  type error(atom, HostName)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.27.16 os version/1

Templates

     os version(?atom)

Description

os version(OSVersion) unifies OSVersion with the operating system version of the machine executing the current GNU Prolog process.

Errors

 OSVersion is neither a variable nor an atom  type error(atom, OSVersion)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.27.17 architecture/1

Templates

     architecture(?atom)

architecture(Architecture) unifies Architecture with the name of the machine executing the current GNU Prolog process.

Errors

 Architecture is neither a variable nor an atom type error(atom, Architecture)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.27.18 shell/2, shell/1, shell/0

Templates

     shell(+atom, ?integer)
     shell(+atom)
     shell

Description

shell(Command, Status) invokes a new shell (named by the SHELL environment variable) passing Command for execution and unifies Status with the result of the execution. If Command is the empty atom '' a new interactive shell is executed. The control is returned to Prolog upon termination of the called process.

shell(Command) is equivalent to shell(Command, 0).
shell is equivalent to shell('', 0).

Errors

 Command is a variable  instantiation error
 Command is neither a variable nor an atom  type error(atom, Command)
 Status is neither a variable nor an integer  type error(integer, Status)

Portability

GNU Prolog predicates.

8.27.19 system/2, system/1

Templates

     system(+atom, ?integer)
     system(+atom)

Description

system(Command, Status) invokes a new default shell passing Command for execution and unifies Status with the result of the execution. The control is returned to Prolog upon termination of the shell process. This predicate is an interface to the C Unix function system(3).

system(Command) is equivalent to system(Command, 0).

Errors

 Command is a variable  instantiation error
 Command is neither a variable nor an atom  type error(atom, Command)
 Status is neither a variable nor an integer  type error(integer, Status)

Portability

GNU Prolog predicates.

8.27.20 spawn/3, spawn/2

Templates

     spawn(+atom, +atom list, ?integer)
     spawn(+atom, +atom list)

Description

spawn(Command, Arguments, Status) executes Command passing as arguments of the command-line each element of the list Arguments and unifies Status with the result of the execution. The control is returned to Prolog upon termination of the command.

spawn(Command, Arguments) is equivalent to spawn(Command, Arguments, 0).

Errors

 Command is a variable  instantiation error
 Command is neither a variable nor an atom  type error(atom, Command)

Arguments is a partial list or a list with an element which is a variable

                                             instantiation error
 Arguments is neither a partial list nor a list  type error(list, Arguments)

an element E of the Arguments list is neither a variable nor an atom

                                             type error(atom, E)
 Status is neither a variable nor an integer  type error(integer, Status)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicates.

8.27.21 popen/3

Templates

     popen(+atom, +io mode, -stream)

Description

popen(Command, Mode, Stream) invokes a new default shell (by creating a pipe) passing Command for execution and associates a stream either to the standard input or the standard output of the created process. if Mode is read (resp. write) an input (resp. output) stream is created and Stream is unified with the stream-term associated. Writing to the stream writes to the standard input of the command while reading from the stream reads the command’s standard output. The stream must be closed using close/2 (section 8.10.7, page 82). This predicate is an interface to the C Unix function popen(3).

Errors

 Command is a variable  instantiation error
 Command is neither a variable nor an atom  type error(atom, Command)
 Mode is a variable  instantiation error

Mode is neither a variable nor an atom type error(atom, Mode)

 Mode is an atom but neither read nor write.  domain error(io mode, Mode)
 Stream is not a variable  uninstantiation error(Stream)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.27.22 exec/5, exec/4

Templates

     exec(+atom, -stream, -stream, -stream, -integer)
     exec(+atom, -stream, -stream, -stream)

Description

exec(Command, StreamIn, StreamOut, StreamErr, Pid) invokes a new default shell passing Command

for execution and associates streams to standard streams of the created process. StreamIn is unified with the stream-term associated with the standard input stream of Command (it is an output stream). StreamOut is unified with the stream-term associated with the standard output stream of Command (it is an input stream). StreamErr is unified with the stream-term associated with the standard error stream of Command (it is an input stream). Pid is unified with the process identifier of the new process. This information is only useful if it is necessary to obtain the status of the execution using wait/2 (section 8.27.25, page 173). Until a call to wait/2 is done the process remains in the system processes table (as a zombie process if terminated). For this reason, if the status is not needed it is preferable to

use exec/4.

exec/4 is similar to exec/5 but the process is removed from system processes as soon as it is terminated.

 Command is a variable  instantiation error
 Command is neither a variable nor an atom  type error(atom, Command)
 StreamIn is not a variable  uninstantiation error(StreamIn)
 StreamOut is not a variable  uninstantiation error(StreamOut)
 StreamErr is not a variable  uninstantiation error(StreamErr)
 Pid is not a variable  uninstantiation error(Pid)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicates.

8.27.23 fork prolog/1

Templates

     fork prolog(-integer)

Description

fork prolog(Pid) creates a child process that differs from the parent process only in its PID. In the parent process Pid is unified with the PID of the child while in the child process Pid is unified with 0. In the parent process, the status of the child process can be obtained using wait/2 (section 8.27.25, page 173). Until a call to wait/2 is done the child process remains in the system processes table (as a zombie process if terminated). This predicate is an interface to the C Unix function fork(2).

Errors

 Pid is not a variable  uninstantiation error(Pid)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.27.24 create pipe/2

Templates

     create pipe(-stream, -stream)

Description

create pipe(StreamIn, StreamOut) creates a pair of streams pointing to a pipe inode. StreamIn is unified with the stream-term associated with the input side of the pipe and StreamOut is unified with the stream-term associated with output side. This predicate is an interface to the C Unix function pipe(2).

 StreamIn is not a variable  uninstantiation error(StreamIn)
 StreamOut is not a variable  uninstantiation error(StreamOut)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability GNU Prolog predicate.

8.27.25 wait/2 Templates

     wait(+integer, ?integer)

Description wait(Pid, Status) waits for the child process whose identifier is Pid to terminate. Status is then unified with the exit status. This predicate is an interface to the C Unix function waitpid(2). Errors

 Pid is a variable  instantiation error

Pid is neither a variable nor an integer type error(integer, Pid)

 Status is neither a variable nor an integer  type error(integer, Status)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability GNU Prolog predicate.

8.27.26 prolog pid/1

Templates

     prolog pid(?integer)

Description prolog pid(Pid) unifies Pid with the process identifier of the current GNU Prolog process. Errors Pid is neither a variable nor an integer type error(integer, Pid) Portability GNU Prolog predicate.

8.27.27 send signal/2

Templates

     send signal(+integer, +integer)
     send signal(+integer, +atom)

Description

send signal(Pid, Signal) sends Signal to the process whose identifier is Pid. Signal can be specified directly as an integer or symbolically as an atom. Allowed atoms depend on the machine (e.g. 'SIGINT',

'SIGQUIT', 'SIGKILL', 'SIGUSR1', 'SIGUSR2', 'SIGALRM',. . . ). This predicate is an interface to the C

Unix function kill(2).

Errors

 Pid is a variable  instantiation error

Pid is neither a variable nor an integer type error(integer, Pid)

 Signal is a variable  instantiation error

Signal is neither a variable nor an integer or an atom

                                             type error(integer, Signal)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.27.28 sleep/1

Templates

     sleep(+number)

Description

sleep(Seconds) puts the GNU Prolog process to sleep for Seconds seconds. Seconds can be an integer or a floating point number (in which case it can be < 1). This predicate is an interface to the C Unix

function usleep(3).

Errors

 Seconds is a variable  instantiation error
 Seconds is neither a variable nor a number  type error(number, Seconds)
 Seconds is a number < 0  domain error(not less than zero, Seconds)

Portability

8.27.29 select/5

Templates

    select(+list, ?list, +list, ?list, +number)

Description

select(Reads, ReadyReads, Writes, ReadyWrites, TimeOut) waits for a number of streams (or file

descriptors) to change status. ReadyReads is unified with the list of elements listed in Reads that have characters available for reading. Similarly ReadyWrites is unified with the list of elements of Writes that are ok for immediate writing. The elements of Reads and Writes are either stream-terms or aliases or integers considered as file descriptors, e.g. for sockets (section 8.28, page 176). Streams that must be tested with select/5 should not be buffered. This can be done at the opening using open/4 (section 8.10.6, page 81) or later using set stream buffering/2 (section 8.10.27, page 93). TimeOut is an upper bound on the amount of time (in milliseconds) elapsed before select/5 returns. If TimeOut ≤ 0 (no timeout) select/5 waits until something is available (either or reading or for writing) and thus can block indefinitely. This predicate is an interface to the C Unix function select(2).

Errors

Reads (or Writes) is a partial list or a list with an element E which is a variable

                                           instantiation error

Reads is neither a partial list nor a list type error(list, Reads) Writes is neither a partial list nor a list type error(list, Writes)

ReadyReads is neither a partial list nor a list  type error(list, ReadyReads)
ReadyWrites is neither a partial list nor a list  type error(list, ReadyWrites)

an element E of the Reads (or Writes) list is neither a stream-term or alias nor an integer

                                           domain error(stream or alias, E)

an element E of the Reads (or Writes) list is not a selectable item

                                           domain error(selectable item, E)

an element E of the Reads (or Writes) list is an integer < 0

                                           domain error(not less than zero, E)

an element E of the Reads (or Writes) list is a stream-tern or alias not associated with an open stream

                                           existence error(stream, E)

an element E of the Reads list is associated with an output stream

                                           permission error(input, stream, E)

an element E of the Writes list is associated with an input stream

                                           permission error(output, stream, E)
TimeOut is a variable  instantiation error
TimeOut is neither a variable nor a number  type error(number, TimeOut)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                           system error(atom explaining the error)

Portability

8.28 Sockets input/output

8.28.1 Introduction

This set of predicates provides a way to manipulate sockets. The predicates are straightforward interfaces to the corresponding BSD-type socket functions. This facility is available if the sockets part of GNU Prolog has been installed. A reader familiar with BSD sockets will understand them immediately otherwise a study of sockets is needed.

The domain is either the atom 'AF INET' or 'AF UNIX' corresponding to the same domains in BSD-type sockets.

An address is either of the form 'AF INET'(HostName, Port) or 'AF UNIX'(SocketName). HostName is

an atom denoting a machine name, Port is a port number and SocketName is an atom denoting a socket.

By default, streams associated with sockets are block buffered. The predicate set stream buffering/2 (section 8.10.27, page 93) can be used to change this mode. They are also text streams by default. Use set stream type/2 (section 8.10.25, page 92) to change the type if binary streams are needed.

8.28.2 socket/2

Templates

     socket(+socket domain, -integer)

Description

socket(Domain, Socket) creates a socket whose domain is Domain (section 8.28, page 176) and unifies Socket with the descriptor identifying the socket. This predicate is an interface to the C Unix function

socket(2).

Errors

 Domain is a variable  instantiation error
 Domain is neither a variable nor an atom  type error(atom, Domain)
 Domain is an atom but not a valid socket domain domain error(socket domain, Domain)
 Socket is not a variable  uninstantiation error(Socket)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.28.3 socket close/1

Templates

     socket close(+integer)

socket close(Socket) closes the socket whose descriptor is Socket. This predicate should not be used if Socket has given rise to a stream, e.g. by socket connect/4 (section 8.28.5, page 178). In that case simply use close/2 (section 8.10.7, page 82) on the associated stream.

Errors

Socket is a variable  instantiation error
Socket is neither a variable nor an integer  type error(integer, Socket)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                           system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.28.4 socket bind/2

Templates

    socket bind(+integer, +socket address)

Description

socket bind(Socket, Address) binds the socket whose descriptor is Socket to the address specified by Address (section 8.28, page 176). If Address if of the form 'AF INET'(HostName, Port) and if HostName is uninstantiated then it is unified with the current machine name. If Port is uninstantiated, it is unified to a port number picked by the operating system. This predicate is an interface to the C Unix function bind(2).

Errors

Socket is a variable  instantiation error
Socket is neither a variable nor an integer  type error(integer, Socket)
Address is a variable  instantiation error
Address is neither a variable nor a valid address domain error(socket address, Address)
Address = 'AF UNIX'(E) and E is a variable  instantiation error
Address = 'AF UNIX'(E) or 'AF INET'(E,  )

and E is neither a variable nor an atom

                                           type error(atom, E)
Address = 'AF UNIX'(E) and E is an atom but

not a valid pathname

                                           domain error(os path, E)
Address = 'AF INET'( , E) and E is neither a

variable nor an integer

                                           type error(integer, E)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                           system error(atom explaining the error)

Portability

8.28.5 socket connect/4

Templates

     socket connect(+integer, +socket address, -stream, -stream)

Description

socket connect(Socket, Address, StreamIn, StreamOut) connects the socket whose descriptor is

Socket to the address specified by Address (section 8.28, page 176). StreamIn is unified with a streamterm associated with the input of the connection (it is an input stream). Reading from this stream gets data from the socket. StreamOut is unified with a stream-term associated with the output of the connection (it is an output stream). Writing to this stream sends data to the socket. The use of select/5 can be useful (section 8.27.29, page 175). This predicate is an interface to the C Unix function

connect(2).

Errors

 Socket is a variable  instantiation error
 Socket is neither a variable nor an integer  type error(integer, Socket)
 Address is a variable  instantiation error
 Address is neither a variable nor a valid address domain error(socket address, Address)
 Address = 'AF UNIX'(E) or 'AF INET'(E,  ) or
 Address = 'AF INET'( , E) and E is a variable

                                             instantiation error
 Address = 'AF UNIX'(E) or 'AF INET'(E,  )

and E is neither a variable nor an atom

                                             type error(atom, E)
 Address = 'AF UNIX'(E) and E is an atom but

not a valid pathname

                                             domain error(os path, E)
 Address = 'AF INET'( , E) and E is neither a

variable nor an integer

                                             type error(integer, E)
 StreamIn is not a variable  uninstantiation error(StreamIn)
 StreamOut is not a variable  uninstantiation error(StreamOut)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.28.6 socket listen/2

Templates

     socket listen(+integer, +integer)

Description

socket listen(Socket, Length) defines the socket whose descriptor is Socket to have a maximum backlog queue of Length pending connections. This predicate is an interface to the C Unix function

listen(2).

 Socket is a variable  instantiation error
 Socket is neither a variable nor an integer  type error(integer, Socket)
 Length is a variable  instantiation error
 Length is neither a variable nor an integer  type error(integer, Length)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicate.

8.28.7 socket accept/4, socket accept/3

Templates

     socket accept(+integer, -atom, -stream, -stream)
     socket accept(+integer, -stream, -stream)

Description

socket accept(Socket, Client, StreamIn, StreamOut) extracts the first connection to the socket

whose descriptor is Socket. If the domain is 'AF INET', Client is unified with an atom whose name is the Internet host address in numbers-and-dots notation of the connecting machine. StreamIn is unified with a stream-term associated with the input of the connection (it is an input stream). Reading from this stream gets data from the socket. StreamOut is unified with a stream-term associated with the output of the connection (it is an output stream). Writing to this stream sends data to the socket. The use of select/5 can be useful (section 8.27.29, page 175). This predicate is an interface to the C Unix function

accept(2).
socket accept(Socket, StreamIn, StreamOut) is equivalent to socket accept(Socket,  ,
StreamIn, StreamOut).

Errors

 Socket is a variable  instantiation error
 Socket is neither a variable nor an integer  type error(integer, Socket)
 Client is not a variable  uninstantiation error(Client)
 StreamIn is not a variable  uninstantiation error(StreamIn)
 StreamOut is not a variable  uninstantiation error(StreamOut)

an operating system error occurs and the value of the os error Prolog flag is error (section 8.22.1, page 148)

                                             system error(atom explaining the error)

Portability

GNU Prolog predicates.

8.28.8 hostname address/2

     hostname address(+atom, ?atom)
     hostname address(?atom, +atom)

Description hostname address(HostName, HostAddress) succeeds if the Internet host address in numbers-and-dots notation of HostName is HostAddress. Hostname can be given as a fully qualified name, or an unqualified name or an alias of the machine. The predicate will fail if the machine name or address cannot be resolved. Errors

 HostName and HostAddress are variables  instantiation error
 HostName is neither a variable nor an atom  type error(atom, HostName)
 HostAddress is neither a variable nor an atom  type error(atom, HostAddress)
 Address is neither a variable nor a valid address domain error(socket address, Address)

Portability GNU Prolog predicate.

8.29 Linedit management

The following predicates are only available if the linedit part of GNU Prolog has been installed.

8.29.1 get linedit prompt/1

Templates

     get linedit prompt(?atom)

Description get linedit prompt(Prompt) succeeds if Prompt is the current linedit prompt, e.g. the top-level prompt is '| ?-'. By default all other reads have an empty prompt. Errors Prompt is neither a variable nor an atom type error(atom, Pred) Portability GNU Prolog predicate.

8.29.2 set linedit prompt/1

Templates

     set linedit prompt(+atom)

Description set linedit prompt(Prompt) sets the current linedit prompt to Prompt. This prompt will be displayed for reads from a terminal (except for top-level reads). Errors

 Prompt is a variable  instantiation error

Prompt is neither a variable nor an atom type error(atom, Pred) Portability GNU Prolog predicate.

8.29.3 add linedit completion/1

Templates

     add linedit completion(+atom)

Description add linedit completion(Word) adds Word in the list of completion words maintained by linedit (section 4.2.6, page 18). Only words containing letters, digits and the underscore character are added (if Word does not respect this restriction the predicate fails). Errors

 Word is a variable  instantiation error

Word is neither a variable nor an atom type error(atom, Word) Portability GNU Prolog predicate.

8.29.4 find linedit completion/2

Templates

     find linedit completion(+atom, ?atom)

Description

find linedit completion(Prefix, Word) succeeds if Word is a word beginning by Prefix and belongs

to the list of completion words maintained by linedit (section 4.2.6, page 18). This predicate is reexecutable on backtracking. Errors

 Prefix is a variable  instantiation error
 Prefix is neither a variable nor an atom  type error(atom, Prefix)

Word is neither a variable nor an atom type error(atom, Word) Portability GNU Prolog predicate.

182  8 PROLOG BUILT-IN PREDICATES

9 Finite domain solver and built-in predicates

9.1 Introduction

The finite domain (FD) constraint solver extends Prolog with constraints over FD. This facility is available if the FD part of GNU Prolog has been installed. The solver is an instance of the Constraint Logic Programming scheme introduced by Jaffar and Lassez in 1987 [7]. Constraints on FD are solved using propagation techniques, in particular arc-consistency (AC). The interested reader can refer to “Constraint Satisfaction in Logic Programming” of P. Van Hentenryck (1989) [8]. The solver is based on the clp(FD) solver [4]. The GNU Prolog FD solver offers arithmetic constraints, boolean constraints, reified constraints and symbolic constraints on an new kind of variables: Finite Domain variables.

9.1.1 Finite Domain variables

A new type of data is introduced: FD variables which can only take values in their domains. The initial domain of an FD variable is 0..fd max integer where fd max integer represents the greatest value that any FD variable can take. The predicate fd max integer/1 returns this value which may be different from the max integer Prolog flag (section 8.22.1, page 148). The domain of an FD variable X is reduced step by step by constraints in a monotonic way: when a value has been removed from the domain of X it will never reappear in the domain of X. An FD variable is fully compatible with both Prolog integers and Prolog variables. Namely, when an FD variable is expected by an FD constraint it is possible to pass a Prolog integer (considered as an FD variable whose domain is a singleton) or a Prolog variable (immediately bound to an initial range 0..fd max integer). This avoids the need for specific type declaration. Although it is not necessary to declare the initial domain of an FD variable (since it will be bound 0..fd max integer when appearing for the fist time in a constraint) it is advantageous to do so and thus reduce as soon as possible the size of its domain: particularly because GNU Prolog, for efficiency reasons, does not check for overflows. For instance, without any preliminary domain definitions for X, Y and Z, the non-linear constraint X*Y#=Z will fail due to an overflow when computing the upper bound of the domain of Z: fd max integer × fd max integer. This overflow causes a negative result for the upper bound and the constraint then fails.

There are two internal representations for an FD variable:

• interval representation: only the min and the max of the variable are maintained. In this representation it is possible to store values included in 0..fd max integer.

• sparse representation: an additional bit-vector is used to store the set of possible values for the variable (i.e. the domain). In this representation it is possible to store values included in 0..vector max. By default vector max is set to 127. This value can be redefined via an environment variable VECTORMAX or via the built-in predicate fd set vector max/1 (section 9.2.3, page 185). The predicate fd vector max/1 returns the current value of vector max (section 9.2.1, page 184).

The initial representation for an FD variable X is always an interval representation and is switched to a sparse representation when a “hole” appears in the domain (e.g. due to an inequality constraint). Once a variable uses a sparse representation it will not switch back to an interval representation even if there are no longer holes in its domain. When this switching occurs some values in the domain of X can be lost since vector max is less than fd max integer. We say that “X is extra-constrained” since X is constrained by the solver to the domain 0..vector max (via an imaginary constraint X #=< vector max ). An extra cstr is associated with each FD variable to indicate that values have been lost due to the switch to a sparse representation. This flag is updated on every operations. The domain of an extra-constrained FD variable is output followed by the @ symbol. When a constraint fails on a extra-constrained variable

a message Warning:  Vector too small - maybe lost solutions (FD Var:N) is displayed (N is the

address of the involved variable).

Example 1 (vector max = 127):

Constraint on X Domain of X    extra cstr Lost values

 X #=< 512      0..512            off  none
 X #\= 10       0..9:11..127      on      128..512
 X #=< 100      0..9:11..100      off  none

In this example, when the constraint X #\= 10 is posted some values are lost, the extra cstr is then switched on. However, posting the constraint X #=< 100 will turn off the flag (no values are lost).

Example 2:

Constraint on X Domain of X                     extra cstr Lost values

 X #=< 512      0..512                             off  none
 X #\= 10       0..9:11..127                       on      128..512
 X #>= 256      Warning:  Vector too small...      on      128..512

In this example, the constraint X #>= 256 fails due to the lost of 128..512 so a message is displayed onto the terminal. The solution would consist in increasing the size of the vector either by setting the environment variable VECTORMAX (e.g. to 512) or using fd set vector max(512).

Finally, bit-vectors are not dynamic, i.e. all vectors have the same size (0..vector max). So the use of fd set vector max/1 is limited to the initial definition of vector sizes and must occur before any constraint. As seen before, the solver tries to display a message when a failure occurs due to a too short vector max. Unfortunately, in some cases it cannot detect the lost of values and no message is emitted. So the user should always take care to this parameter to be sure that it is large to encode any vector.

9.2 FD variable parameters

9.2.1 fd max integer/1

Templates

     fd max integer(?integer)

Description

fd max integer(N) succeeds if N is the current value of fd max integer (section 9.1, page 183).

Errors

N is neither a variable nor an integer type error(integer, N)

Portability

9.2.2 fd vector max/1

Templates

     fd vector max(?integer)

Description fd vector max(N) succeeds if N is the current value of vector max (section 9.1, page 183). Errors N is neither a variable nor an integer type error(integer, N) Portability GNU Prolog predicate.

9.2.3 fd set vector max/1

Templates

     fd set vector max(+integer)

Description fd set vector max(N) initializes vector max based on the value N (section 9.1, page 183). More precisely, on 32 bit machines, vector max is set to the smallest value of (32*k)-1 which is ≥ N. Errors

 N is a variable  instantiation error

N is neither a variable nor an integer type error(integer, N)

 N is an integer < 0  domain error(not less than zero, N)

Portability GNU Prolog predicate.

9.3 Initial value constraints

9.3.1 fd domain/3, fd domain bool/1

Templates

     fd domain(+fd variable list or fd variable, +integer, +integer)
     fd domain(?fd variable, +integer, +integer)
     fd domain bool(+fd variable list)
     fd domain bool(?fd variable)

Description

fd domain(Vars, Lower, Upper) constraints each element X of Vars to take a value in Lower..Upper.

This predicate is generally used to set the initial domain of variables to an interval. Vars can be also a single FD variable (or a single Prolog variable).

fd domain bool(Vars) is equivalent to fd domain(Vars, 0, 1) and is used to declare boolean FD variables.

Errors

Vars is not a variable but is a partial list instantiation error Vars is neither a variable nor an FD variable nor an integer nor a list

                                             type error(list, Vars)

an element E of the Vars list is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, E)
 Lower is a variable  instantiation error

Lower is neither a variable nor an integer type error(integer, Lower)

 Upper is a variable  instantiation error

Upper is neither a variable nor an integer type error(integer, Upper)

Portability

GNU Prolog predicate.

9.3.2 fd domain/2

Templates

     fd domain(+fd variable list, +integer list)
     fd domain(?fd variable, +integer list)

Description

fd domain(Vars, Values) constraints each element X of the list Vars to take a value in the list Values. This predicate is generally used to set the initial domain of variables to a set of values. The domain of each variable of Vars uses a sparse representation. Vars can be also a single FD variable (or a single Prolog variable).

Errors

Vars is not a variable but is a partial list instantiation error Vars is neither a variable nor an FD variable nor an integer nor a list

                                             type error(list, Vars)

an element E of the Vars list is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, E)

Values is a partial list or a list with an element E which is a variable

                                             instantiation error

Values is neither a partial list nor a list type error(list, Values) an element E of the Values list is neither a variable nor an integer

                                             type error(integer, E)

Portability

9.4 Type testing

9.4.1 fd var/1, non fd var/1, generic var/1, non generic var/1

Templates

     fd var(?term)
     non fd var(?term)

                                                 generic var(?term)
                                                 non generic var(?term)

Description fd var(Term) succeeds if Term is currently an FD variable. non fd var(Term) succeeds if Term is currently not an FD variable (opposite of fd var/1). generic var(Term) succeeds if Term is either a Prolog variable or an FD variable. non generic var(Term) succeeds if Term is neither a Prolog variable nor an FD variable (opposite of

generic var/1).

Errors None. Portability GNU Prolog predicate.

9.5 FD variable information

These predicate allow the user to get some information about FD variables. They are not constraints, they only return the current state of a variable.

9.5.1 fd min/2, fd max/2, fd size/2, fd dom/2

Templates

     fd min(+fd variable, ?integer)
     fd max(+fd variable, ?integer)
     fd size(+fd variable, ?integer)
     fd dom(+fd variable, ?integer list)

Description fd min(X, N) succeeds if N is the minimal value of the current domain of X. fd max(X, N) succeeds if N is the maximal value of the current domain of X. fd size(X, N) succeeds if N is the number of elements of the current domain of X. fd dom(X, Values) succeeds if Values is the list of values of the current domain of X.

Errors

 X is a variable  instantiation error

X is neither an FD variable nor an integer type error(fd variable, X) N is neither a variable nor an integer type error(integer, N) an element E of the Vars list is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, E)

Values is neither a partial list nor a list type error(list, Values) Portability GNU Prolog predicate.

9.5.2 fd has extra cstr/1, fd has vector/1, fd use vector/1

Templates

     fd has extra cstr(+fd variable)
     fd has vector(+fd variable)
     fd use vector(+fd variable)

Description fd has extra cstr(X) succeeds if the extra cstr of X is currently on (section 9.1, page 183). fd has vector(X) succeeds if the current domain of X uses a sparse representation (section 9.1, page 183). fd use vector(X) enforces a sparse representation for the domain of X (section 9.1, page 183). Errors

 X is a variable  instantiation error

X is neither an FD variable nor an integer type error(fd variable, X) Portability GNU Prolog predicates.

9.6 Arithmetic constraints

9.6.1 FD arithmetic expressions An FD arithmetic expression is a Prolog term built from integers, variables (Prolog or FD variables), and functors (or operators) that represent arithmetic functions. The following table details the components of an FD arithmetic expression:

FD Expression  Result

Prolog variable domain 0..fd max integer

FD variable X        domain of X
integer number N     domain N..N
+ E                same as E
- E                opposite of E

E1 + E2  sum of E1 and E2

E1 - E2 subtraction of E2 from E1 E1 * E2 multiplication of E1 by E2 E1 / E2 integer division of E1 by E2 (only succeeds if the remainder is 0) E1 ** E2 E1 raised to the power of E2 (E1 or E2 must be an integer)

min(E1,E2)  minimum of E1 and E2
max(E1,E2)  maximum of E1 and E2
dist(E1,E2)  distance, i.e. |E1 - E2|

E1 // E2 quotient of the integer division of E1 by E2 E1 rem E2 remainder of the integer division of E1 by E2

quot rem(E1,E2,R)

quotient of the integer division of E1 by E2

(R is the remainder of the integer division of E1 by E2)

FD expressions are not restricted to be linear. However non-linear constraints usually yield less constraint propagation than linear constraints.

+, -, *, /, //, rem and ** are predefined infix operators. + and - are predefined prefix operators (section 8.14.10, page 113).

Errors

a sub-expression is of the form ** E and E is a variable

                                           instantiation error

a sub-expression E is neither a variable nor an integer nor an FD arithmetic functor

                                           type error(fd evaluable, E)
an expression is too complex  resource error(too big fd constraint)

9.6.2 Partial AC: (#=)/2 - constraint equal, (#\=)/2 - constraint not equal, (#<)/2 - constraint less than, (#=<)/2 - constraint less than or equal, (#>)/2 - constraint greater than, (#>=)/2 - constraint greater than or equal

Templates

    #=(?fd evaluable, ?fd evaluable)
    #\=(?fd evaluable, ?fd evaluable)
    #<(?fd evaluable, ?fd evaluable)
    #=<(?fd evaluable, ?fd evaluable)
    #>(?fd evaluable, ?fd evaluable)
    #>=(?fd evaluable, ?fd evaluable)

Description

FdExpr1 #= FdExpr2 constrains FdExpr1 to be equal to FdExpr2.
FdExpr1 #\= FdExpr2 constrains FdExpr1 to be different from FdExpr2.
FdExpr1 #< FdExpr2 constrains FdExpr1 to be less than FdExpr2.

FdExpr1 #=< FdExpr2 constrains FdExpr1 to be less than or equal to FdExpr2.

FdExpr1 #> FdExpr2 constrains FdExpr1 to be greater than FdExpr2.

FdExpr1 #>= FdExpr2 constrains FdExpr1 to be greater than or equal to FdExpr2. FdExpr1 and FdExpr2 are arithmetic FD expressions (section 9.6.1, page 188). #=, #\=, #<, #=<, #> and #>= are predefined infix operators (section 8.14.10, page 113). These predicates post arithmetic constraints that are managed by the solver using a partial arc-consistency algorithm to reduce the domain of involved variables. In this scheme only the bounds of the domain of variables are updated. This leads to less propagation than full arc-consistency techniques (section 9.6.3, page 190) but is generally more efficient for arithmetic. These arithmetic constraints can be reified (section 9.7, page 191). Errors Refer to the syntax of arithmetic FD expressions for possible errors (section 9.6.1, page 188). Portability GNU Prolog predicates.

9.6.3 Full AC: (#=#)/2 - constraint equal, (#\=#)/2 - constraint not equal, (#<#)/2 - constraint less than, (#=<#)/2 - constraint less than or equal, (#>#)/2 - constraint greater than, (#>=#)/2 - constraint greater than or equal

Templates

     #=#(?fd evaluable, ?fd evaluable)
     #\=#(?fd evaluable, ?fd evaluable)
     #<#(?fd evaluable, ?fd evaluable)
     #=<#(?fd evaluable, ?fd evaluable)
     #>#(?fd evaluable, ?fd evaluable)
     #>=#(?fd evaluable, ?fd evaluable)

Description

FdExpr1 #=# FdExpr2 constrains FdExpr1 to be equal to FdExpr2.
FdExpr1 #\=# FdExpr2 constrains FdExpr1 to be different from FdExpr2.
FdExpr1 #<# FdExpr2 constrains FdExpr1 to be less than FdExpr2.
FdExpr1 #=<# FdExpr2 constrains FdExpr1 to be less than or equal to FdExpr2.
FdExpr1 #># FdExpr2 constrains FdExpr1 to be greater than FdExpr2.

FdExpr1 #>=# FdExpr2 constrains FdExpr1 to be greater than or equal to FdExpr2. FdExpr1 and FdExpr2 are arithmetic FD expressions (section 9.6.1, page 188). #=#, #\=#, #<#, #=<#, #># and #>=# are predefined infix operators (section 8.14.10, page 113).

These predicates post arithmetic constraints that are managed by the solver using a full arc-consistency algorithm to reduce the domain of involved variables. In this scheme the full domain of variables is updated. This leads to more propagation than partial arc-consistency techniques (section 9.6.1, page 188) but is generally less efficient for arithmetic. These arithmetic constraints can be reified (section 9.7.1, page 191).

Errors

Refer to the syntax of arithmetic FD expressions for possible errors (section 9.6.1, page 188).

Portability

GNU Prolog predicates.

9.6.4 fd prime/1, fd not prime/1

Templates

     fd prime(?fd variable)
     fd not prime(?fd variable)

Description

fd prime(X) constraints X to be a prime number between 0..vector max. This constraint enforces a sparse representation for the domain of X (section 9.1, page 183).

fd not prime(X) constraints X to be a non prime number between 0..vector max. This constraint enforces a sparse representation for the domain of X (section 9.1, page 183).

Errors

X is neither an FD variable nor an integer type error(fd variable, X)

Portability

GNU Prolog predicates.

9.7 Boolean and reified constraints

9.7.1 Boolean FD expressions

An boolean FD expression is a Prolog term built from integers (0 for false, 1 for true), variables (Prolog or FD variables), partial AC arithmetic constraints (section 9.6.2, page 189), full AC arithmetic constraints (section 9.6.3, page 190) and functors (or operators) that represent boolean functions. When a subexpression of a boolean expression is an arithmetic constraint c , it is reified. Namely, as soon as the solver detects that c is true (i.e. entailed) the sub-expression has the value 1. Similarly when the solver detects that c is false (i.e. disentailed) the sub-expression evaluates as 0. While neither the entailment nor the disentailment can be detected the sub-expression is evaluated as a domain 0..1. The following table details the components of an FD boolean expression:

FD Expression Result
Prolog variable domain 0..1
FD variable X domain of X, X is constrained to be in 0..1
0 (integer)  0 (false)
1 (integer)  1 (true)

#\ E  not E
E1 #<=> E2     E1 equivalent to E2

E1 #\<=> E2    E1 not equivalent to E2 (i.e. E1 different from E2)
E1 ## E2       E1 exclusive OR E2 (i.e. E1 not equivalent to E2)

E1 #==> E2     E1 implies E2
E1 #\==> E2    E1 does not imply E2
E1 #/\ E2      E1 AND E2
E1 #\/\ E2     E1 NAND E2
E1 #\/ E2      E1 OR E2
E1 #\\/ E2     E1 NOR E2

#<=>, #\<=>, ##, #==>, #\==>, #/\, #\/\, #\/ and #\\/ are predefined infix operators. #\ is a predefined prefix operator (section 8.14.10, page 113).

Errors

a sub-expression E is neither a variable nor an integer (0 or 1) nor an FD boolean functor nor reified constraint

                                           type error(fd bool evaluable, E)
an expression is too complex  resource error(too big fd constraint)

a sub-expression is an invalid reified constraint an arithmetic constraint error (section 9.6.1, page 188)

9.7.2 fd reified in/4

Templates

    fd reified in(?fd variable, +integer, +integer, ?fd variable)

Description

fd reified in(X, Lower, Upper, B) captures the truth value of the constraint X ∈ [Lower..Upper] in the boolean variable B.

Errors

X is neither a variable nor an FD variable nor an integer

                                           type error(fd variable, X)

B is neither a variable nor an FD variable nor an integer

                                           type error(fd variable, B)
Lower is a variable  instantiation error

Lower is neither a variable nor an integer type error(integer, Lower)

Upper is a variable  instantiation error

Upper is neither a variable nor an integer type error(integer, Upper)

9.7.3 (#\)/1 - constraint NOT, (#<=>)/2 - constraint equivalent, (#\<=>)/2 - constraint different, (##)/2 - constraint XOR, (#==>)/2 - constraint imply, (#\==>)/2 - constraint not imply, (#/\)/2 - constraint AND, (#\/\)/2 - constraint NAND, (#\/)/2 - constraint OR, (#\\/)/2 - constraint NOR

Templates

     #\(?fd bool evaluable)
     #<=>(?fd bool evaluable, ?fd bool evaluable)
     #\<=>(?fd bool evaluable, ?fd bool evaluable)
     ##(?fd bool evaluable, ?fd bool evaluable)
     #==>(?fd bool evaluable, ?fd bool evaluable)
     #\==>(?fd bool evaluable, ?fd bool evaluable)
     #/\(?fd bool evaluable, ?fd bool evaluable)
     #\/\(?fd bool evaluable, ?fd bool evaluable)
     #\/(?fd bool evaluable, ?fd bool evaluable)
     #\\/(?fd bool evaluable, ?fd bool evaluable)

Description

#\ FdBoolExpr1 constraints FdBoolExpr1 to be false.
FdBoolExpr1 #<=> FdBoolExpr2 constrains FdBoolExpr1 to be equivalent to FdBoolExpr2.
FdBoolExpr1 #\<=> FdBoolExpr2 constrains FdBoolExpr1 to be equivalent to not FdBoolExpr2.
FdBoolExpr1 ## FdBoolExpr2 constrains FdBoolExpr1 XOR FdBoolExpr2 to be true
FdBoolExpr1 #==> FdBoolExpr2 constrains FdBoolExpr1 to imply FdBoolExpr2.
FdBoolExpr1 #\==> FdBoolExpr2 constrains FdBoolExpr1 to not imply FdBoolExpr2.
FdBoolExpr1 #/\ FdBoolExpr2 constrains FdBoolExpr1 AND FdBoolExpr2 to be true.
FdBoolExpr1 #\/\ FdBoolExpr2 constrains FdBoolExpr1 AND FdBoolExpr2 to be false.
FdBoolExpr1 #\/ FdBoolExpr2 constrains FdBoolExpr1 OR FdBoolExpr2 to be true.
FdBoolExpr1 #\\/ FdBoolExpr2 constrains FdBoolExpr1 OR FdBoolExpr2 to be false.

FdBoolExpr1 and FdBoolExpr2 are boolean FD expressions (section 9.7.1, page 191).

Note that #\<=> (not equivalent) and ## (exclusive or) are synonymous.

These predicates post boolean constraints that are managed by the FD solver using a partial arcconsistency algorithm to reduce the domain of involved variables. The (dis)entailment of reified constraints is detected using either the bounds (for partial AC arithmetic constraints) or the full domain (for full AC arithmetic constraints).

#<=>, #\<=>, ##, #==>, #\==>, #/\, #\/\, #\/ and #\\/ are predefined infix operators. #\ is a predefined prefix operator (section 8.14.10, page 113).

Refer to the syntax of boolean FD expressions for possible errors (section 9.7.1, page 191).

Portability

GNU Prolog predicates.

9.7.4 fd cardinality/2, fd cardinality/3, fd at least one/1, fd at most one/1,
       fd only one/1

Templates

     fd cardinality(+fd bool evaluable list, ?fd variable)
     fd cardinality(+integer, ?fd variable, +integer)
     fd at least one(+fd bool evaluable list)
     fd at most one(+fd bool evaluable list)
     fd only one(+fd bool evaluable list)

Description

fd cardinality(List, Count) unifies Count with the number of constraints that are true in List. This is equivalent to post the constraint B_{1} + B_{2} + ...+ B_{n} #= Count where each variable Bi is a new variable defined by the constraint B_{i} #<=> C_{i} where C_{i} is the ith constraint of List. Each C_{i} must be a boolean FD expression (section 9.7.1, page 191).

fd cardinality(Lower, List, Upper) is equivalent to fd cardinality(List, Count), Lower #=<
Count, Count #=< Upper
fd at least one(List) is equivalent to fd cardinality(List, Count), Count #>= 1.
fd at most one(List) is equivalent to fd cardinality(List, Count), Count #=< 1.
fd only one(List) is equivalent to fd cardinality(List, 1).

Errors

 List is a partial list  instantiation error

List is neither a partial list nor a list type error(list, List) Count is neither an FD variable nor an integer type error(fd variable, Count)

 Lower is a variable  instantiation error

Lower is neither a variable nor an integer type error(integer, Lower)

 Upper is a variable  instantiation error

Upper is neither a variable nor an integer type error(integer, Upper) an element E of the List list is an invalid boolean expression an FD boolean constraint (section 9.7.1, page 191)

Portability

9.8 Symbolic constraints

9.8.1 fd all different/1

Templates

     fd all different(+fd variable list)

Description

fd all different(List) constrains all variables in List to take distinct values. This is equivalent to posting an inequality constraint for each pair of variables. This constraint is triggered when a variable becomes ground, removing its value from the domain of the other variables.

Errors

 List is a partial list  instantiation error

List is neither a partial list nor a list type error(list, List) an element E of the List list is neither a variable nor an integer nor an FD variable

                                             type error(fd variable, E)

Portability

GNU Prolog predicate.

9.8.2 fd element/3

Templates

     fd element(?fd variable, +integer list, ?fd variable)

Description

fd element(I, List, X) constraints X to be equal to the Ith integer (from 1) of List.

Errors

I is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, I)

X is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, X)

List is a partial list or a list with an element E which is a variable

                                             instantiation error

List is neither a partial list nor a list type error(list, List) an element E of the List list is neither a variable nor an integer

                                             type error(integer, E)

Portability

9.8.3 fd element var/3

Templates

     fd element var(?fd variable, +fd variable list, ?fd variable)

Description

fd element var(I, List, X) constraints X to be equal to the Ith variable (from 1) of List. This constraint is similar to fd element/3 (section 9.8.2, page 195) but List can also contain FD variables (rather than just integers).

Errors

I is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, I)

X is neither a variable nor an FD variable nor an integer

                                             type error(fd variable, X)
 List is a partial list  instantiation error

List is neither a partial list nor a list type error(list, List) an element E of the List list is neither a variable nor an integer nor an FD variable

                                             type error(fd variable, E)

Portability

GNU Prolog predicate.

9.8.4 fd atmost/3, fd atleast/3, fd exactly/3

Templates

     fd atmost(+integer, +fd variable list, +integer)
     fd atleast(+integer, +fd variable list, +integer)
     fd exactly(+integer, +fd variable list, +integer)

Description

fd atmost(N, List, V) posts the constraint that at most N variables of List are equal to the value V.

fd atleast(N, List, V) posts the constraint that at least N variables of List are equal to the value V.

fd exactly(N, List, V) posts the constraint that at exactly N variables of List are equal to the value

V.

These constraints are special cases of fd cardinality/2 (section 9.7.4, page 194) but their implementation is more efficient.

N is a variable  instantiation error

N is neither a variable nor an integer type error(integer, N)

V is a variable  instantiation error

V is neither a variable nor an integer type error(integer, V)

List is a partial list  instantiation error

List is neither a partial list nor a list type error(list, List) an element E of the List list is neither a variable nor an FD variable nor an integer

                                           type error(fd variable, E)

Portability

GNU Prolog predicates.

9.8.5 fd relation/2, fd relationc/2

Templates

    fd relation(+integer list list, ?fd variable list)
    fd relationc(+integer list list, ?fd variable list)

Description

fd relation(Relation, Vars) constraints the tuple of variables Vars to be equal to one tuple of the list Relation. A tuple is represented by a list.

Example: definition of the boolean AND relation so that X AND Y ⇔ Z:

    and(X,Y,Z):-
            fd_relation([[0,0,0],[0,1,0],[1,0,0],[1,1,1]], [X,Y,Z]).

fd relationc(Columns, Vars) is similar to fd relation/2 except that the relation is not given as the list of tuples but as the list of the columns of the relation. A column is represented by a list.

Example:

    and(X,Y,Z):-
            fd_relationc([[0,0,1,1],[0,1,0,1],[0,0,0,1]], [X,Y,Z]).

Errors

Relation is a partial list or a list with a sub-term E which is a variable

                                           instantiation error
Relation is neither a partial list nor a list  type error(list, Relation)

an element E of the Relation list is neither a variable nor an integer

                                           type error(integer, E)
Vars is a partial list  instantiation error

Vars is neither a partial list nor a list type error(list, Vars) an element E of the Vars list is neither a variable nor an integer nor an FD variable

                                           type error(fd variable, E)

Portability

9.9 Labeling constraints

9.9.1 fd labeling/2, fd labeling/1, fd labelingff/1

Templates

fd labeling(+fd variable list, +fd labeling option list)
fd labeling(+fd variable, +fd labeling option list)
fd labeling(+fd variable list)
fd labeling(+fd variable)
fd labelingff(+fd variable list)
fd labelingff(+fd variable)

Description

fd labeling(Vars, Options) assigns a value to each variable X of the list Vars according to the list of labeling options given by Options. Vars can be also a single FD variable. This predicate is re-executable on backtracking.

FD labeling options: Options is a list of labeling options. If this list contains contradictory options, the rightmost option is the one which applies. Possible options are:

• variable method(V): specifies the heuristics to select the variable to enumerate:

– standard: no heuristics, the leftmost variable is selected.

– first fail (or ff): selects the variable with the smallest number of elements in its domain. If several variables have the same number of elements the leftmost variable is selected.

– most constrained: like first fail but when several variables have the same number of elements selects the variable that appears in most constraints.

– smallest: selects the variable that has the smallest value in its domain. If there is more than one such variable selects the variable that appears in most constraints.

– largest: selects the variable that has the greatest value in its domain. If there is more than one such variable selects the variable that appears in most constraints.

– max regret: selects the variable that has the greatest difference between the smallest value and the next value of its domain. If there is more than one such variable selects the variable that appears in most constraints.

– random: selects randomly a variable. Each variable is chosen only once.

The default value is standard.

• reorder(true/false): specifies if the variable heuristics should dynamically reorder the list of variable (true) or not (false). Dynamic reordering is generally more efficient but in some cases a static ordering is faster. The default value is true.

• value method(V): specifies the heuristics to select the value to assign to the chosen variable: – min: enumerates the values from the smallest to the greatest (default).

– max: enumerates the values from the greatest to the smallest.

– middle: enumerates the values from the middle to the bounds.

– bounds: enumerates the values from the bounds to the middle.

– random: enumerates the values randomly. Each value is tried only once.

– bisect: recursively creates a choice between X #=< M and X #> M, where M is the midpoint of the domain of the variable. Values are thus tried from the smallest to the greatest. This is also known as domain splitting.

The default value is min.

• backtracks(B): unifies B with the number of backtracks during the enumeration.

fd labeling(Vars) is equivalent to fd labeling(Vars, []).
fd labelingff(Vars) is equivalent to fd labeling(Vars, [variable method(ff)]).

Errors

Vars is a partial list or a list with an element E which is a variable

                                             instantiation error

Vars is neither a partial list nor a list type error(list, Vars) an element E of the Vars list is neither a variable nor an integer nor an FD variable

                                             type error(fd variable, E)

Options is a partial list or a list with an element E which is a variable

                                             instantiation error
 Options is neither a partial list nor a list  type error(list, Options)

an element E of the Options list is neither a variable nor a labeling option

                                             domain error(fd labeling option, E)

Portability

GNU Prolog predicates.

9.10 Optimization constraints

9.10.1 fd minimize/2, fd maximize/2

Templates

     fd minimize(+callable term, ?fd variable)
     fd maximize(+callable term, ?fd variable)

Description

fd minimize(Goal, X) repeatedly calls Goal to find a value that minimizes the variable X. Goal is a Prolog goal that should instantiate X, a common case being the use of fd labeling/2 (section 9.9.1, page 198). This predicate uses a branch-and-bound algorithm with restart: each time call(Goal) succeeds the computation restarts with a new constraint X #< V where V is the value of X at the end of the last call of Goal. When a failure occurs (either because there are no remaining choice-points for Goal or because the added constraint is inconsistent with the rest of the store) the last solution is recomputed since it is optimal.

fd maximize(Goal, X) is similar to fd minimize/2 but X is maximized.

Goal is a variable  instantiation error

Goal is neither a variable nor a callable term type error(callable, Goal) The predicate indicator Pred of Goal does not correspond to an existing procedure and the value of the unknown Prolog flag is error (section 8.22.1, page 148)

                                           existence error(procedure, Pred)

X is neither a variable nor an FD variable nor an integer

                                           type error(fd variable, X)

Portability

10 Interfacing Prolog and C

10.1 Introduction

The foreign code interface allows the use to link Prolog and C in both directions.

A Prolog predicate can call a C function passing different kinds of arguments (input, output or input/output). The interface performs implicit Prolog ↔ C data conversions for simple types (for instance a Prolog integer is automatically converted into a C integer) and provides a set of API (Application Programming Interface) functions to convert more complex types (lists or structures). The interface also performs automatic error detection depending on the actual type of the passed argument. An important feature is the ability to write non-deterministic code in C.

It is also possible to call (or callback) a Prolog predicate from a C function and to manage Prolog nondeterminism: the C code can ask for next solutions, remove all remaining solutions or terminate and keep alternatives for the calling Prolog predicate).

10.2 Including and using gprolog.h

The C code should include gprolog.h which provides a set of C definitions (types, macros, prototypes) associated to the API. Include this files as follows:

     #include <gprolog.h>

If the installation has been correctly done nothing else is needed. If the C compiler/preprocessor cannot locate gprolog.h pass the C compiler option required to specify an additional include directory (e.g.-Iinclude dir) to gplc as follows (section 4.4.3, page 23):

% gplc -C -Iinclude dir ...

The file gprolog.h declares the following C types:

• PlBool as an integer and the constants PL FALSE (i.e. 0) and PL TRUE (i.e. 1).

• PlLong as an integer able to store a pointer (equivalent to intptr t). This type appeared in GNU Prolog 1.4.0 in replacement of long to support Windows 64 bits (where the long type is only 32 bits). This type is used to handle integer types.

• PlULong same as PlLong but unsigned (same as uintptr t).

• PlTerm same as intptr t. This type is used to store general Prolog terms.

New in GNU Prolog 1.3.1 and backward compatibility issues: in GNU Prolog 1.3.1 the API has been modified to protect namespace. The name of public functions, macros, variables and types are now prefixed with Pl , PL or pl . All these prefixes should be avoided by the foreign C-code to prevent name clashes. To ensure a backward compatibility, the names used by the old API are available thanks to a set of #define. However, this deprecated API should not be used by recent code. It is also possible to prevent the definition of the compatibility macros using:

     #define __GPROLOG_FOREIGN_STRICT__
     #include <gprolog.h>

In addition, gprolog.h defines a set of macros:

• GNU PROLOG (as the major version).

   • GPROLOG , GPROLOG MINOR and GPROLOG PATCHLEVEL . Their values are the major version,

minor version, and patch level of GNU Prolog, as integer constants. For example, GNU Prolog 1.3.2

     will define GPROLOG to 1, GPROLOG MINOR to 3, and GPROLOG PATCHLEVEL to 2.

If you need to write code which depends on a specific version, you must be more careful. Recall these macros appeared in GNU Prolog 1.3.1 (undefined before), each time the minor version is increased, the patch level is reset to zero; each time the major version is increased (which happens rarely), the minor version and patch level are reset.

• GPROLOG VERSION : the version as an integer defined as follows: major ∗ 10000 + minor ∗ 100 + patchlevel. For example: version 1.3.2 will result in the value 10302.

• PL PROLOG DIALECT: a C constant string (generally "gprolog"). Appeared in 1.3.2.

• PL PROLOG NAME: a C constant string (generally "GNU Prolog").

• PL PROLOG VERSION: a C constant string associated to the version (e.g. "1.4.0").

• PL PROLOG DATE: a C constant string associated with the date of this version (e.g. "Mar 29 2011".

• PL PROLOG COPYRIGHT: a C constant string associated with the copyright of this version (e.g.

     "Copyright (C) 1999-2021 Daniel Diaz".

Note the above PL PROLOG ... macros are also accessible via Prolog flags thanks to the built-in predicate current prolog flag/2 (section 8.22.2, page 151)

10.3 Calling C from Prolog

10.3.1 Introduction

This interface can then be used to write both simple and complex C routines. A simple routine uses either input or output arguments which type is simple. In that case the user does not need any knowledge of Prolog data structures since all Prolog ↔ C data conversions are implicitly achieved. To manipulate complex terms (lists, structures) a set of functions is provided. Finally it is also possible to write nondeterministic C code.

10.3.2 foreign/2 directive

foreign/2 directive (section 7.1.15, page 50) declares a C function interface. The general form is foreign(Template, Options) which defines an interface predicate whose prototype is Template according to the options given by Options. Template is a callable term specifying the type/mode of each argument of the associated Prolog predicate.

Foreign options: Options is a list of foreign options. If this list contains contradictory options, the rightmost option is the one which applies. Possible options are:

• fct name(F): F is an atom representing the name of the C function to call. By default the name of the C function is the same as the principal functor of Template. In any case, the atom associated with the name of the function must conforms to the syntax of C identifiers.

• return(boolean/none/jump): specifies the value returned by the C function:

– boolean: the type of the function is PlBool (returns PL TRUE on success, PL FALSE otherwise).

– none: the type of the function is void (no returned value).

– jump: the type of the function is void(*)() (returns the address of a Prolog code to execute).

The default value is boolean.

• bip name(Name, Arity): initializes the error context with Name and Arity. If an error occurs this information is used to indicate from which predicate the error occurred (section 6.3.1, page 39). It is also possible to prevent the initialization of the error context using bip name(none). By default Name and Arity are set to the functor and arity of Template.

• choice size(N): this option specifies that the function implements a non-deterministic code. N is an integer specifying the size needed by the non-deterministic C function. This facility is explained later (section 10.3.7, page 205). By default a foreign function is deterministic.

foreign(Template) is equivalent to foreign(Template, []).

Foreign modes and types: each argument of Template specifies the foreign mode and type of the corresponding argument. This information is used to check the type of effective arguments at run-time and to perform Prolog ↔ C data conversions. Each argument of Template is formed with a mode symbol followed by a type name. Possible foreign modes are:

• +: input argument.

• -: output argument.

• ?: input/output argument.

Possible foreign types are:

Foreign type Prolog type  C type Description of the C type
integer  integer  PlLong value of the integer
positive  positive integer  PlLong value of the integer
float  floating point number double value of the floating point number
number  number  double value of the number
atom  atom  PlLong internal key of the atom
boolean  boolean  PlLong value of the boolean (0=false, 1=true)
char  character  PlLong value of (the code of) the character
code  character code  PlLong value of the character-code
byte  byte  PlLong value of the byte
in char  in-character  PlLong value of the character or -1 for end-of-file
in code  in-character code  PlLong value of the character-code or -1 for end-of-file
in byte  in-byte  PlLong value of the byte or -1 for the end-of-file
string  atom  char * C string containing the name of the atom
chars  character list  char * C string containing the characters of the list
codes  character-code list  char * C string containing the characters of the list
term  Prolog term  PlTerm generic Prolog term

Simple foreign type: a simple type is any foreign type listed in the above tabled except term. A simple foreign type is an atomic term (character and character-code lists are in fact lists of constants). Each simple foreign type is converted to/from a C type to simplify the writing of the C function.

Complex foreign type: type foreign type term refers to any Prolog term (e.g. lists, structures…). When such an type is specified the argument is passed to the C function as a PlTerm (GNU Prolog C type equivalent to a PlLong). Several functions are provided to manipulate PlTerm variables (section 10.4, page 210). Since the original term is passed to the function it is possible to read its value or to unify it. So the meaning of the mode symbol is less significant. For this reason it is possible to omit the mode symbol. In that case term is equivalent to +term.

10.3.3 The C function

The type returned by the C function depends on the value of the return foreign option (section 10.3.2, page 202). If it is boolean then the C function is of type PlBool and shall return PL TRUE in case of success and PL FALSE otherwise. If the return option is none the C function is of type void. Finally if it is jump, the function shall return the address of a Prolog predicate and, at the exit of the function, the control is given to that predicate.

The type of the arguments of the C function depends on the mode and type declaration specified in Template for the corresponding argument as explained in the following sections.

10.3.4 Input arguments

An input argument is tested at run-time to check if its type conforms to the foreign type and then it is passed to the C function. The type of the associated C argument is given by the above table (section 10.3.2, page 202). For instance, the effective argument Arg associated with +positive foreign declaration is submitted to the following process:

• if Arg is a variable an instantiation error is raised.

• if Arg is neither a variable nor an integer a type error(integer, Arg) is raised.

   • if Arg is an integer < 0 a domain error(not less than zero, Arg) is raised.

• otherwise the value of Arg is passed to the C is passed to the C function as an integer (PlLong).

When +string is specified the string passed to the function is the internal string of the corresponding atom and should not be modified.

When +term is specified the term passed to the function is the original Prolog term. It can be read and/or unified. It is also the case when term is specified without any mode symbol.

10.3.5 Output arguments

An output argument is tested at run-time to check if its type conforms to the foreign type and it is unified with the value set by the C function. The type of the associated C argument is a pointer to the type given by the above table (section 10.3.2, page 202). For instance, the effective argument Arg associated with -positive foreign declaration is handled as follows:

• if Arg is neither a variable nor an integer a type error(integer, Arg) is raised.

   • if Arg is an integer < 0 a domain error(not less than zero, Arg) is raised.

• otherwise a pointer to an integer (PlLong *) is passed to the C function. If the function returns PL TRUE the integer stored at this location is unified with Arg.

When -term is specified, the function must construct a term into the its corresponding argument (which is of type PlTerm *). At the exit of the function this term will be unified with the actual predicate argument.

10.3.6 Input/output arguments

Basically an input/output argument is treated as in input argument if it is not a variable, as an output argument otherwise. The type of the associated C argument is a pointer to a PlFIOArg (GNU Prolog C type) defined as follows:

     typedef struct
         {
          PlBool is_var;
          PlBool unify;
          union
             {
              PlLong l;
              char  *s;
              double d;
             }value;
         }PlFIOArg;

The field is var is set to PL TRUE if the argument is a variable and PL FALSE otherwise. This value can be tested by the C function to determine which treatment to perform. The field unify controls whether the effective argument must be unified at the exit of the C function. Initially unify is set to the same value as is var (i.e. a variable argument will be unified while a non-variable argument will not) but it can be modified by the C function. The field value stores the value of the argument. It is declared as a C union since there are several kinds of value types. The field s is used for C strings, d for C doubles and l otherwise (int, PlLong, PlTerm). if is var is PL FALSE then value contains the input value of the argument with the same conventions as for input arguments (section 10.3.4, page 204). At the exit of the function, if unify is PL TRUE value must contain the value to unify with the same conventions as for output arguments (section 10.3.5, page 204).

For instance, the effective argument Arg associated with ?positive foreign declaration is handled as follows:

• if Arg is a variable is var and unify are set to PL TRUE else to PL FALSE and its value is copied in

     value.l.

• if Arg is neither a variable nor an integer a type error(integer, Arg) is raised.

   • if Arg is an integer < 0 a domain error(not less than zero, Arg) is raised.

• otherwise a pointer to the PlFIOArg (PlFIOArg *) is passed to the C function. If the function returns PL TRUE and if unify is PL TRUE the value stored in value.l is unified with Arg.

10.3.7 Writing non-deterministic C code

The interface allows the user to write non-deterministic C code. When a C function is non-deterministic, a choice-point is created for this function. When a failure occurs, if all more recent non-deterministic code are finished, the function is re-invoked. It is then important to inform Prolog when there is no more solution (i.e. no more choice) for a non-deterministic code. So, when no more choices remains the function must remove the choice-point. The interface increments a counter each time the function is re-invoked. At the first call this counter is equal to 0. This information allows the function to detect its first call. When writing non-deterministic code, it is often useful to record data between consecutive re-invocations of the function. The interface maintains a buffer to record such an information. The size of this buffer is given by choice size(N) when using foreign/2 (section 10.3.2, page 202). This size is the number of (consecutive) PlLongs needed by the C function. Inside the function it is possible to call the following functions/macros:

int  Pl_Get_Choice_Counter(void)
TYPE Pl_Get_Choice_Buffer (TYPE)
void Pl_No_More_Choice    (void)

The macro Pl Get Choice Counter() returns the value of the invocation counter (0 at the first call).

The macro Pl Get Choice Buffer(TYPE) returns a pointer to the buffer (casted to TYPE ).

The function Pl No More Choice() deletes the choice point associated with the function.

10.3.8 Example: input and output arguments

All examples presented here can be found in the ExamplesC sub-directory of the distribution, in the files examp.pl (Prolog part) and examp c.c (C part).

Let us define a predicate first occurrence(A, C, P) which unifies P with the position (from 0) of the first occurrence of the character C in the atom A. The predicate must fail if C does not appear in A.

In the prolog file examp.pl:

:- foreign(first occurrence(+string, +char, -positive)).

In the C file examp c.c:

#include <string.h>
#include <gprolog.h>
PlBool
first_occurrence(char *str, PlLong c, PlLong *pos)
{
  char *p;
  p = strchr(str, c);
  if (p == NULL)                    /* C does not appear in A */
    return PL_FALSE;                /* fail */
  *pos = p - str;                   /* set the output argument */
  return PL_TRUE;                   /* succeed */
}

The compilation produces an executable called examp:

% gplc examp.pl examp c.c

Examples of use:

| ?- first_occurrence(prolog, p, X).
X = 0
| ?- first_occurrence(prolog, k, X).
no
| ?- first_occurrence(prolog, A, X).
{exception: error(instantiation_error,first_occurrence/3)}

     | ?- first_occurrence(prolog, 1 ,X).
     {exception: error(type_error(character,1),first_occurrence/3)}

10.3.9 Example: non-deterministic code

We here define a predicate occurrence(A, C, P) which unifies P with the position (from 0) of one occurrence of the character C in the atom A. The predicate will fail if C does not appear in A. The predicate is re-executable on backtracking. The information that must be recorded between two invocations of the function is the next starting position in A to search for C.

In the prolog file examp.pl:

     :- foreign(occurrence(+string, +char, -positive), [choice size(1)]).

In the C file examp c.c:

     #include <string.h>
     #include <gprolog.h>
     PlBool
     occurrence(char *str, PlLong c, PlLong *pos)
     {
       char **info_pos;
       char *p;
       info_pos = Pl_Get_Choice_Buffer(char **); /* recover the buffer */
       if (Pl_Get_Choice_Counter() == 0) /* first invocation ? */
         *info_pos = str;
       p = strchr(*info_pos, c);
       if (p == NULL)                    /* c does not appear */
         {
           Pl_No_More_Choice();          /* remove choice-point */
           return PL_FALSE;              /* fail */
         }
       *pos = p - str;                   /* set the output argument */
       *info_pos = p + 1;                /* update next starting pos */
       return PL_TRUE;                   /* succeed */
     }

The compilation produces an executable called examp:

     % gplc examp.pl examp c.c

Examples of use:

| ?- occurrence(prolog, o, X).

X = 2 ?  (here the user presses ; to compute another solution)
X = 4 ?  (here the user presses ; to compute another solution)
no  (no more solution)

| ?- occurrence(prolog, k, X).
no

In the first example when the second (the last) occurrence is found (X=4) the choice-point remains and the failure is detected only when another solution is requested (by pressing ;). It is possible to improve this behavior by deleting the choice-point when there is no more occurrence. To do this it is necessary to do one search ahead. The information stored is the position of the next occurrence. Let us define such a behavior for the predicate occurrence2/3.

In the prolog file examp.pl:

:- foreign(occurrence2(+string, +char, -positive), [choice size(1)]).

In the C file examp c.c:

#include <string.h>
#include <gprolog.h>
PlBool
occurrence2(char *str, PlLong c, PlLong *pos)
{
  char **info_pos;
  char *p;
  info_pos = Pl_Get_Choice_Buffer(char **); /* recover the buffer */
  if (Pl_Get_Choice_Counter() == 0) /* first invocation ? */
    {
      p = strchr(str, c);
      if (p == NULL)                /* C does not appear at all */
        {
          Pl_No_More_Choice();      /* remove choice-point */
          return PL_FALSE;          /* fail */
        }
      *info_pos = p;
    }
                                    /* info_pos = an occurrence */
  *pos = *info_pos - str;           /* set the output argument */
  p = strchr(*info_pos + 1, c);
  if (p == NULL)                    /* no more occurrence */
    Pl_No_More_Choice();            /* remove choice-point */
  else
    *info_pos = p;                  /* else update next solution */
  return PL_TRUE;                   /* succeed */
}

Examples of use:

| ?- occurrence2(prolog, l, X).

X = 3  (here the user is not prompted since there is no more alternative)

| ?- occurrence2(prolog, o, X).

X = 2 ?  (here the user presses ; to compute another solution)
X = 4  (here the user is not prompted since there is no more alternative)

10.3.10 Example: input/output arguments

We here define a predicate char ascii(Char, Code) which converts in both directions the character Char and its character-code Code. This predicate is then similar to char code/2 (section 8.19.4, page 130).

In the prolog file examp.pl:

:- foreign(char ascii(?char, ?code)).

In the C file examp c.c:

#include <gprolog.h>
PlBool
char_ascii(PlFIOArg *c, PlFIOArg *ascii)
{
  if (!c->is_var)                  /* Char is not a variable */
    {
      ascii->unify = PL_TRUE;      /* enforce unif. of Code */
      ascii->value.l = c->value.l; /* set Code */
      return PL_TRUE;              /* succeed */
    }
  if (ascii->is_var)               /* Code is also a variable */
    Pl_Err_Instantiation();        /* emit instantiation_error */
  c->value.l = ascii->value.l;     /* set Char */
  return PL_TRUE;                  /* succeed */
}

If Char is instantiated it is necessary to enforce the unification of Code since it could be instantiated. Recall that by default if an input/output argument is instantiated it will not be unified at the exit of the function (section 10.3.6, page 205). If both Char and Code are variables the function raises an instantiation error. The way to raise Prolog errors is described later (section 10.5, page 217).

The compilation produces an executable called examp:

% gplc examp.pl examp c.c

Examples of use:

| ?- char_ascii(a, X).
X = 97

     | ?- char_ascii(X, 65).
     X = 'A'
     | ?- char_ascii(a, 12).
     no
     | ?- char_ascii(X, X).
     {exception: error(instantiation_error,char_ascii/2)}
     | ?- char_ascii(1, 12).
     {exception: error(type_error(character,1),char_ascii/2)}

10.4 Manipulating Prolog terms

10.4.1 Introduction

In the following we presents a set of functions to manipulate Prolog terms. For simple foreign terms the functions manipulate simple C types (section 10.3.2, page 202).

Functions managing lists handle an array of 2 elements (of type PlTerm) containing the terms corresponding to the head and the tail of the list. For the empty list NULL is passed as the array. These functions require to flatten a list in each sub-list. To simplify the management of proper lists (i.e. lists terminated by []) a set of functions is provided that handle the number of elements of the list (an integer) and an array whose elements (of type PlTerm) are the elements of the list. The caller of these functions must provide the array.

Functions managing compound terms handle a functor (the principal functor of the term), an arity N ≥ 0 and an array of N elements (of type PlTerm) containing the sub-terms of the compound term. Since a list is a special case of compound term (functor = '.' and arity=2) it is possible to use any function managing compound terms to deal with a list but the error detection is not the same. Indeed many functions check if the Prolog argument is correct. The name of a read or unify function checking the Prolog arguments is of the form Name Check(). For each of these functions there is a also check-free version called Name(). We then only present the name of checking functions.

10.4.2 Managing Prolog atoms

Each atom has a unique internal key (an integer) which corresponds to its index in the GNU Prolog atom table. It is possible to obtain the information about an atom and to create new atoms using:

     char   *Pl_Atom_Name           (int atom)
     int     Pl_Atom_Length         (int atom)
     PlBool  Pl_Atom_Needs_Quote    (int atom)
     PlBool  Pl_Atom_Needs_Scan     (int atom)
     PlBool  Pl_Is_Valid_Atom       (int atom)
     int     Pl_Create_Atom         (const char *str)
     int     Pl_Create_Allocate_Atom(const char *str)
     int     Pl_Find_Atom           (const char *str)
     int     Pl_Atom_Char           (char c)

     int     Pl_Atom_Nil            (void)
     int     Pl_Atom_False          (void)
     int     Pl_Atom_True           (void)
     int     Pl_Atom_End_Of_File    (void)

The function Pl Atom Name(atom) returns the internal string of atom (this string should not be modified). The function Pl Atom Length(atom) returns the length (of the name) of atom.

The function Pl Atom Needs Scan(atom) indicates if the canonical form of atom needs to be quoted as done by writeq/2 (section 8.14.6, page 108). In that case Pl Atom Needs Scan(atom) indicates if this simply comes down to write quotes around the name of atom or if it necessary to scan each character of the name because there are some non-printable characters (or included quote characters). The function Pl Is Valid Atom(atom) is true only if atom is the internal key of an existing atom.

The function Pl Create Atom(str) adds a new atom whose name is the content of str to the system and returns its internal key. If the atom already exists its key is simply returned. The string str passed to the function should not be modified later. The function Pl Create Allocate Atom(str) is provided when this condition cannot be ensured. It simply makes a dynamic copy of str (using strdup(3)).

The function Pl Find Atom(str) returns the internal key of the atom whose name is str or -1 if it does not exist.

All atoms corresponding to a single character already exist and their key can be obtained via the function Pl Atom Char. For instance Pl Atom Char('.') is the atom associated with '.' (this atom is the functor of lists). The other functions return the internal key of frequently used atoms: [], false, true

and end of file.

10.4.3 Reading Prolog terms

The name of all functions presented here are of the form Pl Rd Name Check(). They all check the validity of the Prolog term to read emitting appropriate errors if necessary. Each function has a check-free version

called Pl Rd Name().

Simple foreign types: for each simple foreign type (section 10.3.2, page 202) there is a read function (used by the interface when an input argument is provided):

     PlLong  Pl_Rd_Integer_Check       (PlTerm term)
     PlLong  Pl_Rd_Positive_Check      (PlTerm term)
     int     Pl_Rd_C_Int_Check         (PlTerm term)
     int     Pl_Rd_C_Int_Positive_Check(PlTerm term)
     double  Pl_Rd_Float_Check         (PlTerm term)
     double  Pl_Rd_Number_Check        (PlTerm term)
     int     Pl_Rd_Atom_Check          (PlTerm term)
     int     Pl_Rd_Boolean_Check       (PlTerm term)
     int     Pl_Rd_Char_Check          (PlTerm term)
     int     Pl_Rd_In_Char_Check       (PlTerm term)
     int     Pl_Rd_Code_Check          (PlTerm term)
     int     Pl_Rd_In_Code_Check       (PlTerm term)
     int     Pl_Rd_Byte_Check          (PlTerm term)
     int     Pl_Rd_In_Byte_Check       (PlTerm term)
     char   *Pl_Rd_String_Check        (PlTerm term)
     char   *Pl_Rd_Chars_Check         (PlTerm term)
     char   *Pl_Rd_Codes_Check         (PlTerm term)
     int     Pl_Rd_Chars_Str_Check     (PlTerm term, char *str)

     int     Pl_Rd_Codes_Str_Check     (PlTerm term, char *str)

The function Pl Rd C Int Check() is similar to Pl Rd Integer Check() but returns a C int instead of a PlLong. If the Prolog integer does not fit into a C int, a representation error is raised (this can occur on 64-bits machine with int on 32-bits). Similarly for Pl Rd C Int Positive Check().

All functions returning a C string (char *) use a same buffer. The function Pl Rd Chars Str Check() is similar to Pl Rd Chars Check() but accepts as argument a string to store the result and returns the length of that string (which is also the length of the Prolog list). Similarly for Pl Rd Codes Str Check().

Complex terms: the following functions return the sub-arguments (terms) of complex terms as an array of PlTerm except Pl Rd Proper List Check() which returns the size of the list read (and initializes the array element). Refer to the introduction of this section for more information about the arguments of complex functions (section 10.4.1, page 210).

     int     Pl_Rd_Proper_List_Check(PlTerm term, PlTerm *arg)
     PlTerm *Pl_Rd_List_Check       (PlTerm term)
     PlTerm *Pl_Rd_Compound_Check   (PlTerm term, int *functor, int *arity)
     PlTerm *Pl_Rd_Callable_Check   (PlTerm term, int *functor, int *arity)

10.4.4 Unifying Prolog terms

The name of all functions presented here are of the form Pl Un Name Check(). They all check the validity of the Prolog term to unify emitting appropriate errors if necessary. Each function has a check-free version

called Pl Un Name().

Simple foreign types: for each simple foreign type (section 10.3.2, page 202) there is an unify function (used by the interface when an output argument is provided):

     PlBool Pl_Un_Integer_Check (PlLong n,        PlTerm term)
     PlBool Pl_Un_Positive_Check(PlLong n,        PlTerm term)
     PlBool Pl_Un_Float_Check   (double n,        PlTerm term)
     PlBool Pl_Un_Number_Check  (double n,        PlTerm term)
     PlBool Pl_Un_Atom_Check    (int atom,        PlTerm term)
     PlBool Pl_Un_Boolean_Check (int b,           PlTerm term)
     PlBool Pl_Un_Char_Check    (int c,           PlTerm term)
     PlBool Pl_Un_In_Char_Check (int c,           PlTerm term)
     PlBool Pl_Un_Code_Check    (int c,           PlTerm term)
     PlBool Pl_Un_In_Code_Check (int c,           PlTerm term)
     PlBool Pl_Un_Byte_Check    (int b,           PlTerm term)
     PlBool Pl_Un_In_Byte_Check (int b,           PlTerm term)
     PlBool Pl_Un_String_Check  (const char *str, PlTerm term)
     PlBool Pl_Un_Chars_Check   (const char *str, PlTerm term)
     PlBool Pl_Un_Codes_Check   (const char *str, PlTerm term)

The function Pl Un Number Check(n, term) unifies term with an integer if n is an integer, with a floating point number otherwise. The function Pl Un String Check(str, term) creates the atom correspond-

ing to str and then unifies term with it (same as Pl Un Atom Check(Pl Create Allocate Atom(str),
term)).

The following functions perform a general unification (between 2 terms). The second one performs a occurs-check test (while the first one does not).

     PlBool Pl_Unif(PlTerm term1, PlTerm term2)
     PlBool Pl_Unif_With_Occurs_Check(PlTerm term1, PlTerm term2)

Complex terms: the following functions accept the sub-arguments (terms) of complex terms as an array of PlTerm. Refer to the introduction of this section for more information about the arguments of complex functions (section 10.4.1, page 210).

PlBool Pl_Un_Proper_List_Check(int size, PlTerm *arg, PlTerm term)
PlBool Pl_Un_List_Check       (PlTerm *arg, PlTerm term)
PlBool Pl_Un_Compound_Check   (int functor, int arity, PlTerm *arg,
                               PlTerm term)
PlBool Pl_Un_Callable_Check   (int functor, int arity, PlTerm *arg,
                               PlTerm term)

All these functions check the type of the term to unify and return the result of the unification. Generally if an unification fails the C function returns PL FALSE to enforce a failure. However if there are several arguments to unify and if an unification fails then the C function returns PL FALSE and the type of other arguments has not been checked. Normally all error cases are tested before doing any work to be sure that the predicate fails/succeeds only if no error condition is satisfied. So a good method is to check the validity of all arguments to unify and later to do the unification (using check-free functions). Obviously if there is only one to unify it is more efficient to use a unify function checking the argument. For the other cases the interface provides a set of functions to check the type of a term.

Simple foreign types: for each simple foreign type (section 10.3.2, page 202) there is check-for-unification function (used by the interface when an output argument is provided):

void Pl_Check_For_Un_Integer (PlTerm term)
void Pl_Check_For_Un_Positive(PlTerm term)
void Pl_Check_For_Un_Float   (PlTerm term)
void Pl_Check_For_Un_Number  (PlTerm term)
void Pl_Check_For_Un_Atom    (PlTerm term)
void Pl_Check_For_Un_Boolean (PlTerm term)
void Pl_Check_For_Un_Char    (PlTerm term)
void Pl_Check_For_Un_In_Char (PlTerm term)
void Pl_Check_For_Un_Code    (PlTerm term)
void Pl_Check_For_Un_In_Code (PlTerm term)
void Pl_Check_For_Un_Byte    (PlTerm term)
void Pl_Check_For_Un_In_Byte (PlTerm term)
void Pl_Check_For_Un_String  (PlTerm term)
void Pl_Check_For_Un_Chars   (PlTerm term)
void Pl_Check_For_Un_Codes   (PlTerm term)

Complex terms: the following functions check the validity of complex terms:

void Pl_Check_For_Un_List    (PlTerm term)
void Pl_Check_For_Un_Compound(PlTerm term)
void Pl_Check_For_Un_Callable(PlTerm term)
void Pl_Check_For_Un_Variable(PlTerm term)

The function Pl Check For Un List(term) checks if term can be unified with a list. This test is done for the entire list (not only for the functor/arity of term but also recursively on the tail of the list). The function Pl Check For Un Variable(term) ensures that term is not currently instantiated. These functions can be defined using functions to test the type of a Prolog term (section 10.4.6, page 214) and functions to raise Prolog errors (section 10.5, page 217). For instance Pl Check For Un List(term) is defined as follows:

void Pl_Check_For_Un_List(PlTerm term)
{
 if (!Pl_Builtin_List_Or_Partial_List(term))
     Pl_Err_Type(type_list, term);
}

10.4.5 Creating Prolog terms

These functions are provided to creates Prolog terms. Each function returns a PlTerm containing the created term.

Simple foreign types: for each simple foreign type (section 10.3.2, page 202) there is a creation function:

PlTerm Pl_Mk_Integer (PlLong n)
PlTerm Pl_Mk_Positive(PlLong n)
PlTerm Pl_Mk_Float   (double n)
PlTerm Pl_Mk_Number  (double n)
PlTerm Pl_Mk_Atom    (int atom)
PlTerm Pl_Mk_Boolean (int b)
PlTerm Pl_Mk_Char    (int c)
PlTerm Pl_Mk_In_Char (int c)
PlTerm Pl_Mk_Code    (int c)
PlTerm Pl_Mk_In_Code (int c)
PlTerm Pl_Mk_Byte    (int b)
PlTerm Pl_Mk_In_Byte (int b)
PlTerm Pl_Mk_String  (const char *str)
PlTerm Pl_Mk_Chars   (const char *str)
PlTerm Pl_Mk_Codes   (const char *str)

The function Pl Mk Number(n, term) initializes term with an integer if n is an integer, with a floating point number otherwise. The function Pl Mk String(str) first creates an atom corresponding to str and then returns that Prolog atom (i.e. equivalent to Pl Mk Atom(Pl Create Allocate Atom(str))).

Complex terms: the following functions accept the sub-arguments (terms) of complex terms as an array of PlTerm. Refer to the introduction of this section for more information about the arguments of complex functions (section 10.4.1, page 210).

PlTerm Pl_Mk_Proper_List(int size, const PlTerm *arg)
PlTerm Pl_Mk_List       (PlTerm *arg)
PlTerm Pl_Mk_Compound   (int functor, int arity, const PlTerm *arg)
PlTerm Pl_Mk_Callable   (int functor, int arity, const PlTerm *arg)

10.4.6 Testing the type of Prolog terms

The following functions test the type of a Prolog term. Each function corresponds to a type testing built-in predicate (section 8.1.1, page 55).

PlBool Pl_Builtin_Var                 (PlTerm term)
PlBool Pl_Builtin_Non_Var             (PlTerm term)
PlBool Pl_Builtin_Atom                (PlTerm term)
PlBool Pl_Builtin_Integer             (PlTerm term)
PlBool Pl_Builtin_Float               (PlTerm term)
PlBool Pl_Builtin_Number              (PlTerm term)
PlBool Pl_Builtin_Atomic              (PlTerm term)
PlBool Pl_Builtin_Compound            (PlTerm term)
PlBool Pl_Builtin_Callable            (PlTerm term)
PlBool Pl_Builtin_List                (PlTerm term)
PlBool Pl_Builtin_Partial_List        (PlTerm term)
PlBool Pl_Builtin_List_Or_Partial_List(PlTerm term)
PlBool Pl_Builtin_Fd_Var              (PlTerm term)

     PlBool Pl_Builtin_Non_Fd_Var          (PlTerm term)
     PlBool Pl_Builtin_Generic_Var         (PlTerm term)
     PlBool Pl_Builtin_Non_Generic_Var     (PlTerm term)
     int    Pl_Type_Of_Term                (PlTerm term)
     PlLong   Pl_List_Length                 (PlTerm list)

The function Pl Type Of Term(term) returns the type of term, the following constants can be used to test this type (e.g. in a switch instruction):

• PL PLV: Prolog variable.

• PL FDV: finite domain variable.

• PL INT: integer.

• PL FLT: floating point number.

• PL ATM: atom.

• PL LST: list.

• PL STC: structure

The tag PL LST means a term whose principal functor is '.' and whose arity is 2 (recall that the empty list is the atom []). The tag PL STC means any other compound term.

The function Pl List Length(list) returns the number of elements of the list (0 for the empty list). If list is not a list this function returns -1.

10.4.7 Comparing Prolog terms

The following functions compares Prolog terms. Each function corresponds to a comparison built-in predicate (section 8.3.2, page 57).

     PlBool Pl_Builtin_Term_Eq (PlTerm term1, PlTerm term2)
     PlBool Pl_Builtin_Term_Neq(PlTerm term1, PlTerm term2)
     PlBool Pl_Builtin_Term_Lt (PlTerm term1, PlTerm term2)
     PlBool Pl_Builtin_Term_Lte(PlTerm term1, PlTerm term2)
     PlBool Pl_Builtin_Term_Gt (PlTerm term1, PlTerm term2)
     PlBool Pl_Builtin_Term_Gte(PlTerm term1, PlTerm term2)

All these functions are based on a general comparison function returning a negative integer if term1 is less than term2, 0 if they are equal and a positive integer otherwise:

     PlLong Term_Compare(PlTerm term1, PlTerm term2)

Finally, the following function gives an access to the compare/3 built-in (section 8.3.3, page 58) unifying cmp with the atom <, = or > depending on the result of the comparison of term1 and term2.

     PlBool Pl_Builtin_Compare(PlTerm cmp, PlTerm term1, PlTerm term2)

10.4.8 Term processing

The following functions give access to the built-in predicates: functor/3 (section 8.4.1, page 59), arg/3 (section 8.4.2, page 59) and (=..)/2 (section 8.4.3, page 60).

     PlBool Pl_Builtin_Functor(PlTerm term, PlTerm functor, PlTerm arity)
     PlBool Pl_Builtin_Arg(PlTerm arg_no, PlTerm term, PlTerm sub_term)
     PlBool Pl_Builtin_Univ(PlTerm term, PlTerm list)

The following functions make a copy of a Prolog term:

     void Pl_Copy_Term           (PlTerm *dst_term, const PlTerm *src_term)
     void Pl_Copy_Contiguous_Term(PlTerm *dst_term, const PlTerm *src_term)
     int  Pl_Term_Size           (PlTerm term)

The function Pl Copy Term(dst term, src term) makes a copy of the term located at src term and stores it from the address given by dst term. The result is a contiguous term. If it can be ensured that the source term is a contiguous term (i.e. result of a previous copy) the function Pl Copy Contiguous Term() can be used instead (it is faster). In any case, sufficient space should be available for the copy (i.e. from dst term). The function Pl Term Size(term) returns the number of PlTerm needed by term.

The following function is an utility to display a term to the console, similarly to the built-in predicate write/1 (section 8.14.6, page 108).

     void Pl_Write(PlTerm term)

This Pl Write function can be used for debugging purpose. However, it is more flexible to receive the content of the write/1 as a C string. This can be achieved by the following functions (using repectively write/1, writeq/1, write canonical/1 and display/1 (section 8.14.6, page 108) to obtain a textual representation of the term). These functions return a dynamically allocated C string (using malloc(3)) which can be freed by the user when no longer needed.

     char *Pl_Write_To_String(PlTerm term)
     char *Pl_Writeq_To_String(PlTerm term)
     char *Pl_Write_Canonical_To_String(PlTerm term)
     char *Pl_Display_To_String(PlTerm term)

Finally the following function performs the opposite converstion: given a C string it returns the associated Prolog term. It uses read term/2 (section 8.14.1, page 104) with the option end of term(eof) (thus the C string does not need to terminate by a dot).

     PlTerm Pl_Read_From_String(const char *str)

10.4.9 Comparing and evaluating arithmetic expressions

The following functions compare arithmetic expressions. Each function corresponds to a comparison built-in predicate (section 8.6.3, page 69).

     PlBool Pl_Builtin_Eq (PlTerm expr1, PlTerm expr2)
     PlBool Pl_Builtin_Neq(PlTerm expr1, PlTerm expr2)
     PlBool Pl_Builtin_Lt (PlTerm expr1, PlTerm expr2)
     PlBool Pl_Builtin_Lte(PlTerm expr1, PlTerm expr2)
     PlBool Pl_Builtin_Gt (PlTerm expr1, PlTerm expr2)
     PlBool Pl_Builtin_Gte(PlTerm expr1, PlTerm expr2)

The following function evaluates the expression expr and stores its result as a Prolog number (integer or floating point number) in result:

     void Pl_Math_Evaluate(PlTerm expr, PlTerm *result)

This function can be followed by a read function (section 10.4.3, page 211) to obtain the result.

10.5 Raising Prolog errors

The following functions allows a C function to raise a Prolog error. Refer to the section concerning Prolog errors for more information about the effect of raising an error (section 6.3, page 39).

10.5.1 Managing the error context

When one of the following error function is invoked it refers to the implicit error context (section 6.3.1, page 39). This context indicates the name and the arity of the concerned predicate. When using a foreign/2 declaration this context is set by default to the name and arity of the associated Prolog predicate. This can be controlled using the bip name option (section 10.3.2, page 202). In any case, the following functions can also be used to modify this context:

     void Pl_Set_C_Bip_Name  (const char *functor, int arity)
     void Pl_Unset_C_Bip_Name(void)

The function Pl Set C Bip Name(functor, arity) initializes the context of the error with functor and arity (if arity<0 only functor is significant). The function Pl Unset C Bip Name() removes such an initialization (the context is then reset to the last Functor/Arity set by a call to set bip name/2 (section 8.22.3, page 151). This is useful when writing a C routine to define a context for errors occurring in this routine and, before exiting to restore the previous context.

10.5.2 Instantiation error

The following function raises an instantiation error (section 6.3.2, page 39):

     void Pl Err Instantiation(void)

10.5.3 Uninstantiation error

The following function raises an uninstantiation error (section 6.3.3, page 40):

     void Pl Err Uninstantiation( PlTerm culprit)

10.5.4 Type error

The following function raises a type error (section 6.3.4, page 40):

     void Pl Err Type(int atom type, PlTerm culprit)

atom type is (the internal key of) the atom associated with the expected type. For each type name T there is a corresponding predefined atom stored in a global variable whose name is of the form pl type T . culprit is the argument which caused the error.

Example: x is an atom while an integer was expected: Pl Err Type(pl type integer, x).

10.5.5 Domain error

The following function raises a domain error (section 6.3.5, page 40):

     void Pl Err Domain(int atom domain, PlTerm culprit)

atom domain is (the internal key of) the atom associated with the expected domain. For each domain name D there is a corresponding predefined atom stored in a global variable whose name is of the form domain D . culprit is the argument which caused the error.

Example: x is < 0 but should be ≥ 0: Pl Err Domain(pl domain not less than zero, x).

10.5.6 Existence error

The following function raises an existence error (section 6.3.6, page 41):

     void Pl Err Existence(int atom object, PlTerm culprit)

atom object is (the internal key of) the atom associated with the type of the object. For each object name O there is a corresponding predefined atom stored in a global variable whose name is of the form pl existence O . culprit is the argument which caused the error.

Example: x does not refer to an existing source: Pl Err Existence(pl existence source sink, x).

10.5.7 Permission error

The following function raises a permission error (section 6.3.7, page 41):

     void Pl Err Permission(int atom operation, int atom permission, PlTerm culprit)

atom operation is (the internal key of) the atom associated with the operation which caused the error. For each operation name O there is a corresponding predefined atom stored in a global variable whose name is of the form pl permission operation O . atom permission is (the internal key of) the atom associated with the tried permission. For each permission name P there is a corresponding predefined atom stored in a global variable whose name is of the form pl permission type P . culprit is the argument which caused the error.

Example: reading from an output stream x: Pl Err Permission(pl permission operation input,
pl permission type stream, x).

10.5.8 Representation error

The following function raises a representation error (section 6.3.8, page 41):

     void Pl Err Representation(int atom limit)

atom limit is (the internal key of) the atom associated with the reached limit. For each limit name L there is a corresponding predefined atom stored in a global variable whose name is of the form

pl representation L .
Example: an arity too big occurs: Pl Err Representation(pl representation max arity).

10.5.9 Evaluation error

The following function raises an evaluation error (section 6.3.9, page 42):

     void Pl Err Evaluation(int atom error)

atom error is (the internal key of) the atom associated with the error. For each evaluation error name E there is a corresponding predefined atom stored in a global variable whose name is of the form

pl evaluation E .
Example: a division by zero occurs: Pl Err Evaluation(pl evaluation zero divisor).

10.5.10 Resource error

The following function raises a resource error (section 6.3.10, page 42):

     void Pl Err Resource(int atom resource)

atom resource is (the internal key of) the atom associated with the resource. For each resource error name R there is a corresponding predefined atom stored in a global variable whose name is of the form

pl resource R .
Example: too many open streams: Pl Err Resource(pl resource too many open streams).

10.5.11 Syntax error

The following function raises a syntax error (section 6.3.11, page 42):

     void Pl Err Syntax(int atom error)

atom error is (the internal key of) the atom associated with the error. There is no predefined syntax error atoms.

Example: a / is expected: Pl Err Syntax(Pl Create Atom("/ expected")).

The following function emits a syntax error according to the value of the syntax error Prolog flag (section 8.22.1, page 148). This function can then return (if the value of the flag is either warning or fail). In that case the calling function should fail (e.g. returning PL FALSE). This function accepts a file name (the empty string C "" can be passed), a line and column number and an error message string. Using this function makes it possible to further call the built-in predicate syntax error info/4 (section 8.14.4, page 107):

     void Pl Emit Syntax Error(char *file name, int line, int column, char *message)
Example: a / is expected: Pl Emit Syntax Error("data", 10, 30, "/ expected").

10.5.12 System error

The following function raises a system error (4.3.11, page *):

     void Pl Err System(int atom error)

atom error is (the internal key of) the atom associated with the error. There is no predefined system error atoms.

Example: an invalid pathname is given: Pl Err System(Pl Create Atom("invalid path name")).

The following function emits a system error associated with an operating system error according to the value of the os error Prolog flag (section 8.22.1, page 148). This function can then return (if the value of the flag is either warning or fail). In that case the calling function should fail (e.g. returning PL FALSE).

The following function uses the value of the errno C library variable (basically it calls Pl Err System with the result of strerror(errno)).

     void Pl Os Error(void)

Example: if a call to the C Unix function chdir(2) returns -1 then call Os Error().

10.6 Calling Prolog from C

10.6.1 Introduction

The following functions allows a C function to call a Prolog predicate:

     void   Pl_Query_Begin        (PlBool recoverable)
     int    Pl_Query_Call         (int functor, int arity, PlTerm *arg)
     int    Pl_Query_Start        (int functor, int arity, PlTerm *arg,
                                   PlBool recoverable)
     int    Pl_Query_Next_Solution(void)
     void   Pl_Query_End          (int op)
     PlTerm Pl_Get_Exception      (void)
     void   Pl_Exec_Continuation  (int functor, int arity, PlTerm *arg)
     void   Pl_Throw              (PlTerm ball)

The invocation of a Prolog predicate should be done as follows:

• open a query using Pl Query Begin()

• compute the first solution using Pl Query Call()

• eventually compute next solutions using Pl Query Next Solution()

• close the query using Pl Query End()

The function Pl Query Begin(recoverable) is used to initialize a query. The argument recoverable shall be set to PL TRUE if the user wants to recover, at the end of the query, the memory space consumed by the query (in that case an additional choice-point is created). All terms created in the heap, e.g. using Pl Mk ... family functions (section 10.4.5, page 214), after the invocation of Pl Query Begin() can be recovered when calling Pl Query End(PL TRUE) (see below).

The function Pl Query Call(functor, arity, arg) calls a predicate passing arguments. It is then used to compute the first solution. The arguments functor, arity and arg are similar to those of the functions handling complex terms (section 10.4.1, page 210). This function returns:

• PL FAILURE (a constant equal to PL FALSE, i.e. 0) if the query fails.

• PL SUCCESS (a constant equal to PL TRUE, i.e. 1) in case of success. In that case the argument array arg can be used to obtain the unification performed by the query.

• PL EXCEPTION (a constant equal to 2). In that case function Pl Get Exception() can be used to obtained the exceptional term raised by throw/1 (section 7.2.4, page 53).

The function Pl Query Start(functor, arity, arg, recoverable) is a shorthand equivalent to a call
to Pl Query Begin(recoverable) followed by a call to Pl Query Call(functor, arity, arg).

The function Pl Query Next Solution() is used to compute a new solution. It must be only used if the result of the previous solution was PL SUCCESS. This functions returns the same kind of values as

Pl Query Call() (see above).

The function Pl Query End(op) is used to finish a query. This function mainly manages the remaining alternatives of the query. However, even if the query has no alternatives this function must be used to correctly finish the query. The value of op is:

• PL RECOVER: to recover the memory space consumed by the query. After that the state of Prolog stacks is exactly the same as before opening the query. To use this option the query must have been initialized specifying PL TRUE for recoverable (see above).

• PL CUT: to cut remaining alternatives. The effect of this option is similar to a cut after the query.

• PL KEEP FOR PROLOG: to keep the alternatives for Prolog. This is useful when the query was invoked in a foreign C function. In that case, when the predicate corresponding to the C foreign function is invoked a query is executed and the remaining alternatives are then available as alternatives of that predicate.

Note that several queries can be nested since a stack of queries is maintained. For instance, it is possible to call a query and before terminating it to call another query. In that case the first execution of Pl Query End() will finish the second query (i.e. the inner) and the next execution of Pl Query End() will finish the first query.

The function Pl Exec Continuation(functor, arity, arg) replaces the current calculus by the execution of the specified predicate. The arguments functor, arity and arg are similar to those of the functions handling complex terms (section 10.4.1, page 210).

Finally the function Pl Throw(ball) throws an exception. See the throw/1 control construct for more information on exceptions (section 7.2.4, page 53). Note that Pl Throw(ball) is logically equivalent (but

faster) to Pl Exec Continuation(Pl Find Atom("throw"), 1, &ball) .

10.6.2 Example: my call/1 - a call/1 clone

We here define a predicate my call(Goal) which acts like call(Goal) except that we do not handle exceptions (if an exception occurs the goal simply fails):

In the prolog file examp.pl:

     :- foreign(my call(term)).

In the C file examp c.c:

     #include <string.h>
     #include <gprolog.h>
     PlBool
     my_call(PlTerm goal)
     {
       PlTerm *arg;
       int functor, arity;

  int result;
  arg = Pl_Rd_Callable_Check(goal, &functor, &arity);
  Pl_Query_Begin(PL_FALSE);
  result = Pl_Query_Call(functor, arity, arg);
  Pl_Query_End(PL_KEEP_FOR_PROLOG);
  return (result == PL_SUCCESS);
}

The compilation produces an executable called examp:

% gplc examp.pl examp c.c

Examples of use:

| ?- my call(write(hello)).
hello
| ?- my call(for(X,1,3)).

X = 1 ?  (here the user presses ; to compute another solution)
X = 2 ?  (here the user presses ; to compute another solution)
X = 3  (here the user is not prompted since there is no more alternative)

| ?- my call(1).
{exception: error(type error(callable,1),my call/1)}
| ?- my call(call(1)).
no

When my call(1) is called an error is raised due to the use of Pl Rd Callable Check(). However the error raised by my call(call(1)) is ignored and PL FALSE (i.e. a failure) is returned by the foreign function.

To really simulate the behavior of call/1 when an exception is recovered it should be re-raised to be captured by an earlier handler. The idea is then to execute a throw/1 as the continuation. This is what it is done by the following code:

#include <string.h>
#include <gprolog.h>
PlBool
my_call(PlTerm goal)
{
  PlTerm *args;
  int functor, arity;
  int result;
  args = Pl_Rd_Callable_Check(goal, &functor, &arity);
  Pl_Query_Begin(PL_FALSE);
  result = Pl_Query_Call(functor, arity, args);
  Pl_Query_End(PL_KEEP_FOR_PROLOG);
  if (result == PL_EXCEPTION)
    {

    PlTerm except = Pl_Get_Exception();
    Pl_Throw(except);
    // equivalent to Pl_Exec_Continuation(Find_Atom("throw"), 1, &except);
  }
return result;
}

The following code propagates the error raised by call/1.

| ?- my call(call(1)).
{exception: error(type error(callable,1),my call/1)}

Finally note that a simpler way to define my call/1 is to use Pl Exec Continuation() as follows:

#include <string.h>
#include <gprolog.h>
PlBool
my_call(PlTerm goal)
{
PlTerm *args;
int functor, arity;
args = Pl_Rd_Callable_Check(goal, &functor, &arity);
Pl_Exec_Continuation(functor, arity, args);
return PL_TRUE;
}

10.6.3 Example: recovering the list of all operators

We here define a predicate all op(List) which unifies List with the list of all currently defined operators

as would be done by: findall(X,current op( , ,X),List).

In the prolog file examp.pl:

:- foreign(all op(term)).

In the C file examp c.c:

#include <string.h>
#include <gprolog.h>
PlBool
all_op(PlTerm list)
{
PlTerm op[1024];
PlTerm args[3];
int n = 0;
int result;
Pl_Query_Begin(PL_TRUE);
args[0] = Pl_Mk_Variable();
args[1] = Pl_Mk_Variable();
args[2] = Pl_Mk_Variable();
result = Pl_Query_Call(Find_Atom("current_op"), 3, args);

       while (result)
         {
           op[n++] = Pl_Mk_Atom(Pl_Rd_Atom(args[2])); /* arg[2]: the name of the op */
           result = Pl_Query_Next_Solution();
         }
       Pl_Query_End(PL_RECOVER);
       return Pl_Un_Proper_List_Check(n, op, list);
     }

Note that we know here that there is no source for exception. In that case the result of Pl Query Call and Pl Query Next Solution can be considered as a boolean.

The compilation produces an executable called examp:

     % gplc examp.pl examp c.c

Example of use:

     | ?- all_op(L).
     L = [:-,:-,\=,=:=,#>=,#<#,@>=,-->,mod,#>=#,**,*,+,+,',',...]
     | ?- findall(X,current_op(_,_,X),L).
     L = [:-,:-,\=,=:=,#>=,#<#,@>=,-->,mod,#>=#,**,*,+,+,',',...]

10.7 Defining a new C main() function

GNU Prolog allows the user to define his own main() function. This can be useful to perform several tasks before starting the Prolog engine. To do this simply define a classical main(argc, argv) function. The following functions can then be used:

     int    Pl_Start_Prolog         (int argc, char *argv[])
     void   Pl_Stop_Prolog          (void)
     void   Pl_Reset_Prolog         (void)
     PlBool Pl_Try_Execute_Top_Level(void)

The function Pl Start Prolog(argc, argv) initializes the Prolog engine (argc and argv are the commandline variables). This function collects all linked objects (issued from the compilation of Prolog files) and initializes them. The initialization of a Prolog object file consists in adding to appropriate tables new atoms, new predicates and executing its system directives. A system directive is generated by the Prolog to WAM compiler to reflect a (user) directive executed at compile-time such as op/3 (section 7.1.11, page 49). Indeed, when the compiler encounters such a directive it immediately executes it and also generates a system directive to execute it at the start of the executable. When all system directives have been executed the Prolog engine executes all initialization directives defined with initialization/1 (section 7.1.14, page 50). The function returns the number of user directives (i.e. initialization/1) executed. This function must be called only once.

The function Pl Stop Prolog() stops the Prolog engine. This function must be called only once after all Prolog treatment have been done.

The function Pl Reset Prolog() reinitializes the Prolog engine (i.e. reset all Prolog stacks).

The function Pl Try Execute Top Level() executes the top-level if linked (section 4.4.3, page 23) and returns PL TRUE. If the top-level is not present the functions returns PL FALSE.

Here is the definition of the default GNU Prolog main() function:

static int
Main_Wrapper(int argc, char *argv[])
{
  int nb_user_directive;
  PlBool top_level;
  nb_user_directive = Pl_Start_Prolog(argc, argv);
  top_level = Pl_Try_Execute_Top_Level();
  Pl_Stop_Prolog();
  if (top_level || nb_user_directive)
    return 0;
  fprintf(stderr,
          "Warning: no initial goal executed\n"
          "   use a directive :- initialization(Goal)\n"
          "   or remove the link option --no-top-level"
          " (or --min-bips or --min-size)\n");
  return 1;
}
int
main(int argc, char *argv[])
{
  return Main_Wrapper(argc, argv);
}

Note that under some circumstances it is necessary to encapsulate the code of main() inside an intermediate function called by main(). Indeed, some C compilers (e.g. gcc) treats main() particularly, producing an incompatible code w.r.t GNU Prolog. So it is a good idea to always use a wrapper function as shown above.

10.7.1 Example: asking for ancestors

In this example we use the following Prolog code (in a file called new main.pl):

parent(bob,   mary).
parent(jane,  mary).
parent(mary,  peter).
parent(paul,  peter).
parent(peter, john).
anc(X, Y):-
        parent(X, Y).
anc(X, Z) :-
        parent(X, Y),
        anc(Y, Z).

The following file (called new main c.c) defines a main() function read the name of a person and displaying all successors of that person. This is equivalent to the Prolog query: anc(Result, Name).

static int
Main_Wrapper(int argc, char *argv[])
{
int func;
PlTerm arg[10];
char str[100];
char *sol[100];
int i, nb_sol = 0;
PlBool res;
Pl_Start_Prolog(argc, argv);
func = Pl_Find_Atom("anc");
for (;;)
  {
    printf("\nEnter a name (or 'end' to finish): ");
    fflush(stdout);
    scanf("%s", str);
    if (strcmp(str, "end") == 0)
break;
    Pl_Query_Begin(PL_TRUE);
    arg[0] = Pl_Mk_Variable();
    arg[1] = Pl_Mk_String(str);
    nb_sol = 0;
    res = Pl_Query_Call(func, 2, arg);
    while (res)
{
sol[nb_sol++] = Pl_Rd_String(arg[0]);
res = Pl_Query_Next_Solution();
}
    Pl_Query_End(PL_RECOVER);
    for (i = 0; i < nb_sol; i++)
printf("  solution: %s\n", sol[i]);
    printf("%d solution(s)\n", nb_sol);
  }
Pl_Stop_Prolog();
return 0;
}
int
main(int argc, char *argv[])
{
return Main_Wrapper(argc, argv);
}

The compilation produces an executable called new main:

% gplc new main.pl new main c.c

Examples of use:

Enter a name (or 'end' to finish): john
solution: peter
solution: bob
solution: jane
solution: mary
solution: paul
5 solution(s)
Enter a name (or 'end' to finish): mary
solution: bob
solution: jane
2 solution(s)
Enter a name (or 'end' to finish): end

228  10 INTERFACING PROLOG AND C

References

[1] H. Aït-Kaci. “Warren’s Abstract Machine, A Tutorial Reconstruction”.

Logic Programming Series, MIT Press, 1991.

http://web.archive.org/web/20071225092145/www.vanx.org/archive/wam/wam.html

[2] W.F. Clocksin and C.S. Mellish. Programming in Prolog, Springer-Verlag, 1981.

[3] P. Codognet and D. Diaz. “wamcc: Compiling Prolog to C”.

In 12th International Conference on Logic Programming, Tokyo, Japan, MIT Press, 1995.

http://cri-dist.univ-paris1.fr/diaz/publications/WAMCC/iclp95.pdf

[4] P. Codognet and D. Diaz. “Compiling Constraint in clp(FD)”.

Journal of Logic Programming, Vol. 27, No. 3, June 1996.

http://cri-dist.univ-paris1.fr/diaz/publications/CLP-FD/jlp96.pdf

[5] D. Diaz and P. Codognet. “Design and Implementation of the GNU Prolog System”. Journal of Functional and Logic Programming, Vol. 2001, No. 6, October 2001.

http://cri-dist.univ-paris1.fr/diaz/publications/GNU-PROLOG/jflp01.pdf

[6] Information technology - Programming languages - Prolog - Part 1: General Core. ISO/IEC 13211-1, 1995.

[7] J. Jaffar and J-L. Lassez. “Constraint Logic Programming”.

In Principles Of Programming Languages, Munich, Germany, January 1987.

[8] P. Van Hentenryck. “Constraint Satisfaction in Logic Programming”.

Logic Programming Series, The MIT Press, 1989.

[9] D. H. D. Warren. “An Abstract Prolog Instruction Set”.

Technical Report 309, SRI International, Oct. 1983.

230  REFERENCES

\underline{INDEX 231}

Index

!/0, 51, 52

'.'/2, 153
(',')/2, 51
(*->)/2, 51
(-->)/2, 123
(->)/2, 51
(;)/2, 51
(=)/2, 56
(=..)/2, 60
(=:=)/2, 69

(==)/2, 57, 136

(=<)/2, 69
(=\=)/2, 69
(@=<)/2, 57
(@<)/2, 57
(@>)/2, 57
(@>=)/2, 57

(#/\)/2 (FD), 193 (#=)/2 (FD), 189

(#==>)/2 (FD), 193

(#=#)/2 (FD), 190 (#=<)/2 (FD), 189

(#=<#)/2 (FD), 190

(##)/2 (FD), 193 (#<)/2 (FD), 189

(#<=>)/2 (FD), 193

(#<#)/2 (FD), 190 (#>)/2 (FD), 189 (#>=)/2 (FD), 189

(#>=#)/2 (FD), 190

(#>#)/2 (FD), 190 (#\)/1 (FD), 193 (#\/)/2 (FD), 193

(#\/\)/2 (FD), 193

(#\=)/2 (FD), 189

(#\==>)/2 (FD), 193
(#\=#)/2 (FD), 190
(#\<=>)/2 (FD), 193
(#\\/)/2 (FD), 193
(is)/2, 69
(<)/2, 69
(>)/2, 69
(>=)/2, 69
(\+)/1, 126
(\=)/2, 57
(\==)/2, 57

--, 13

--assembly, 23
--aux-father, 29
--aux-father2, 29
--c-compiler, 24

                                               --cmd-line, 29
                                               --comment, 24, 24
                                               --compile-msg, 24
                                               --consult-file, 13
                                               --cstr-size, 25
                                               --decode, 29
                                               --demangling, 29
                                               --encode, 29
                                               --entry-goal, 13
                                               --fast-math, 24, 68
                                               --fd-to-c, 23
                                               --fixed-sizes, 21, 25
                                               --foreign-only, 24
                                               --global-size, 25
                                               --gui-console, 25

--help, 13, 23, 29

                                               --init-goal, 13, 28
                                               --keep-void-inst, 24
                                               --linker, 25
                                               --local-size, 21, 25
                                               --mangling, 29
                                               --max-atom, 25
                                               --min-bips, 25
                                               --min-fd-bips, 25
                                               --min-pl-bips, 25
                                               --min-reg-opt, 24
                                               --min-size, 25
                                               --mini-assembly, 23
                                               --new-top-level, 25, 28
                                               --no-call-c, 24
                                               --no-debugger, 25, 25
                                               --no-del-temp, 23
                                               --no-demangling, 23
                                               --no-fd-lib, 25
                                               --no-inline, 24
                                               --no-opt-last-subterm, 24
                                               --no-redef-error, 24
                                               --no-reg-opt, 24
                                               --no-reorder, 24
                                               --no-singl-warn, 24
                                               --no-susp-warn, 24
                                               --no-top-level, 25
                                               --object, 23
                                               --output, 23
                                               --pic, 24
                                               --pl-state, 24, 152
                                               --printf, 29
                                               --query-goal, 13
                                               --relax, 29
                                               --statistics, 24
                                               --strip, 25

\underline{232 INDEX}

--temp-dir, 23
--trail-size, 25
--verbose, 23
--version, 13, 23, 29
--wam-comment, 24
--wam-for-byte-code, 23
--wam-for-native, 23

-A, 24 -C, 24 -D, 29 -E, 29 -F, 23 -L, 25 -M, 23, 29 -P, 29 -S, 23 -W, 23 -c, 23 -h, 23 -o, 23 -s, 25 -v, 23 -w, 23

abolish/1, 73

abort/0, 18, 34, 126

absolute file name (property), 165

absolute file name/2, 48, 81, 153, 154, 158,

159, 160, 162–167

acyclic term/1, 61
add linedit completion/1, 181
add stream alias/2, 78, 90
add stream mirror/2, 79, 91
address bits (flag), 149

alias (option), 81 alias (property), 84 append (mode), 81

append/1, 121
append/3, 135

arch (flag), 149

architecture/1, 168
arg/3, 59

argument selector, 142

argument counter/1, 160
argument list/1, 14, 161
argument value/2, 14, 160

argv (flag), 149

asserta/1, 71
assertz/1, 71
at end of stream/0, 85
at end of stream/1, 85
atom/1, 55
atom chars/2, 131
atom codes/2, 131

                                               atom concat/3, 129
                                               atom length/2, 129
                                               atom property/2, 134
                                               atomic/1, 55

back quotes (flag), 14, 104, 150, 152 back quotes (token), 106 backtracks (FD option), 199

                                               bagof/3, 77
                                               between/3, 128

binary (option), 81, 92

                                               bind variables/2, 64

bip name (option), 203, 217 bisect (FD option), 198 block (option), 82, 93 block device (permission), 165 bof (whence), 87 boolean (option), 202, 204 bounded (flag), 149 bounds (FD option), 198 break/0, 18, 34, 126 buffering (option), 81 buffering (property), 84 built in (property), 48, 75 built in/0 (directive), 47 built in/1 (directive), 47 built in fd (property), 48, 75 built in fd/0 (directive), 47 built in fd/1 (directive), 47

c cc (flag), 149

                                               c cc version data (flag), 149

c cflags (flag), 149 c ldflags (flag), 149

                                               call/1, 52
                                               call/2-11, 126
                                               call det/2, 126
                                               call with args/1-11, 126
                                               callable/1, 55

catch/3, 31, 39, 53

                                               change directory/1, 162
                                               char code/2, 130, 209

char conversion (flag), 104, 116, 150, 152

                                               char conversion/2 (directive), 49
                                               char conversion/2, 50, 116
                                               character count/2, 87

character device (permission), 165 choice size (option), 203, 205

                                               clause/2, 73
                                               close/1, 82

close/2, 82, 171, 177

                                               close input atom stream/1, 95
                                               close input chars stream/1, 95
                                               close input codes stream/1, 95
                                               close output atom stream/2, 96

close output chars stream/2, 96
close output codes stream/2, 96
compare/3, 58
compiled at (flag), 149

completion, 19, 181

compound/1, 55

consult/1, 17, 21, 23, 153 control constructs, 51

control construct (property), 75
copy term/2, 60
cpu time/1, 155
create pipe/2, 172

current (whence), 87

current alias/2, 90
current atom/1, 134
current bip name/2, 39, 151
current char conversion/2, 116
current input/1, 79
current mirror/2, 92
current op/3, 115
current output/1, 79
current predicate/1, 72, 74
current prolog flag/2, 151
current stream/1, 83
date time/1, 167

debug (flag), 150 debug/0 (debug), 17, 18, 31 debugging/0 (debug), 31, 34

decompose file name/4, 159

Definite clause grammars, see DCG

delete/3, 136
delete directory/1, 162
delete file/1, 163

demangling, 26 dialect (flag), 149 directory (permission), 165

directory files/2, 162
discontiguous/1 (directive), 46
display/1, 108

display/2, 108, 119, 120

display to atom/2, 118
display to chars/2, 119
display to codes/2, 120

double quotes (flag), 104, 150, 152 dynamic (property), 75 dynamic/1 (directive), 45, 70

elif/1 (directive), 48 else/0 (directive), 48 end of stream (property), 84 end of term (option), 104 endif/0 (directive), 48

ensure linked/1 (directive), 47
ensure loaded/1 (directive), 49

                                               environ/2, 161

eof (whence), 87

eof action (option), 81 eof action (property), 84 eof code (option), 81, 93 error (option), 81, 93, 104 escape sequence, 14, 134, 150, 150 exclude (option), 65

                                               exec/4, 171
                                               exec/5, 171

execute (permission), 164

                                               expand term/2, 124

extended (token), 106

extra-constrained, see extra cstr extra cstr (FD), 183, 188

fail (option), 104 fail/0, 51, 127

                                               false/0, 126

fct name (option), 202

                                               fd all different/1 (FD), 195
                                               fd at least one/1 (FD), 194
                                               fd at most one/1 (FD), 194
                                               fd atleast/3 (FD), 196
                                               fd atmost/3 (FD), 196
                                               fd cardinality/2 (FD), 194, 196
                                               fd cardinality/3 (FD), 194

fd dom/2 (FD), 187

                                               fd domain/2 (FD), 186
                                               fd domain/3 (FD), 185
                                               fd domain bool/1 (FD), 185
                                               fd element/3 (FD), 195
                                               fd element var/3 (FD), 196
                                               fd exactly/3 (FD), 196
                                               fd has extra cstr/1 (FD), 188
                                               fd has vector/1 (FD), 188
                                               fd labeling/1 (FD), 198
                                               fd labeling/2 (FD), 198, 199
                                               fd labelingff/1 (FD), 198

fd max/2 (FD), 187

                                               fd max integer (FD), 183, 184
                                               fd max integer/1 (FD), 184
                                               fd maximize/2 (FD), 199

fd min/2 (FD), 187

                                               fd minimize/2 (FD), 199
                                               fd not prime/1 (FD), 191
                                               fd only one/1 (FD), 194
                                               fd prime/1 (FD), 191
                                               fd reified in/4 (FD), 192
                                               fd relation/2 (FD), 197
                                               fd relationc/2 (FD), 197
                                               fd set vector max/1 (FD), 183, 185
                                               fd size/2 (FD), 187
                                               fd use vector/1 (FD), 188

fd var/1 (FD), 187

fd vector max/1 (FD), 183, 185

fifo (permission), 165

file exists/1, 164

file name (property), 84

file permission/2, 164
file property/2, 165
find linedit completion/2, 181
findall/3, 76
findall/4, 76

first fail (FD option), 198 flag, see Prolog flag

flatten/2, 138
float/1, 55
flush output/0, 83
flush output/1, 79, 83
for/3, 128
forall/2, 126

force (option), 83 foreign/1 (directive), 50, 202 foreign/2 (directive), 50, 202

fork prolog/1, 172
format/2, 110

format/3, 110, 119, 120

format to atom/3, 118
format to chars/3, 119
format to codes/3, 120

from (option), 64 full (debug), 32

functor/3, 59

g array (global var.), 142 g array auto (global var.), 142 g array extend (global var.), 142

g array size/2, 144
g assign/2, 143
g assignb/2, 143
g dec/1, 144
g dec/2, 144
g dec/3, 144
g deco/2, 144
g inc/1, 144
g inc/2, 144
g inc/3, 144
g inco/2, 144
g link/2, 143
g read/2, 143
g reset bit/2, 145
g set bit/2, 145
g test reset bit/2, 145
g test set bit/2, 145
generic var/1 (FD), 187
get/1, 122
get0/1, 122

                                               get byte/1, 101
                                               get byte/2, 78, 101
                                               get char/1, 97
                                               get char/2, 97
                                               get code/1, 97
                                               get code/2, 97, 98
                                               get key/1, 98
                                               get key/2, 98
                                               get key no echo/1, 98
                                               get key no echo/2, 98
                                               get linedit prompt/1, 180
                                               get print stream/1, 113
                                               get seed/1, 156

gplc, 23, 26, 28, 29, 152

                                               ground/1, 55

GUI console, 20, 25

half (debug), 32 halt/0, 14, 18, 126

                                               halt/1, 126

hash (property), 134 hash code, 62, 134

                                               hexgplc, 29

home (flag), 149 host (flag), 149 host cpu (flag), 149

                                               host name/1, 168

host os (flag), 149

                                               host vendor (flag), 149
                                               hostname address/2, 179

if/1 (directive), 48

ignore ops (option), 108 include/1 (directive), 48 infix op (property), 134 initialization/1 (directive), 26, 27, 50, 224 input (property), 84

                                               integer/1, 55
                                               integer rounding function (flag), 68, 149

interpreter, see top-level

                                               is absolute file name/1, 158
                                               is list/1, 55
                                               is relative file name/1, 158

jump (option), 202, 204

                                               keysort/1, 140
                                               keysort/2, 140

largest (FD option), 198

                                               last/2, 138
                                               last modification (property), 165
                                               last read start line column/2, 107

leash/1 (debug), 32, 33 length (property), 134

length/2, 139

line (option), 82, 93

line count/2, 88, 89
line position/2, 88

linedit, 18, 98, 180, 181

list/1, 55
list or partial list/1, 55
listing/0, 154

listing/1, 34, 112, 154 load/1, 17, 23, 25, 153 loose (debug), 32

lower upper/2, 130

MA, 22 main(), 28, 224

make directory/1, 162
maplist/2-8, 140

max (FD option), 198 max arity (flag), 149 max atom (flag), 149 max depth (option), 109 max integer (flag), 149, 183

max list/2, 139

max regret (FD option), 198 max unget (flag), 100, 103, 149

member/2, 135
memberchk/2, 135
meta predicate (property), 75

middle (FD option), 198 min (FD option), 198

min integer (flag), 149
min list/2, 139

mini-assembly, 11, 22, 29 mirror (option), 81 mirror (property), 84 mode (property), 84 monofile (property), 75

most constrained (FD option), 198
msort/1, 140
msort/2, 140

multifile (property), 75 multifile/1 (directive), 46, 47

multifile warning (flag), 150, 152

name demangling, 29 name mangling, 26, 28

name/2, 133
name query vars/2, 64
name singleton vars/1, 63, 112

namevars (option), 16, 64, 108 native code (property), 75 needs quotes (property), 134 needs scan (property), 134

new atom/1, 133
new atom/2, 133

next (option), 64

                                               nl/0, 100
                                               nl/1, 100

nodebug/0 (debug), 31, 34

                                               non fd var/1 (FD), 187
                                               non generic var/1 (FD), 187

none (debug), 32

none (option), 81, 93, 202, 204

                                               nonvar/1, 55

nospy/1 (debug), 32, 34

                                               nospyall/0 (debug), 32

notrace/0 (debug), 31

                                               nth/3, 139
                                               number/1, 55
                                               number atom/2, 132
                                               number chars/2, 132
                                               number codes/2, 132

numbervars (option), 16, 64, 108

                                               numbervars/1, 64, 112
                                               numbervars/3, 64
                                               once/1, 126

op/3 (directive), 49

op/3, 49, 113

                                               open/3, 81

open/4, 78, 81, 92, 93, 175

                                               open input atom stream/2, 94
                                               open input chars stream/2, 94
                                               open input codes stream/2, 94
                                               open output atom stream/1, 96
                                               open output chars stream/1, 96
                                               open output codes stream/1, 96

os error (flag), 150, 220

                                               os version/1, 168

output (property), 84

                                               partial list/1, 55
                                               peek byte/1, 102
                                               peek byte/2, 102
                                               peek char/1, 99
                                               peek char/2, 99
                                               peek code/1, 99
                                               peek code/2, 99

permission (property), 165

                                               permutation/2, 137
                                               phrase/2, 125
                                               phrase/3, 125
                                               popen/3, 78, 170
                                               portray/1, 113
                                               portray clause/1, 112
                                               portray clause/2, 112, 154

portrayed (option), 109 position (property), 84 postfix op (property), 134

                                               predicate property/2, 75

prefix/2, 137

prefix op (property), 134 print/1, 108, 111 print/2, 108, 113, 119, 120

print to atom/2, 118
print to chars/2, 119
print to codes/2, 120

priority (option), 109 private (property), 75 Prolog flag, 14, 38, 50, 68, 74, 100, 103, 104, 106,

116, 148, 151, 152, 183, 219, 220

prolog copyright (flag), 149
prolog date (flag), 149

prolog file (property), 75

prolog file name/2, 153, 159

prolog line (property), 75

prolog name (flag), 149
prolog pid/1, 173
prolog version (flag), 149

PrologScript, 18, 153 public (property), 75 public/1 (directive), 45, 71 punct (token), 106

put/1, 123
put byte/1, 103
put byte/2, 103
put char/1, 100
put char/2, 100
put code/1, 100
put code/2, 100

quoted (option), 16, 108

random (FD option), 198

random/1, 157
random/3, 157
randomize/0, 156

read (mode), 81 read (permission), 164 read/1, 104, 107 read/2, 104, 107, 117, 118

read atom/1, 105, 107

read atom/2, 105, 107, 116

read from atom/2, 117
read from chars/2, 117
read from codes/2, 118
read integer/1, 105, 107
read integer/2, 105, 107, 116
read number/1, 105, 107
read number/2, 105, 107, 116
read pl state file/1, 152
read term/2, 104, 107

read term/3, 104, 107, 116–118

read term from atom/3, 14, 104, 117
read term from chars/3, 117

                                               read term from codes/3, 118
                                               read token/1, 106, 107

read token/2, 106, 107, 116–118

                                               read token from atom/2, 117
                                               read token from chars/2, 117
                                               read token from codes/2, 118

real file name (property), 165

                                               real time/1, 155

regular (permission), 165

                                               remove stream mirror/2, 79, 91, 91
                                               rename file/2, 163

reorder (FD option), 198

                                               repeat/0, 127

reposition (option), 81 reposition (property), 84 reset (option), 81, 93

                                               retract/1, 72
                                               retractall/1, 72

return (option), 202, 204

                                               reverse/2, 136

search (permission), 164

                                               see/1, 121
                                               seeing/1, 121
                                               seek/4, 86
                                               seen/0, 122
                                               select/3, 136

select/5, 79, 175, 178, 179

                                               send signal/2, 174
                                               set bip name/2, 39, 151, 217
                                               set input/1, 78, 80
                                               set linedit prompt/1, 180
                                               set output/1, 78, 80
                                               set prolog flag/2 (directive), 50
                                               set prolog flag/2, 50, 148
                                               set seed/1, 156
                                               set stream buffering/2, 79, 93, 175, 176
                                               set stream eof action/2, 93
                                               set stream line column/3, 89
                                               set stream position/2, 79, 86
                                               set stream type/2, 92, 176
                                               setarg/3, 63
                                               setarg/4, 63
                                               setof/3, 77

shebang support, 18, 153

                                               shell/0, 169
                                               shell/1, 169
                                               shell/2, 169
                                               singleton warning (flag), 150, 152

singletons (option), 64, 65, 104

size (property), 165

                                               skip/1, 122
                                               sleep/1, 174

smallest (FD option), 198 socket (permission), 165

socket/2, 176
socket accept/3, 179
socket accept/4, 179
socket bind/2, 177
socket close/1, 176
socket connect/4, 78, 177, 178
socket listen/2, 178

soft-cut, 52

sort/1, 140
sort/2, 140

space args (option), 109

spawn/2, 170
spawn/3, 170

spy/1 (debug), 32, 34

spypoint condition/3 (debug), 32, 34

standard (FD option), 198 static (property), 75

statistics/0, 155
statistics/2, 155
stop/0, 126
stream line column/3, 89
stream position/2, 85, 86
stream property/2, 84, 85, 86, 91, 92

strict iso (flag), 38, 74, 150 string (token), 106

sub atom/5, 129
sublist/2, 138
subsumes term/2, 61
subtract/3, 136
succ/2, 70
suffix/2, 137
sum list/2, 139
suspicious warning (flag), 150, 152

syntax error (flag), 104, 150, 219 syntax error (option), 104

syntax error info/4, 107, 219
system/1, 169
system/2, 169
system time/1, 155
tab/1, 123
tell/1, 121
telling/1, 121
temporary file/3, 167
temporary name/2, 166
term expansion/2, 124
term hash/2, 62
term hash/4, 62
term ref/2, 65
term variables/2, 61
term variables/3, 61

text (option), 81, 92 throw/1, 31, 39, 53, 220, 221 tight (debug), 32

                                               told/0, 122

top-level, 13, 18, 25, 27, 126, 180, 181, 224

                                               top level/0, 13, 126

trace/0 (debug), 17, 18, 31

                                               true/0, 51

type (option), 81

type (property), 84, 165

                                               unget byte/1, 102
                                               unget byte/2, 102
                                               unget char/1, 100
                                               unget char/2, 100
                                               unget code/1, 100
                                               unget code/2, 100
                                               unify with occurs check/2, 56

unix (flag), 149 unknown (flag), 150 unknown (permission), 165

                                               unlink/1, 163

user (property), 75

user, 120, 121, 153, 158, 159

                                               user error, 78

user input, 78, 82, 120–122 user output, 78, 82, 120–122

                                               user time/1, 155

value method (FD option), 198

var (token), 106

                                               var/1, 55
                                               variable method (FD option), 198

variable names (option), 64, 65, 104, 108 variables (option), 104 vector max (FD), 183, 185, 191 version (flag), 149

                                               version data (flag), 149
                                               wait/2, 173

WAM, 11, 21, 22, 35

wam debug/0 (debug), 31, 35

warning (option), 104

Warren Abstract Machine, see WAM

                                               working directory/1, 162

write (mode), 81

write (permission), 164

write/1, 108, 111 write/2, 108, 119, 120

                                               write canonical/1, 108, 111
                                               write canonical/2, 108, 119, 120
                                               write canonical to atom/2, 118
                                               write canonical to chars/2, 119
                                               write canonical to codes/2, 120
                                               write pl state file/1, 25, 152
                                               write term/2, 108

write term/3, 16, 33, 108, 119, 120

write term to atom/3, 118
write term to chars/3, 119
write term to codes/3, 120
write to atom/2, 118
write to chars/2, 119
write to codes/2, 120
writeq/1, 108, 111

writeq/2, 108, 119, 120, 211

writeq to atom/2, 118
writeq to chars/2, 119
writeq to codes/2, 120